### v1

# Efficient Approximate Degenerate Ordered Statistics Decoding for Quantum Codes via Reliable Subset Reduction

Abstract
Efficient decoding of quantum codes is crucial for achieving high-performance quantum error correction. In this paper, we introduce the concept of approximate degenerate decoding and integrate it with ordered statistics decoding (OSD). Previously, we proposed a reliability metric that leverages both hard and soft decisions from the output of belief propagation (BP), which is particularly useful for identifying highly reliable subsets of variables. Using the approach of reliable subset reduction, we reduce the effective problem size. Additionally, we identify a degeneracy condition that allows high-order OSD to be simplified to order-0 OSD. By integrating these techniques, we present an ADOSD algorithm that significantly improves OSD efficiency in the code capacity noise model. We demonstrate the effectiveness of our BP+ADOSD approach through extensive simulations on a varity of quantum codes, including generalized hypergraph-product codes, topological codes, lift-connected surface codes, and bivariate bicycle codes. The results indicate that the BP+ADOSD decoder outperforms existing methods, achieving higher error thresholds and enhanced performance at low error rates. Additionally, we validate the efficiency of our approach in terms of computational time, demonstrating that ADOSD requires, on average, the same amount of time as two to three BP iterations on surface codes at a depolarizing error rate of around $1\%$. All the proposed algorithms are compared using single-threaded CPU implementations.
I Introduction
Reliable quantum communication is a critical area of research, essential for scaling quantum systems, enabling the exchange of quantum information across distances, and facilitating multiparty protocols. Quantum states are inherently fragile, requiring the implementation of quantum error correction to mitigate the effects of noise and decoherence [2, 3]. In this work, we explore advanced decoding techniques for quantum error-correcting codes to enhance the reliability and efficiency of quantum communication.
Quantum stabilizer codes, analogous to classical linear block codes, allow efficient encoding and binary syndrome decoding [4, 5, 6]. A notable class of stabilizer codes is quantum low-density parity-check (LDPC) codes, which are preferred for their high code rates and feasible syndrome measurements from low-weight stabilizers. These codes can be decoded using belief propagation (BP), similar to classical LDPC codes [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. A quantum LDPC code of the Calderbank–Shor–Steane (CSS) type [20, 21] can be decoded by the BP algorithm for binary codes (BP2) by treating $X$ and $Z$ errors separately. When $X$ and $Z$ errors exhibit correlations, such as in depolarizing channels, a quaternary BP algorithm (BP4) could be more effective, as it better exploits the error correlations during decoding. An important feature of quantum codes is their binary error syndromes, unlike classical nonbinary codes. Consequently, the complexity of BP4 can be reduced by restricting message passing in the Tanner graph [12] to scalar messages, which encode the likelihood ratios associated with Pauli commutation relations.
The decoding problem for quantum codes with high degeneracy is particularly challenging for BP, as seen in topological codes [22, 23, 24, 25, 26, 27, 28]. To address these challenges, several remedies for the BP algorithm have been introduced, including techniques such as random perturbation [8], enhanced feedback [9, 10], check-matrix augmentation [11], message normalization and offset [12, 13], and trapping set analysis [29, 30]. These methods aim to improve the decoding accuracy and convergence of the BP algorithm. A recently developed modified BP4 algorithm, known as MBP4, incorporates additional memory effects and has demonstrated the ability to handle topological codes [31]. The adaptive version of MBP4, termed AMBP4, further improves BP performance while maintaining nearly linear time complexity [31]. Nonetheless, its practical application involves an unknown parameter search and may not be fully parallelized.
We investigate ordered statistic decoding (OSD) in quantum coding theory as a complementary approach to BP. Fossorier and Lin [32] first proposed OSD in classical coding theory as an approach to approximate maximum-likelihood decoding [33, 34, 35, 36, 37, 38, 39, 40, 41]. The OSD algorithm searches for possible solutions by sorting the reliability measures of symbols while maintaining controllable complexity [32, 42, 43, 44, 45]. In cases where BP fails to find a valid error for a given syndrome, syndrome-based OSD can be utilized [46]. Moreover, OSD has been demonstrated to be particularly useful for degenerate quantum codes by Panteleev and Kalachev [42].
The core function of OSD is Gaussian elimination on the parity-check matrix of dimension $m\times n$, which identifies the independent unreliable bits and has a complexity of ${\cal O}(nm^{2})$. The remaining reliable bits can be used to regenerate the most probable error. In an order-$w$ OSD, the set of reliable bits can be flipped up to $w$ bits, generating a list of probable solutions. However, this technique incurs at least a cubic time complexity overhead from Gaussian elimination, along with additional costs for high-order examination, despite its effectiveness.
Fossorier et al[46] identified specific candidate pruning conditions to reduce ineffective high-order OSD processing, as such operations do not yield improved error candidates. Many techniques have been proposed in classical coding theory to reduce the complexity of high-order OSD, by focusing on more probable candidates, thereby reducing the number of OSD candidates required [47, 48, 49, 50, 51, 52].
[…]
However, there is a great deal of room for improvement over the current postprocessing techniques in the literature under the code capacity noise model. These methods primarily rely on binary BP to handle quaternary errors, neglecting error correlations, which suggests that binary decoders are inherently suboptimal. In contrast, the quaternary AMBP4 decoder in [31] has been shown to outperform all the aforementioned decoders without requiring postprocessing. For supporting examples, please refer to Figure 2, Figure 4, and Figure 7 in the simulation section.
Since OSD acts as a complementary strategy to BP for cases where BP fails, BP-OSD is naturally expected to surpass the performance of AMBP4. In this paper, we propose a BP4-based OSD to achieve this improvement. Additionally, we introduce an efficient OSD algorithm tailored for quantum codes.
First, the sorting process plays a pivotal role in OSD [32]. For efficient implementation, we consider binary representations of an $n$-qubit quantum code and thus an $n$-qubit Pauli error variable can be represented by $2n$ binary variables. Previously we proposed a reliability metric that takes into account the hard-decision history obtained during BP4 iterations and incorporates soft messages from the last iteration [1]. The $2n$ binary error variables of an error vector are then sorted according to the reliability metric. Specifically, the length of the last run in a bit’s hard-decision history is the primary term in our metric. Through simulations, we demonstrate that incorporating the hard-decision history to determine the reliability order for OSD enhances decoding performance. Our OSD algorithm based on this reliability metric will be denoted as OSD4.
To develop an efficient OSD algorithm, it is crucial to reduce the problem size in OSD and minimize the computational burden of Gaussian elimination. To address this challenge, we propose an approach called reliable subset reduction. In the code capacity error decoding problem of an $[[n,k]]$ quantum code, we need to solve a linear system of equations with $2n$ binary variables and $n-k$ independent constraints. The core idea is that if the solution for a subset of the variables can be determined with confidence, the linear system can be reduced, thereby lowering the computation burden of both Gaussian elimination and the OSD process. So the primary task is to identify a highly reliable subset. An excellent choice are the bits that exhibit nearly constant hard decision history vectors during BP4 iterations and show high soft reliability in the final iteration. At low error rates, there is a large portion of highly reliable bits. Eliminating these bits would significantly accelerate the OSD process. Moreover, resources can be allocated to the less reliable parts, similar to the approaches in [42, 43]. We will demonstrate that in our simulations of surface codes with distances from 11 to 15, the problem sizes are reduced to less than 30% of the original in over 90% of the trials at an error rate of around 2%. We note that the ambiguity clustering decoder [56] and LSD [57] operate at a different stage from our approach, and these algorithms could be applied after reliable subset reduction.
Next, we consider approximate degenerate OSD in quantum codes. A typical OSD generates a list of error candidates for a given error syndrome and selects the one with the minimum weight, which is optimal for classical codes. For quantum codes, however, the optimal decoding criterion involves choosing the error coset with the highest probability, a criterion that is challenging to implement in practice. To address this, we consider approximate degenerate decoding in OSD by focusing only on relatively low-weight terms within an error coset.
Additionally, we introduce a degenerate error candidate pruning condition for stabilizer codes. We show that certain bit-flips in OSD correspond to multiplying stabilizers, resulting to degenerate error candidates. These bit-flips may alter the error weight but do not change the error coset, allowing them to be ignored in approximate degenerate OSD. For improving OSD efficiency, using a minimum weight decoding criterion is beneficial. A key observation is that if all bit-flips correspond to multiplying a stabilizer, there is no need to implement high-order OSD. We will demonstrate that this is indeed the case at low error rates by simulations in Section VI-G. We note that a degeneracy condition is also utilized in ambiguity clustering to facilitate the resolution of subproblems [56].
Integrating all the mentioned techniques, we propose an approximate degenerate OSD algorithm, abbreviated as ADOSD4. We conduct simulations of MBP4 and our OSD schemes OSD4 and ADOSD4 on several quantum codes under depolarizing errors in the code capacity noise model, including BB codes [55], rotated toric codes and rotated surface codes [24], (6.6.6) and (4.8.8) color codes [27], twisted XZZX codes on a torus [58], generalized hypergraph-product (GHP) codes [42, 43], and lift-connected surface (LCS) codes [59].
The results demonstrate that our proposed schemes outperform previous approaches in the literature, both in terms of error thresholds for topological codes and low error rate performance (logical error rate around $10^{-6}$). Both our MBP4+OSD4 and MBP4+ADOSD4 decoders outperform other binary BP-OSD schemes in these quantum codes. We achieve thresholds of approximately 17.5%–17.7% for the toric, surface, and twisted XZZX codes, and thresholds of 15%–15.42% for the (4.8.8) and (6.6.6) color codes, as summarized in Table II. Comparisons of our results with existing threshold values for the toric codes are provided in Table III. We also compare the performance of various reliability metrics on the GHP code, justifying our choice of reliability metric.
Finally, we analyze the time efficiency of our OSD schemes for various topological codes of comparable distances to the $[[882,48,16]]$ GHP code using single-threaded CPU implementations. The simulation results, summarized in Table IV, show that the time consumption of ADOSD4 is approximately 1% to 3% of that for order-2 OSD4 with similar decoding performance. Moreover, Table VI presents the time consumption of ADOSD4 in terms of BP iterations for surface codes with distances from 11 to 15. Specifically, the average time consumption per codeword for ADOSD4 is approximately equivalent to 2 to 3 BP4 iterations, whereas a successful BP trial typically takes 2 to 3 iterations on average. This suggests that ADOSD4 is an excellent complement to BP, especially since ADOSD4 is used only when BP fails.
This paper is organized as follows. We introduce quantum stabilizer codes in Section II. Section III discusses reliability metrics and the associated OSD algorithm. The reliable subset reduction method is presented in Section IV. We then cover approximate degenerate decoding in Section V, and provide simulation results in Section VI. Finally, we conclude in Sec. VII.
II Quantum stabilizer codes
In this section, we review the basic of stabilizer codes and define the relevant notation [4, 5, 6].
Let $I=\begin{bmatrix}1&0\\ 0&1\end{bmatrix},$ $X=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},$ $Z=\begin{bmatrix}1&0\\ 0&-11\end{bmatrix},$ and $Y=iXZ$ denote the Pauli matrices. Consider the $n$-fold Pauli group
$\mathcal{G}_{n}\triangleq\left\{cM_{1}\otimes\dots\otimes M_{n}:c\in\{\pm 1,\pm i\},M_{j}\in\{I,X,Y,Z\}\right\}.$
Every nonidentity Pauli operator in $\mathcal{G}_{n}$ has eigenvalues ±$\pm 1$. Any two Pauli operators either commute or anticommute with each other. The weight of a Pauli operator $E\in{\cal G}_{n}$, denoted $\mathrm{wt}\left(E\right)$, refers to the number of its nonidentity components.
We consider Pauli errors in this paper, assuming independent depolarizing errors with rate $\epsilon$. Each qubit independently experiences an $X$, $Y$, or $Z$ error with probability $\epsilon/3$ and no error with probability $1-\epsilon$. Therefore, at a small error rate $\epsilon$, low-weight Pauli errors are more likely to occur.
A stabilizer group $\mathcal{S}$ is an Abelian subgroup in $\mathcal{G}_{n}$ such that $-I^{\otimes n}\not\in\mathcal{S}$ [6]. Suppose that $\mathcal{S}$ is generated by $n-k$ independent generators. Then $\mathcal{S}$ defines an $[[n,k,d]]$ stabilizer code $\mathcal{C}(\mathcal{S})$ that encodes $k$ logical qubits into $n$ physical qubits:
$\mathcal{C}(\mathcal{S})=\left\{\ket{\psi}\in\mathbb{C}^{2^{n}}:S\ket{\psi}=\ket{\psi},~{}\forall\,S\in\mathcal{S}\right\}.$
The elements in $\mathcal{S}$ are called stabilizers. The parameter $d$ represents the minimum distance of the code such that any Pauli error of weight less than $d$ is detectable.
An error $E\in{\cal G}_{n}$ can be detected by ${\cal C}({\cal S})$ through stabilizer measurements if $E$ anticommutes with some of its stabilizers. Suppose that $\{S_{i}\}_{i=1}^{m}$, where $m\geq n-k$, is a set of stabilizers that generates ${\cal S}$. Their binary measurement outcomes are referred to as the error syndrome of $E$. Let $N({\cal S})\subset{\cal G}_{n}$ denote the normalizer group of ${\cal S}$, which consists of the Pauli operators that commute with all stabilizers. Consequently, if an error is in $N({\cal S})$, it will have zero syndrome and cannot be detected. Note that if an error is a stabilizer, it has no effect on the code space. An element in the normalizer group $N({\cal S})$ that is not a stabilizer, up to a phase, is called a nontrivial logical operator as it changes the logical state of a code . Therefore, the minimum distance of ${\cal C}({\cal S})$ is defined as the minimum weight of a nontrivial logical operator. It can be observed that $E$ and $ES$ for $S\in{\cal S}$ have the same effects on the codespace. They are called degenerate to each other. We say that a quantum code is highly degenerate if it has many stabilizers with low weight relative to its minimum distance.
A syndrome decoding problem can be stated as follows: given an error syndrome of an unknown Pauli error $E$, find the most probable $\hat{E}\in{\cal G}_{n}$ that matches the syndrome. An error estimate $\hat{E}$ is considered valid if it matches the syndrome or $\hat{E}E\in{\cal S}$. Two decoding criteria are typically considered: minimum weight decoding and degenerate maximum likelihood decoding [60, 61, 62]. In minimum weight decoding, the error $\hat{E}$ with the smallest weight is chosen, while in degenerate maximum likelihood decoding, the error $\hat{E}$ whose coset $\hat{E}\mathcal{S}$ has the minimum coset probability is selected. However, degenerate maximum likelihood decoding has exponential complexity and is impractical due to the need for coset enumeration. We propose the following criterion.
Definition 1.
(Approximate Degenerate Decoding)
A $\delta$-approximate degenerate decoding ($\delta$-ADD) criterion aims to find an error estimate $\hat{E}$ that matches the given syndrome, maximizing the dominant terms of its coset probability
$\sum_{S\in{\cal S},\mathrm{wt}\left(S\right)\leq\delta}\Pr\big{\{}\hat{E}S\big{\}}.$ (1)
Note that 00-ADD reduces to minimum weight decoding, while $n$-ADD corresponds precisely to degenerate maximum likelihood decoding.
II-A Binary representations
Without loss of generality, we assume that each stabilizer in $\{S_{i}\}_{i=1}^{m}$ is of the form $S_{i}=S_{i1}\otimes S_{i2}\otimes\dots\otimes S_{in}$, where $S_{ij}\in\{I,X,Y,Z\}$. Then we can study the decoding problem in the binary vector space [4, 6], using a mapping $\varphi:$
$I\mapsto\left[\begin{array}[]{c|c}0&0\end{array}\right],X\mapsto\left[\begin{array}[]{c|c}1&0\end{array}\right],Z\mapsto\left[\begin{array}[]{c|c}0&1\end{array}\right],Y\mapsto\left[\begin{array}[]{c|c}1&1\end{array}\right].$
A check matrix of the stabilizer code ${\cal C}({\cal S})$ is a binary matrix ${\mathbf{S}}=[{\mathbf{S}}_{ij}]\in\{0,1\}^{m\times 2n},$ where
$\left[\begin{array}[]{c|c}{\mathbf{S}}_{ij}&{\mathbf{S}}_{i(j+n)}\end{array}\right]=\varphi(S_{ij}).$
An $n$-fold Pauli operator $E\in{\cal G}_{n}$ can also be represented by $\varphi(E)\triangleq{\mathbf{E}}=\left[\begin{array}[]{c|c}{\mathbf{E}}^{X}&{\mathbf{E}}^{Z}\end{array}\right]\in\{0,1\}^{2n}$, where ${\mathbf{E}}^{X}$ and ${\mathbf{E}}^{Z}$ are the indicator vectors of $X$ and $Z$ components of $E$, respectively. Then the error syndrome of $E$ is
${\mathbf{z}}={\mathbf{S}}\Lambda{\mathbf{E}}^{\top}\in\{0,1\}^{m\times 1},$ (2)
where ${\mathbf{E}}^{\top}$ denotes the transpose of ${\mathbf{E}}$, $\Lambda=\left[\begin{array}[]{c|c}\mathbf{0}&{\mathbf{I}}_{n}\\ {\mathbf{I}}_{n}&\mathbf{0}\end{array}\right],$ ${\mathbf{I}}_{n}$ is the $n\times n$ identity matrix, and $\mathbf{0}$ is the zero matrix of appropriate dimensions.
Consequently, the decoding problem is to solve this system of linear equations with $2n$ binary variables ${\mathbf{E}}_{j}$ that are most probable. Note that ${\mathbf{S}}$ is of rank $n-k$ so the degree of freedom of this binary system is $n+k$.
II-B Belief Propagation (BP) Decoding
We explain how a quaternary BP algorithm works in the following [12], which is computed in linear domain. A log-likelihood version of BP can be found in [13].
Given a check matrix ${\mathbf{S}}$, a syndrome ${\mathbf{z}}$, and a depolarizing error rate $\epsilon$, BP performs the following steps:
  1. 1.
    An initial error distribution ${\mathbf{p}}_{i}\in\mathbb{R}^{4}$ for qubit $i\in\{1,2,\dots,n\}$ is chosen for BP.
  2. 2.
    The belief distribution vector ${\mathbf{q}}_{i}=(q_{i}^{I},q_{i}^{X},q_{i}^{Y},q_{i}^{Z})$ will be updated by BP, using the messages of parity-check satisfaction and the channel statistics ${\mathbf{p}}_{i}$ for $i=1,2,\ldots,n$.
  3. 3.
    At each iteration, an error estimate $\hat{{\mathbf{E}}}\in\{0,1\}^{2n}$ is made according to the hard decision on ${\mathbf{q}}_{i}$. If a valid error estimate is generated, it will be accepted as the solution; otherwise, go to the next iteration starting from step 2).
Steps 2) and 3) will be iterated for a maximum number of $T$ iterations, where $T$ is chosen in advance. If no valid estimate is obtained after $T$ iterations, BP will claim a failure.
Typically, the initial error distribution ${\mathbf{p}}_{i}$ is defined as $({1-\epsilon},\frac{\epsilon}{3},\frac{\epsilon}{3},\frac{\epsilon}{3})$. However, one may instead choose a fixed initialization by a given value $\epsilon_{0}\neq\epsilon$ for BP as long as BP performs well at this $\epsilon_{0}$. See more discussion on fixed initialization in [63, 13, 31].
III Order-statistic decoding based on quaternary belief propagation
In the aforementioned syndrome decoding problem, we aim to solve a system of linear equations given by (2), for a given parity-check matrix ${\mathbf{S}}$ and an error syndrome ${\mathbf{z}}$ with $2n$ binary variables. The linear system has $n-k$ independent constraints, impliying that there are $n+k$ degrees of freedom in the vector ${\mathbf{E}}$. Given the values of these $n+k$ bits, one can reconstruct ${\mathbf{E}}$. If BP fails to provide a valid error estimate after $T$ iterations, OSD will be utilized.
In the context of OSD, a critical step involves identifying and sorting the accurate and reliable coordinates to establish $n+k$ linearly independent bits. Then, we can use these coordinates to generate a valid error that matches the syndrome. To accomplish this, it is necessary to establish the concept of reliability order as a means of prioritizing the coordinates based on their reliability.
III-A Reliability order based on BP4
To utilize the quaternary distribution generated by BP4 for each qubit error in OSD, we propose a method that employs the hard-decision history from all BP iterations and the output probabilities from the last BP iteration.
Suppose that BP fails to provide a valid error after a maximum of $T$ iterations. At this point, we have the output distribution ${\mathbf{q}}_{i}=(q_{i}^{I},q_{i}^{X},q_{i}^{Y},q_{i}^{Z})$, $i=1,2,\dots,n$ from the last BP iteration. Additionally, we have the hard-decision results for all $T$ iterations, denoted by ${\mathbf{h}}_{i}^{(j)}\in\{I,X,Y,Z\}$ for $i=1,\dots,n$ and $j=1,\dots,T$. For example, ${\mathbf{h}}_{i}^{(1)},{\mathbf{h}}_{i}^{(2)},\dots,{\mathbf{h}}_{i}^{(T)}$ represent the hard-decision result history of BP at qubit $i$ and ${\mathbf{h}}_{1}^{(j)},{\mathbf{h}}_{2}^{(j)},\dots,{\mathbf{h}}_{n}^{(j)}$ represent the hard-decision results of BP at the $j$-th iteration.
Our objective is to assign reliability for each Pauli $X$ and $Z$ error. Following the reliability definitions in our previous paper [1], we define two types of measures in the following.
Definition 2.
Define a hard-decision reliability vector $\bm{\eta}=(\bm{\eta}_{1},\dots,\bm{\eta}_{n})\in\mathbb{Z}^{n}$ such that $\bm{\eta}_{i}$ represents the number of consecutive iterations during which the error at qubit $i$ remains unchanged until the final hard decision is made.
In other words, $\bm{\eta}_{i}$ is the length of the last run in the string ${\mathbf{h}}_{i}^{(1)},{\mathbf{h}}_{i}^{(2)},\dots,{\mathbf{h}}_{i}^{(T)}$. For example, if the hard-decision outputs at the first qubit over the last five iterations are $X,Y,Z,Z,Z$, then the final hard decision output is $Z$ and $\bm{\eta}_{1}=3$. This means that the error at qubit 1 persisted unchanged as $Z$ for three iterations when BP stops.
For implementation, storing the entire hard-decision history ${\mathbf{h}}$ is unnecessary. One can record only the hard-decision outputs at the latest iteration and uses the latest belief vector ${\mathbf{q}}_{i}$ to update the values of $\{\bm{\eta}_{i}:i=1,\dots,n\}$.
Our modified BP4 algorithm, which incorporates hard-decision reliability measures, is presented in Algorithm 1.
Algorithm 1 BP4
Input: check matrix ${\mathbf{S}}\in\{0,1\}^{m\times 2n}$, syndrome ${\mathbf{z}}\in\{0,1\}^{m\times 1}$, depolarizing error rate $\epsilon$, maximum number of iterations $T$.
Output: A valid error estimate $\hat{{\mathbf{E}}}\in\{0,1\}^{2n}$ and belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$.
Initialization:
Choose initial error distributions ${\mathbf{p}}_{i}\in\mathbb{R}^{4}$ for $i=1,\dots,n$. Let $\bm{\eta}_{i}=1$ for $i=1,\dots,n$. Let ${\mathbf{h}}_{i}=I$ for $i=1,\dots,n$. Let ${\mathbf{q}}_{i}=\mathbf{0}$ for $i=1,\dots,n$.
Steps:
for $j=1$ to $T$ do
     Update $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, using ${\mathbf{S}},{\mathbf{z}},$ and $\{{\mathbf{p}}_{i}\}_{i=1}^{n}$.
     for $i=1$ to $n$ do
         if  ${\mathbf{h}}_{i}=\operatorname{HardDecision}({\mathbf{q}}_{i})$ then $\bm{\eta}_{i}\leftarrow\bm{\eta}_{i}+1$;
         else ${\mathbf{h}}_{i}\leftarrow\operatorname{HardDecision}({\mathbf{q}}_{i})$. $\bm{\eta}_{i}=1$.
         end if
     end for
     if  the hard-decision ${{\mathbf{h}}}$ matches the syndrome ${\mathbf{z}}$ then
         return $\varphi({\mathbf{h}})$ and $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$.
     end if
end for
return $\mathbf{0}$ and $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$. $\triangleright$ BP fails.
When BP fails, we are left with belief distributions ${\mathbf{q}}_{i}=(q_{i}^{I},q_{i}^{X},q_{i}^{Y},q_{i}^{Z})$ for $i=1,2,\dots,n$, which will be utilized in the reliability measure.
Definition 3.
Define two soft reliability functions
$\phi^{X}(i)$ $=\max\{q_{i}^{X}+q_{i}^{Y},q_{i}^{I}+q_{i}^{Z}\},$
$\phi^{Z}(i)$ $=\max\{q_{i}^{Z}+q_{i}^{Y},q_{i}^{I}+q_{i}^{X}\},$
for $i\in\{1,2,\dots,n\}$.
Note that $\{q_{i}^{X}+q_{i}^{Y},q_{i}^{I}+q_{i}^{Z}\}$ is a binary distribution indicating the likelihood of an $X$ error occuring at the $i$-th qubit and similarly for $\{q_{i}^{X}+q_{i}^{Y},q_{i}^{I}+q_{i}^{Z}\}$.
Now we define a reliability order for the bits in ${\mathbf{E}}=\left[\begin{array}[]{c|c}{\mathbf{E}}^{X}&{\mathbf{E}}^{Z}\end{array}\right]\in\{0,1\}^{2n}$ based on the outputs of BP4, using the aforementioned hard and soft reliability functions.
Definition 4.
For $1\leq i,j\leq n$ and $a,b\in\{X,Z\}$, error bit ${\mathbf{E}}_{i}^{a}$ is said to be more reliable than error bit ${\mathbf{E}}_{j}^{b}$ if $\bm{\eta}_{i}>\bm{\eta}_{j}$, or if $\phi^{a}(i)\geq\phi^{b}(j)$ when $\bm{\eta}_{i}=\bm{\eta}_{j}$.
This definition means that the hard-decision history reliability is used to rank the error coordinates, with the soft reliability serving as a tie-breaker.
Let ${\mathbf{E}}$ be the hard-decision output from BP4 after $T$ iterations. If ${\mathbf{E}}$ does not match the syndrome, OSD will be applied. We now present our OSD algorithm based on the reliability orders as in Algorithm 2, which is also referred to as OSD4-00. If the bits in the reliable subset are correct, the output of OSD4-00 is the required solution.
Algorithm 2 OSD4-0
Input: check matrix ${\mathbf{S}}\in\{0,1\}^{m\times 2n}$ of rank $n-k$, syndrome ${\mathbf{z}}\in\{0,1\}^{m\times 1}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, hard-decision vector ${\mathbf{E}}\in\{0,1\}^{2n}$, and belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$.
Output: A valid error estimate $\hat{{\mathbf{E}}}\in\{0,1\}^{2n}$.
Steps:
1:Sort the error bits in ascending order of reliability. Construct a column permutation function $\pi$ corresponding to the sorting order. Calculate $\pi({\mathbf{E}})$ and $\pi({\mathbf{S}}\Lambda)$.
2:Perform Gaussian elimination on $\pi({\mathbf{S}}\Lambda)$ and let the output be denoted by $R(\pi({\mathbf{S}}\Lambda))$, where $R$ is the corresponding row operations in the Gaussian elimination. If necessary, apply a column permutation function $\mu$ to ensure that the first $n-k$ columns of the resulting matrix are linearly independent. Thus we have
$\mu(R(\pi({\mathbf{S}}\Lambda)))=\left[\begin{array}[]{cc}{\mathbf{I}}_{n-k}&{\mathbf{A}}\\ \mathbf{0}&\mathbf{0}\end{array}\right],$ (5)
for some ${\mathbf{A}}\in\{0,1\}^{(n-k)\times(n+k)}$, where the last $m-(n-k)$ rows are all zeros.
3:Let ${\mathbf{z}}^{\prime}=R({\mathbf{z}})$ so that the parity-check condition (2) becomes ${\mathbf{z}}^{\prime}=\mu(R(\pi({\mathbf{S}}\Lambda)))(\mu(\pi(\hat{{\mathbf{E}}})))^{\top}$. Remove the last $m-(n-k)$ entries of ${\mathbf{z}}^{\prime}$.
4:Suppose $\mu(\pi({\mathbf{E}}))=\left[{\mathbf{E}}^{U}~{}~{}{\mathbf{E}}^{R}\right]$, where ${\mathbf{E}}^{U}$ represents the independent and unreliable subset of $n-k$ bits and ${\mathbf{E}}^{R}$ is the reliable subset of $n+k$ bits. return
$\hat{{\mathbf{E}}}\triangleq\pi^{-1}\left(\mu^{-1}\left(\left[{{\mathbf{z}}^{\prime}}^{\top}\oplus{{\mathbf{E}}^{R}}{\mathbf{A}}^{\top}\qquad{\mathbf{E}}^{R}\right]\right)\right).$ (6)
For comparison with conventional OSD procedures that do not incorporate the history of BP hard-decision outputs across iterations, we also define a reliability order that accounts for only the soft reliability functions of BP4.
Definition 5.
Error bit ${\mathbf{E}}_{i}^{a}$ is said to be more reliable than error bit ${\mathbf{E}}_{j}^{b}$ in the marginal distribution if $\phi^{a}(i)\geq\phi^{b}(j)$.
An OSD algorithm that follows a reliability order depending only on the belief distributions from the last BP iteration is referred to as mOSD. The OSD based on Definition 5 is referred to as mOSD4.
Note that the reliability metric in Definition 5 is distinct from those in [42, 43]. Other alternative metrics can also be considered, such as
[…]
The reliability of qubit $i$ is assessed based on the quaternary entropy of its belief distribution $\mathbf{q}_{i}$ (Entropy) or based on the maximum probability of the distribution (Max). For our purpose, the hard-decision output of BP is first sorted in the Pauli basis based on the reliability metric (either Entropy or Max) and then transformed into a binary vector for OSD [42]. We will compare these reliability metrics in Section VI-D.
III-B OSD4-$w$
If some bits in the reliable subset are incorrect, additional processing is necessary. We may flip up to $w$ bits in the reliable subset and generate the corresponding error estimate. If this error estimate is valid, it is an error candidate. This process is repeated for all $\sum_{i=0}^{w}\binom{n+k}{i}$ possibilities, and the output will be a valid error candidate that is optimal with respect to $\delta$-ADD for a certain $\delta$. This type of OSD algorithm is referred to as order-$w$ OSD4 (or OSD4-$w$ for short).
In order to explore all $\sum_{i=0}^{w}\binom{n+k}{i}$ possibilities, we propose to use the depth-first search (DFS) algorithm. Figure 1 illustrates the idea of using DFS to traverse all bit strings of length 4 and with weight at most 2. One can see that a child node is obtained by flipping one zero bit from its parent node. Consequently, one can recursively generate all $\sum_{i=0}^{w}\binom{n+k}{i}$ possibilities by bit flipping and encoding.
Lemma 6.
Given an OSD4-0 output vector, flipping one of its reliable bits to generate an error candidate takes time ${\cal O}(n)$. Similarly, generating an error candidate corresponding to a child node from its parent node also takes ${\cal O}(n)$ time.
Proof.
Consider the vector $\Big{[}{\mathbf{z}}^{\prime\top}\oplus{\mathbf{E}}^{R}{\mathbf{A}}^{\top}\,\Big{|}{\mathbf{E}}_{1}^{R}\,{\mathbf{E}}_{2}^{R}\,\cdots\,{\mathbf{E}}_{n+k}^{R}\Big{]}$ generated by $\operatorname{OSD}_{4}$-0 in (6), where ${\mathbf{E}}_{j}^{R}$ are reliable bits. Suppose that the matrix ${\mathbf{A}}$ in (5) has columns ${\mathbf{A}}_{1},{\mathbf{A}}_{2},\dots,{\mathbf{A}}_{n+k}\in\{0,1\}^{n-k}$. Then flipping the bit ${\mathbf{E}}_{i}^{R}$ generates the vector
$\left[{\mathbf{z}}^{\prime\top}\oplus{\mathbf{E}}^{R}{\mathbf{A}}^{\top}\oplus{\mathbf{A}}_{i}^{\top}\middle|{\mathbf{E}}_{1}^{R}\,{\mathbf{E}}_{2}^{R}\,\cdots\Big{(}1\oplus{\mathbf{E}}_{i}^{R}\Big{)}\cdots\,{\mathbf{E}}_{n+k}^{R}\right],$ (9)
which is a valid error candidate after permutations. It can be checked that this calculation takes time $\mathcal{O}(n)$. ∎
To efficiently navigate through these possibilities, we can utilize the technique described in Lemma 6 combining with the DFS algorithm. Then all $\sum_{i=0}^{w}\binom{n+k}{i}$ candidates can be recursively generated.
We prioritize flipping the least reliable bits first. If an error with a lower Hamming weight is encountered, we update the solution accordingly. However, if the error has the same Hamming weight as the current one, we retain the previous solution since the bit flipped in the previous step is deemed less reliable.
00001000110010101001010001100101001000110001
Figure 1: A tree with nodes consisting of all bit strings of length 4 and with at most weight 2. The root node is $0$. A child node is obtained by flipping one of the zero bits from its parent node that come after the last bit that is one.
III-C Complexity of OSD4
We analyze the complexity of OSD based on 00-ADD in the following. The complexity of OSD4-0 is dominated by the Gaussian elimination step, which is then ${\cal O}(2n\times m^{2})={\cal O}(n^{3})$, assuming $m={\cal O}(n)$. For OSD4-$w$ with $w>0$, the total number of possible flips is $\sum_{j=0}^{w}\binom{n+k}{j}$, which can be expressed as ${\cal O}((n+k)^{w})$ or simply ${\cal O}(n^{w})$.
In Lemma 6, when we flip a bit in the reliable part, it requires only $\mathcal{O}(n)$ calculations to find an error candidate. Additionally, by running through $\sum_{j=0}^{w}\binom{n+k}{j}$ possibilities using the DFS algorithm, the OSD4-$w$ algorithm has a complexity of $\mathcal{O}(n^{3}+n^{w+1})$, which simplifies to $\mathcal{O}(n^{3})$ when $w\leq 2$.
The above OSD complexity is only required when BP fails to converge. BP has a worst-case time complexity of $O(n\gamma T)$ using a sequential schedule, where $n$ is the number of error variables, $\gamma$ is the mean column weight of the check matrix, and $T$ is the maximum number of iterations. However, BP typically converges quickly, within $O(\log N)$ iterations, for sparse matrices. This is also the case for MBP4 as demonstrated in Table VI in the simulation section. Thus MBP4 has a nearly linear complexity in $n$ from simulations and has a high probability of convergence for many spare-graph quantum codes [12, 13].
Algorithm 3 HRSR
Input: check matrix ${\mathbf{S}}\in\{0,1\}^{m\times 2n}$ of rank $n-k$, syndrome ${\mathbf{z}}\in\{0,1\}^{m\times 1}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, hard-decision vector ${\mathbf{E}}\in\{0,1\}^{2n}$, belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, and soft reliability threshold $\theta$.
Output: shortened variables $\tilde{{\mathbf{z}}}$, $\tilde{{\mathbf{S}}}$, ${\mathbf{E}}^{H}$, and $\sigma$.
Steps:
1:Use $\bm{\eta}$, ${{\mathbf{q}}_{i}}_{i=1}^{n}$, and $\theta$ to identify the highly reliable bits in ${\mathbf{E}}$ based on Definition 7. Assume that there are $v$ such highly reliable bits.
2:Let $\sigma$ be a column permutation such that $\sigma({\mathbf{E}})=\begin{bmatrix}{\mathbf{E}}^{\prime}&{\mathbf{E}}^{H}\end{bmatrix}$, where ${\mathbf{E}}^{H}$ is the highly reliable subset of $v$ bits, and ${\mathbf{E}}^{\prime}$ is the remaining $2n-v$ bits. Let $\sigma(\hat{{\mathbf{E}}})=\begin{bmatrix}\hat{{\mathbf{E}}}^{\prime}&{\mathbf{E}}^{H}\end{bmatrix}$ be a permuted error estimate, which contains the highly reliable bits in ${\mathbf{E}}^{H}$.
3:Let $\tau$ be a row permutation such that $\tau(\sigma({\mathbf{S}}\Lambda))=\begin{bmatrix}\tilde{{\mathbf{S}}}&{\mathbf{B}}\\ \mathbf{0}&{\mathbf{C}}\\ \end{bmatrix}$ for $\tilde{{\mathbf{S}}}\in\{0,1\}^{m^{\prime}\times(2n-v)}$ with $m^{\prime}\leq m$. The $m-m^{\prime}$ rows $\begin{bmatrix}\mathbf{0}&{\mathbf{C}}\end{bmatrix}$ correspond to parity checks related only to the highly reliable bits. Assume that $\tau({\mathbf{z}})=\begin{bmatrix}{\mathbf{z}}^{\prime\prime}\\ {\mathbf{z}}^{H}\end{bmatrix}$, where ${\mathbf{z}}^{\prime\prime}\in\{0,1\}^{m^{\prime}\times 1}$, and ${\mathbf{z}}^{H}\in\{0,1\}^{(m-m^{\prime})\times 1}$ are the syndrome bits corresponding to $\begin{bmatrix}\mathbf{0}&{\mathbf{C}}\end{bmatrix}$.
4:Verify the parity check condition related to the highly reliable bits.
if ${\mathbf{z}}^{H}\neq{\mathbf{C}}\big{(}{\mathbf{E}}^{H}\big{)}^{\top}$ then return failure;
else we have a new parity-check condition
$\tilde{{\mathbf{z}}}=$ $\tilde{{\mathbf{S}}}\big{(}\hat{{\mathbf{E}}}^{\prime}\big{)}^{\top},$ (10)
     where $\tilde{{\mathbf{z}}}={\mathbf{z}}^{\prime\prime}-{\mathbf{B}}\big{(}{\mathbf{E}}^{H}\big{)}^{\top}$.
     if (10) is solvable then return $\tilde{{\mathbf{z}}}$, $\tilde{{\mathbf{S}}}$, ${\mathbf{E}}^{H}$, and $\sigma$;
     else return failure.
     end if
end if
IV Efficient OSD with highly reliable subset reduction
As we increase the order $w$ of OSD, we can expect an improvement in the decoder’s performance. In particular, when $w=n+k$ and $\delta=n$, OSD4-$(n+k)$ using $n$-ADD becomes a maximum likelihood decoder. However, increasing the order $w$ also leads to an increase in time complexity. To address this issue, we propose an approach to reduce the problem size for OSD. Recall that OSD is used to solve a linear system with $2n$ binary variables and $n-k$ constraints. If a subset of binary variables is known, the problem size–and consequently the computational complexity–can be reduced. The number of binary variables in the reduced linear system is referred to as the effective length for OSD, which corresponds to the length of the remaining binary error variable vector.
We note that OSD is applied when BP fails to find a valid error that matches the syndrome. During the $T$ iterations of message passing in BP, some coordinates become relatively more reliable than others. We have the following definition of highly reliable coordinates.
Definition 7.
An error bit ${\mathbf{E}}_{i}^{a}$, where $i\in{1,\dots,n}$ and $a\in{X,Z}$, is considered highly reliable if its hard-decision reliability equals $T$ or $T+1$, and its soft reliability $\phi^{a}(i)$ satisfies $\phi^{a}(i)\geq\theta$, where $\theta$ is a predetermined reliability threshold.
An error variable has hard reliability $T$ if BP has a constant output at all iterations; a hard reliability of $T+1$ means the corresponding error variable is the identity during the BP process after being initialized to the identity.
The soft reliability threshold $\theta$ is chosen close to one to ensure that a highly reliable coordinate not only exhibits an almost constant hard-decision history but also has a high soft reliability value.
These highly reliable variables are crucial in the OSD process. Identifying these variable nodes allows us to greatly reduce the computation burden of OSD.
In topological codes, such as toric or surface codes, an isolated Pauli error refers to a single-qubit Pauli error (including identity) with no neighboring nonidentity Pauli error within a radius of two qubits. These isolated Pauli errors can be decoded using BP4 with high confidence, contributing to the highly reliable parts in the output of BP.
These highly reliable parts provide outcomes that can be trusted and will remain unchanged throughout the OSD4 procedure with high probability. Consequently, we propose to exclude these highly reliable parts from the parity check relations. By doing so, we can avoid unnecessary calculations and prioritize our focus on the unreliable part. Our highly reliable subset reduction (HRSR) algorithm is given in Algorithm 3.
Equation (10) follows from the parity-check condition (2) as
$\begin{bmatrix}{\mathbf{z}}^{\prime\prime}\\ {\mathbf{z}}^{H}\end{bmatrix}=$ $\tau(\sigma({\mathbf{S}}\Lambda))\begin{bmatrix}\hat{{\mathbf{E}}}^{\prime}&{\mathbf{E}}^{H}\end{bmatrix}^{\top}$
$=$ $\begin{bmatrix}\tilde{{\mathbf{S}}}\big{(}\hat{{\mathbf{E}}}^{\prime}\big{)}^{\top}+{\mathbf{B}}\big{(}{\mathbf{E}}^{H}\big{)}^{\top}&{\mathbf{C}}\big{(}{\mathbf{E}}^{H}\big{)}^{\top}\end{bmatrix}$
and using the conditions ${\mathbf{z}}^{H}={\mathbf{C}}\big{(}{\mathbf{E}}^{H}\big{)}^{\top}$ and $\tilde{{\mathbf{z}}}={\mathbf{z}}^{\prime\prime}-{\mathbf{B}}\big{(}{\mathbf{E}}^{H}\big{)}^{\top}$.
If HRSR returns a reduced linear system as in (10), we then use OSD4-$w$ to solve it with hard-decision BP output ${\mathbf{E}}^{\prime}$ and corresponding reliabilities $\bm{\eta}$, $\phi^{X}$, and $\phi^{Z}$. Suppose that OSD4-$w$ returns $\hat{{\mathbf{E}}}^{\prime}$. Then the error estimate is $\sigma^{-1}\left(\begin{bmatrix}\hat{{\mathbf{E}}}^{\prime}&{\mathbf{E}}^{H}\end{bmatrix}\right)$.
Reducing the problem size may cause the reduced check matrix to lose the ability to represent all valid codewords if its dimension becomes too small. To address this, we set the soft reliability threshold $\theta$ very close to one to ensure that highly reliable bits are accurate. Furthermore, in Step 4 of Algorithm 3, we verify that the removed highly reliable bits align with the corresponding syndrome bits. If they do not, the algorithm reports a failure.
This method effectively reduces the overhead in OSD, especially at low error rates as will be shown in the simulation section. Furthermore, as the effective length in OSD is reduced, we can apply higher order OSD if only a fixed amount of computation is allowed. Consider, for instance, the case where the number of error candidates to be tested in OSD is $\Gamma$. For the original decoding problem, OSD4-$2$ has
$\Gamma=\binom{n+k}{0}+\binom{n+k}{1}+\binom{n+k}{2}.$ (11)
Using the HRSR algorithm for $v$ highly reliable bits, we have a reduced system of linear equations with $2n-v$ binary variables. Assume the rank of the reduced system is $n-k-(m-m^{\prime})$. Then we can apply OSD4-$w$ under the constraint of $\Gamma$ candidates, with $w$ given by
$w=\operatorname*{arg\,max}_{x}\sum_{i=0}^{x}\binom{n+k-v+(m-m^{\prime})}{i}\leq\Gamma.$ (12)
Note that even though the number of tested candidates is fixed, the complexity of OSD following the HRSR algorithm is lower due to the reduced effective length.
V Approximate degenerate OSD
In this section, we discuss how degeneracy in quantum codes helps reduce the computation of OSD. Recall that in OSD-$w$, $\sum_{i=0}^{w}\binom{n+k}{i}$ candidates are generated, and the one that is optimal with respect to $\delta$-ADD is chosen as the output. As shown in Lemma 6, a new candidate can be generated from an error candidate by flipping a bit in the reliable subset and adding a certain vector to the unreliable subset. This new candidate must belong to the same error coset, and thus the bit flipping in the reliable subset corresponds to an operator in the normalizer group. We have the following degenerate error candidate pruning conditions.
Lemma 8.
Consider an $[[n,k,d]]$ stabilizer code with a check matrix given by $\begin{bmatrix}{\mathbf{I}}_{n-k}&{\mathbf{A}}\end{bmatrix}$ after Gaussian elimination and a column permutation as in Equation (5), where ${\mathbf{A}}\in\{0,1\}^{(n-k)\times(n+k)}$ are associated with the reliable bits of OSD4-0 output.
1) When using $n$-ADD, only the reliable bits corresponding to nontrivial logical operators need to be considered in OSD4-$w$.
2) If each error corresponding to flipping a reliable bit of OSD4-0 output is a stabilizer, then OSD4-$w$ is equivalent to OSD4-0 for any $w$.
Proof.
1)
In optimal degenerate decoding ($n$-ADD), different stabilizer cosets are compared. Therefore, multiplying a stabilizer by an error candidate does not create a new nontrivial candidate.
2) The statement is straightforward, as flipping these reliable bits results in only degenerate errors to the OSD4-0 output.
Exploiting Lemma 8-1) with $n$-ADD makes OSD4-$w$ more effective. By incorporating with the HRSR algorithm, we can work on a shortened check matrix $\tilde{{\mathbf{S}}}$ and higher order of OSD can be executed. Consequently, (12) can be improved to $w=\operatorname*{arg\,max}_{x}\sum_{i=0}^{x}\binom{u}{i}\leq\Gamma,$ where $u\leq n+k-v+(m-m^{\prime})$ is the number of columns in the shortened check matrix corresponding to nontrivial logical operators.
For checking whether flipping a bit is nontrivial can be done by checking whether its induced error anticommutes with the permuted logical operators. If an error commutes with all logical operators, it belongs to the stabilizer group. This aligns with the standard procedure in QEC simulations for calculating the logical error rate. We note that the computational complexity of verifying such a bit-flipping event is $O((n-k)k)$. Consequently, checking all the columns of $A$ has a computational complexity of roughly the same order as Gaussian elimination. This approach is both effective and relatively efficient for implementing high-order OSD.
However, the complexity of $n$-ADD is not practically feasible. We propose to use $\delta$-ADD by choosing $\delta$ as the highest weight of a set of low-weight stabilizer generators. For the case of a rotated surface code [24], the stabilizer generators are of weight $2$ or $4$, and the next lowest stabilizer weight is $6$. Thus, $n$-ADD can be well approximated by $4$-ADD at error rates lower than $1\%$. Assume that OSD4444-0 generates a list of error candidates $\mathcal{E}$. Then, $4$-ADD outputs:
$\hat{{\mathbf{E}}}=\operatorname*{arg\,max}_{{\mathbf{E}}}\sum_{i=0}^{n-k}\Pr\big{\{}{\mathbf{E}}+{\mathbf{S}}_{i}\big{\}},$ (13)
where $\mathbf{S}_{0}=I$ and $\{\mathbf{S}_{i}\}_{i=1}^{n-k}$ is a set of stabilizer generators for the rotated surface code. For practical implementation, let $W(x)$ be the (Pauli) weight enumerator of the set $\{\mathbf{E}+\mathbf{S}_{i}:i=0,\dots,n-k\}$, and then the calculation of (13) can be approximated by using the dominating terms in $W(\epsilon/(1-\epsilon))$ for error rate $\epsilon<1\%$.
For codes with algebraic structures, the first few lowest-weight terms can be derived, but determining the complete weight enumerator for a general code is an NP-hard problem. As an alternative, we propose selecting a set of $m\geq n-k$ low-weight stabilizer generators and calculating the weight distribution of (${m\choose c}$ stabilizers, where $c$ is a constant independent of $n$. This method provides a good approximation of the dominant terms for practical purposes. To maintain an overall complexity of $O(n^{3})$, the value of (${m\choose c}$ should be chosen accordingly.
To achieve optimal execution complexity, 00-ADD is often used. Lemma 8 suggests focusing on the columns of $A$ corresponding to nontrivial logical operators when $n$-ADD is used. However, in the scenario of 00-ADD, a stabilizer may influence the output of $\text{OSD}_{4}$-$w$ since multiplying a stabilizer can reduce the weight of the error, potentially providing a better candidate in 00-ADD, even though the error coset remains unchanged. Thus, $\text{OSD}_{4}$-$w$ cannot simply flip bits corresponding to those nontrivial logical operators as in the case of Lemma 8-1).
The following provides a sufficient condition to determine whether flipping a reliable bit corresponds to a stabilizer.
Lemma 9.
Consider an $[[n,k,d]]$ stabilizer code with a check matrix given by $\begin{bmatrix}{\mathbf{I}}_{n-k}&{\mathbf{A}}\end{bmatrix}$ after Gaussian elimination and a column permutation as in Equation (5), where ${\mathbf{A}}\in\{0,1\}^{(n-k)\times(n+k)}$ are associated with the reliable bits of OSD4-0 output. If the weight of the $j$-th column of ${\mathbf{A}}$ is less than $d-1$, then flipping the $j$-th reliable bit of an error candidate corresponds to multiplying the error candidate by a stabilizer.
Proof.
Suppose that ${\mathbf{E}}=\begin{bmatrix}{\mathbf{F}}\ \Big{|}&{\mathbf{E}}_{1}^{R}&{\mathbf{E}}_{2}^{R}&\cdots&{\mathbf{E}}_{n+k}^{R}\end{bmatrix}$ is a valid error candidate, where ${\mathbf{E}}_{i}\in\{0,1\}$ for $i=1,\dots,n$ and ${\mathbf{F}}\in\{0,1\}^{n-k}$. By (9), flipping the bit ${\mathbf{E}}_{j}^{T}$ corresponds to adding ${\mathbf{E}}$ by the vector $\begin{bmatrix}{\mathbf{A}}_{j}^{\top}&0\cdots 0&1&0\cdots 0\end{bmatrix}\in\{0,1\}^{2n}$, where ${\mathbf{A}}_{j}$ is the $j$-th column of ${\mathbf{A}}$. This vector corresponds to a Pauli operator of weight at most $d-1$ and, therefore, is a stabilizer since the minimum distance of the code is $d$.
By Lemma 8-2) and Corollary 9, we establish the following operational degenerate error candidate pruning condition.
Corollary 10.
Let $\tilde{S}\in\{0,1\}^{m^{\prime}\times(2n-v)}$ be a shortened check matrix by the HRSR algorithm. Suppose that $\tilde{S}$ can be transformed into $\begin{bmatrix}{\mathbf{I}}&\tilde{A}\end{bmatrix}$ after Gaussian elimination and a column permutation, where $\tilde{A}$ is associated with the reliable bits. If all the columns of $\tilde{{\mathbf{A}}}$ are of weight less than $d-1$, then OSD4-$w$ on the corresponding reduced linear system is equivalent to OSD4-0 for any $w$.
The operational degenerate error pruning condition in Corollary 10 provides a more operationally efficient method to determine whether order-$w$ OSD for $w>0$ is nontrivial by evaluating the weight of each column. This can be done in $O((n-k)\times(n+k))$.
If there are many highly reliable bits, the number of stabilizers associated with these bits would be high. Consequently, the shortened check matrix $\tilde{\mathbf{S}}$ from the HRSR algorithm will have fewer rows and lower column weights. Therefore, Corollary 10 can be applied with high probability when the error rate is low.
We combine all the aforementioned techniques into the approximate degenerate OSD (ADOSD) algorithm, as detailed in Algorithm 4, using 00-ADD. It is important to note that if decoding performance is of higher priority, $\delta$-ADD can be utilized, and Lemma 9 can be integrated into Algorithm 4.
Algorithm 4 ADOSD4
Input: check matrix ${\mathbf{S}}\in\{0,1\}^{m\times 2n}$ of rank $n-k$, syndrome ${\mathbf{z}}\in\{0,1\}^{m\times 1}$, BP output distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, soft reliability threshold $\theta$, hard-decision vector ${\mathbf{E}}\in\{0,1\}^{2n}$, maximum number of OSD flips $\Gamma$, and backup OSD order $w$.
Output: A valid error estimate $\hat{{\mathbf{E}}}\in\{0,1\}^{2n}$.
Steps:
Apply the HRSR algorithm.
if  HRSR returns failure, then
     use OSD4-$w$ to find an error estimate $\hat{{\mathbf{E}}}\in\{0,1\}^{2n}$ for the original problem, and return $\hat{{\mathbf{E}}}$;
else HRSR returns ${\mathbf{z}}^{\prime}$, $\tilde{{\mathbf{S}}}$, ${\mathbf{E}}^{H}$, and $\sigma$.
     if all the columns of $\tilde{{\mathbf{S}}}$ in reduced row echelon form have weight less than $d-1$then $w\leftarrow 0$;
     else determine $w$ by (12).
     end if
     Use OSD4-$w$ on ${\mathbf{z}}^{\prime}$, $\tilde{S}$, and the given data to find an error estimate $\hat{{\mathbf{E}}}^{\prime}\in\{0,1\}^{2n-v}$, and return $\sigma^{-1}\left(\begin{bmatrix}\hat{{\mathbf{E}}}^{\prime}&{\mathbf{E}}^{H}\end{bmatrix}\right)$.
end if
Remark 11.
We can incorporate the idea of checking the commutation relations with permuted logical operators for column-induced errors of weight at least $d$ into Algorithm 4.
For instance, we first identify columns with weights less than $d-1$ and then verify the commutation relations between the permuted logical operators and the errors induced by the remaining columns.
As will be shown in Table VII, using the strategy outlined in Corollary 10, OSD-00 is applied in the majority of trials when the error rate is low.
decoder source note
MBP4+ADOSD4 this paper based on Algorithm 4
MBP4+OSD4-$w$ this paper order-$w$ OSD4 with reliability metric from Definition 4
MBP4+mOSD4-$w$ this paper order-$w$ OSD4 with reliability metric from Definition 5
AMBP4 [31] No postprocessing. Can act as a predecoder for ADOSD4 and OSD4
BP+mOSD-($w$,$\lambda$) [42] binary BP+ order-w OSD4 with additional flips on $\lambda$ positions
BP-OSD-CS($\lambda$) [43] binary BP-OSD-0 with a greed search on (${\lambda\choose 2}$ positions
BP+GDG [54] binary BP with guided decimal guessing based on a decision tree
TABLE I: Various decoders compared in this paper.
VI Simulation results
We simulate the performance of various BP+OSD schemes on LCS codes [59], GHP codes [42, 43], BB codes [55], and various 2D topological codes, including toric, rotated surface, color codes and twisted XZZX codes [22, 23, 24, 25, 26, 58, 28] (see [64, Table I]).
The decoding performance of a quantum code ${\cal C}({\cal S})$ and its decoder is assessed through the logical error rate (LER). An error $E\in{\cal G}_{n}$ is classified as a logical error if it is a nontrivial logical operator.
The simulation parameters are as follows: The maximum number of BP iterations, $T$, is set to $100$ unless otherwise specified. In most cases, a much smaller value of $T$ suffices for good decoding performance due to the rapid convergence of MBP4[31], which typically completes in just a few iterations for codes with lengths under one thousand. For instance, as shown in Table V, When MBP4 successfully decodes the $[[225,1,15]]$ rotated surface code, it requires an average of only $2.41$ iterations at a depolarizing rate of $\epsilon=3.3\%$. Furthermore, our MBP4+ADOSD4 scheme demonstrates robustness to variations in the maximum number of iterations, as discussed in Section VI-B.
For each data point in the plots, at least 100 logical error events are collected.
In Sections VI-D and VI-E, we first demonstrate the performance of the proposed BP+OSD schemes. Then, we analyze the execution time for these schemes in Section VI-G, illustrating the efficiency of ADOSD.
In the simulations of ADOSD4, $\Gamma$ is chosen according to (11) for the purpose of comparison with OSD-2. The soft reliability threshold is set to $\theta=0.999995$. Based on our simulations, selecting $\theta$ within the range of $0.999$ to $0.999995$ results in similar performance.
For comparison, we consider the following decoders: (1) the AMBP4 decoder [31] using a random serial schedule. (2) mOSD-($w$, $\lambda$), a partial order-$w$ OSD approach introduced in [42], which selects $\sum_{j=0}^{w}\binom{\lambda}{j}$ candidates for $\lambda$ less reliable bits based on the reliability metric Max. (3) the BP-OSD-CS($\lambda$) decoder in [43]. (4) BP-GDG algorithm in [54]. These decoders are summarized in Table I.
We observe that both parallel and serial schedules in BP yield similar results for toric codes and surface codes. However, the parallel schedule outperforms the serial schedule on the [[882, 48, 16]] GHP code. We believe that this is because that the parallel schedule typically results in unsettled oscillations in BP, which can be resolved by OSD, whereas the serial schedule may make premature decisions on some bits, leading to incorrect convergence in the decoding process. Therefore, we consider the parallel schedule for the following simulations. Similar observations are also made in [54, 65].
VI-A Optimization of the parameter $\alpha$ in MBP4 for ADOSD.
In this subsection, we demonstrate how to optimize the parameters of our BP+OSD schemes using LCS codes [59]. LCS codes are a specialized type of quantum code that combines the features of surface codes and lift-product constructions. Decoding LCS codes presents the most challenging problem among all the quantum codes considered in this paper. These codes cannot be effectively handled by our previous AMBP4 decoder [31], even though AMBP4 outperforms binary BP-OSD decoders [42, 43] on all other codes.
In the following, we consider the MBP4 decoder introduced in [31] as a predecoder for OSD. MBP4 includes a parameter $\alpha$ that controls the step size of message updates during BP iterations: when $\alpha<1$, the step size is enlarged; when $\alpha>1$, the step size is reduced. For $\alpha=1$, MBP4 is equivalent to the refined BP4 in [12]. We will specify the value of $\alpha$ only when $\alpha\neq 1$ is used.
Thus, optimizing the choice of $\alpha$ is crucial. For most nondegenerate codes, $\alpha=1$ is sufficient to achieve good decoding performance. However, for degenerate codes, selecting $\alpha>1$ can result in more stable message passing, leading to improved OSD performance. This optimization should be performed for each specific code, although the same value of $\alpha$ typically applies across an entire code family.
Figure 2 demonstrates the performance of MBP4+ADOSD4444 on the $[[175,7,7]]$ LCS code with various values of $\alpha$. The performance of MBP4+ADOSD4444 improves significantly and stabilizes as $\alpha$ increases from $1$ to approximately $1.6$.
For reference, we provide a curve $y=ax^{4}$ for some $a$ if errors of weight up to half of the code distance $\lfloor\frac{d-1}{2}\rfloor$ can be corrected. Therefore, MBP4+ADOSD4444 with $\alpha=1.6$ matches this prediction around LER $10^{-7}$. On the other hand, AMBP meets an earlier error floor around LER $10^{-3}$.
Figure 2 also includes a curve labeled MLE (independent X/Z), which is an estimate of the maximum likelihood binary decoding performance using integer programming, as reported in [59]. Since the decoding of X and Z errors is performed independently, this MLE curve represents the optimal performance only under the assumption that X and Z errors are uncorrelated. In [59], the performance of the BP-OSD-CS decoding algorithm from [43] was shown to approximately match the MLE curve. However, it is evident that MBP4+ADOSD4 significantly outperforms this MLE curve.
We observe a similar performance of MBP4+ADOSD4 for $[[15,3,3]]$, $[[65,5,5]]$ and $[[369,9.9]]$ LCS codes as well, while AMBP has an even higher error floor around LER $10^{-2}$ for the $[[369.9,9]]$ LCS code and a lower error floor around $10^{-5}$ for the $[[65,5,5]]$ LCS code.
Figure 2: MBP4+ADOSD4 decoding with various values of $\alpha$ for the $[[175,7,7]]$ LCS code. The curve MLE is taken from [59]. The notation $\alpha\in(1.6,-0.01,0.5)$ means that $\alpha$ is tested in the sequence $1.60,1.59,1.58,\dots,0.51,0.50$.
VI-B Robustness of the maximum iterations in MBP4+ADOSD4
Figure 3: MBP4+ADOSD4 decoding with various maximum number of iterations $T$ for the $[[175,7,7]]$ LCS code.
Next, we demonstrate that the performance of ADOSD4 remains robust even with variations in the maximum number of iterations in MBP4. We note that in ADOSD4, OSD-$w$ is invoked when HRSR fails. As a result, the overall algorithm maintains high accuracy.
However, the failures of the HRSR algorithm can impact the simulation time, as OSD-$w$ is significantly more time-consuming.
Figure 3 illustrates the performance of MBP4+ADOSD4 decoding with varying maximum numbers of iterations, $T=1,2,3,4,5,10,20,50,100,200$. The curve for $T=4$ is already good enough and smooth. While increasing the number of iterations results in slightly improved performance, the curves converge closely around an LER of $2\times 10^{-7}$ to $3\times 10^{-7}$.
In the extreme case of $T=1$, there are few highly reliable bits or valid highly reliable bits, causing the decoder to essentially revert to performing ordinary OSD4 with increased complexity.
We have demonstrated that the performance of ADOSD4 remains robust even with variations in the maximum number of iterations in MBP4. To balance accuracy and efficiency, we typically select $T$ within the range of $50$ to $100$.
VI-C Bivariate bicycle codes
A class of QLDPC codes, referred to as BB codes, was proposed in [55], demonstrating a high error threshold and a significantly higher code rate compared to topological codes. Several attempts at decoding this code family have been presented in [53, 54] under the code capacity noise model. However, the BP+CB decoder in [53] does not perform as well as BP-OSD-0. The BP+GDG decoder in [54] shows improved performance, surpassing BP-OSD-0 and being comparable to BP2-OSD-CS10 over independent X errors, as shown in [54, Figure 4].
Figure 4 presents the simulation results of MBP4+ADOSD4 and AMBP4 decoding for BB codes, both of which demonstrate significantly better performance than GDG decoding. For instance, consider the [[144,12,12]] BB code. At an $X$ error rate of $0.02$ (corresponding to a depolarizing rate of about $0.03$), BP+GDG achieves an LER of approximately $10^{-4}$. In contrast, both MBP4+ADOSD4 and AMBP4 achieve an LER below $4\times 10^{-6}$ at the same depolarizing rate. Moreover, MBP4+ADOSD4 outperforms AMBP4.
For the [[144,12,12]] code, we also include the decoding performance of MBP4+OSD4-2, which is nearly identical to that of MBP4+ADOSD4.
Figure 4: AMBP4 and MBP4+ADOSD4 decoding performance for various BB codes.
VI-D Generalized hypergraph-product codes
The $[[882,48,16]]$ GHP code [42] is highly-degenerate, which causes conventional BP to perform poorly on this code. With minimum distance of 16, we can gauge the effectiveness of a good decoding algorithm on this GHP code.
The performance curves of different combinations of MBP and OSD are shown in Figure 5, where $\alpha$ is either fixed at 1.6 or chosen by $\alpha(\epsilon)$. It is suggested in [31, Fig. 3] (arXiv version) that $\alpha$ and $\epsilon$ follow the relationship $\alpha(\epsilon)=-0.16\log_{10}(\epsilon)-0.48$, with the coefficients $-0.16$ and $-0.48$ determined through pre-simulations on the [[882, 48]] code. For comparison, we also plot the performance of AMBP4 [31] and the BP+mOSD-($w$, $\lambda$[42].
Figure 5: Comparison of various BP-OSD schemes on the $[[882,48]]$ GHP code. The curve BP+mOSD-(15,15) is from [42].
The results are summarized as follows.
  1. 1.
    MBP4 can be greatly enhanced by OSD.
  2. 2.
    MBP4+OSD4-00 outperforms MBP4+mOSD4-00 (without hard-decision reliability) by roughly half an order of performance. This suggests that the hard-decision history of BP is crucial in determining the reliability order.
  3. 3.
    MBP4+ADOSD4 performs slightly better than MBP4+OSD4-2, while also reducing the time consumption, as shown in Table IV in Section VI-G.
  4. 4.
    Both MBP4+ADOSD4 and MBP+OSD4-2 improve upon BP+mOSD-($15$, $15$[42] by more than an order of magnitude at $\epsilon=0.05$. Note that mOSD-($15$, $15$) essentially enumerates all the 15 less reliable bits.
Figure 6: Comparison of various metrics for reliability order in MBP4-OSD4 on the [[882, 48, 16]] GHP code.
In this paper, we propose using integrated hard-decision and soft reliabilities as a reliability metric (Definition 4). Herein we also simulate the performances of MBP4+mOSD based on the reliability metrics of Entropy or Max as shown in Figure 6, including mOSD4  as well. Two conclusions are as follows:
  1. 1.
    The reliability metric in Definition 4, which incorporates historical message information, outperforms other metrics that rely solely on the last belief distributions.
  2. 2.
    The performances of mOSD based on Definition 5, Entropy, or Max are mostly comparable; mOSD-$w$ based on Definition 5 is slightly better than the other two metrics at lower error rate for $w=2$.
We have conducted additional simulations on a [[625,25,8]] GHP code [43], as shown in Figure 7. In these examples, MBP4+ADOSD4 slightly outperforms both AMBP4 and MBP4+OSD4-2, with all three decoders performing significantly better than BP-OSD-CS60 in [43].
Figure 7: Decoding performance of the [[625,25,8]]] GHP code. Note that the performance curve for BP-OSD-CS60 is taken from [43, Figure 5], with its independent X/Z error rate $p$ rescaled by a factor of $1.5$ to approximate depolarizing error performance.
VI-E 2D topological codes
Next, we consider the decoding performance of our BP+OSD schemes on various 2D topological codes. Since AMBP4 is also capable of decoding various families of topological codes, it serves as a useful reference for evaluating the performance of the BP+OSD approach. It is important to examine the decoding performance when the logical error rate is around $10^{-6}$, as this can reveal whether an iterative decoder exhibits an error floor behavior.
Figure 8 illustrates the performance curves of MBP4+ADOSD4 and AMBP4 on the (4.8.8) color codes with distances ranging from 3 to 17. The results show that MBP4+ADOSD4 using $\alpha=1.6$ improves AMBP4 starting from $d=5$.
The advantages of our BP+OSD schemes over AMBP4 are also evident in rotated toric, rotated surface, XZZX, and (6.6.6) color codes. Additionally, MBP4+OSD4-2 and MBP4+ADOSD4exhibit nearly identical performance on these codes.
Figure 8: AMBP4 vs. MBP4+OSD4-2 on the (4.8.8) color codes.
To sum up, ADOSD4 and OSD4-2 exhibit nearly identical performances on toric, surface, and twisted XZZX codes, with OSD4-2 performing slightly better on the (6.6.6) and (4.8.8) color codes. The threshold analysis for these codes in the next subsection will reflect these results as well. We remark, however, the time consumption for ADOSD4 is significantly lower than that for OSD4-2, as will be shown in Section VI-G.
VI-F Threshold analysis for the topological codes
The decoding performance of a family of 2D topological codes is typically evaluated by the its simulated error threshold, which can be estimated using the scaling ansatz method [66, 67], which is explained as follows.
Suppose that the threshold value of a code family is $\epsilon_{\mathrm{th}}$. For a code in this family of distance $d$, if the logical error rate is $p_{L}$ at physical error rate $\epsilon$, then
$p_{L}=f(d^{\frac{1}{\nu}}(\epsilon-\epsilon_{\mathrm{th}})),$ (14)
where $f$ is a low-degree polynomial and $\nu$ is the critical exponent. The scaling ansatz method involves fitting this equation with the best values of $\epsilon_{\mathrm{th}}$, $\nu$, and $f$.
Figure 9: The threshold of BP4+OSD4-2 on the surface codes is about $17.67\%$ from the simulations. The dashed line stands for no error correction.
Figure 10: The accuracy of the threshold estimation about MBP4+OSD4-2 on the rotated surface codes can be demonstrated using the critical scaling approach. The estimation process begins with $d=7$ and continues up to $d=17$, ensuring that no biased information influences the results.
The threshold analysis for rotated surface codes is presented in Figure 9 and Figure 10. For each data point, we collected at least 10,000 logical error events. The analysis reveals a threshold of approximately 17.67% for BP4+OSD4-2 on surface codes. Similar analyses are conducted for toric, twisted XZZX, (6.6.6), and (4.8.8) color codes, using both BP4+OSD4-2 and BP4+ADOSD4. In the simulations, the data is fitted using a low-degree polynomial.
Note that MBP4 with $\alpha=1$ is used in the threshold analyses. Optimizing the value of $\alpha(\epsilon)$ for MBP can enhance overall performance.
The summarized results, including thresholds obtained with AMBP4, are shown in Table II. These results are consistent with the simulations in the previous subsection. Notably, BP4+OSD4-2 shows a significant enhancement over AMBP4 for surface and color codes.
TABLE II: Thresholds for 2D topological codes subjecte to depolarizing errors. The parameter for MBP4 is set to $\alpha=1.0$.
[…]
We also compare the thresholds of our decoder with those of other known decoders in Table III, including minimum-weight perfect matching (MWPM) [68, 69]. In [43], BP2+mOSD-(2,60) achieved a threshold of $p=9.9\%$ on toric codes for bit-flip errors. We estimate that it has a threshold of $\epsilon_{\mathrm{th}}=9.9\%\times\frac{3}{2}=14.85\%$ for depolarizing errors [1].
TABLE III: Thresholds of several decoders over depolarizing errors.
[…]
: estimated from the threshold of 9.9% in the bit-flip channel [43].
VI-G Time consumption
TABLE IV: Comparison of the time consumption between ADOSD4 and OSD4-2 at logical error rate about $2\times 10^{-6}$.
Code BP failure ADOSD4 OSD4-2 time LER
rate time ($\mu$s) time ($\mu$s) ratio ratio
toric 27.76% 88.63 4712.9 1.88% $\approx 0$dB
surface 67.69% 71.97 3549 2.82% $\approx 0$dB
(6.6.6) 10.32% 35.42 3812 0.93% -2.3dB
(4.8.8) 7.17% 26.50 1265 2.09% -3.9dB
XZZX 11.93% 22.29 1001.3 2.23% $\approx$ 0dB
GHP 0.86% 1188.7 45910.3 2.59% 2.49dB
: the column ‘time ratio’ represents the ratio of MBP4+ADOSD4 time to OSD4-2 time.
: the column ‘LER ratio’ represents the ratio of the logical error rate using ADOSD4 to that using OSD4-2. If the LER ratio is greater than 0dB, it means that ADOSD4 is better.
In the previous subsections, we have demonstrated the effectiveness of our BP+OSD schemes. In this subsection we analyze the time efficiency of our OSD schemes.
VI-G1 Time Consumption of ADOSD4 vs. OSD4-2
First, we compare the time consumption of OSD4-2 and ADOSD4 for various topological codes of comparable distances with the [[882, 48,16]] GHP code, including the [[324,2,18]] rotated toric code, the [[289,1,17]] rotated surface code, the [[[217, 1,17]] (6.6.6) color code, the [[161, 1,17]] (4.8.8) color code, and the [[145, 1,17]] twisted XZZX code. For each code, the simulated physical error rate is chosen so that the logical error rate of MBP4+ADOSD4 is approximately $2\times 10^{-6}$. We fix $\alpha=1$ in this subsection and the decoding performance can be further improved.
In the simulations, when BP fails, the BP outputs will be handled by ADOSD4 and OSD4-2, respectively, We note that ADOSD4 is invoked only when MBP4 fails. and their time consumption statistics will be counted separately in the following tables. For example, ADOSD4 time only counts the time to run Algorithm 4 when MBP4 fails.
The simulation results are summarized in Table IV and the unit of time is microseconds ($\mu$s) measured using a single-thread desktop computer. It can be observed that the time consumption of ADOSD4 is roughly $1\%$ to $3\%$ of the time consumption for OSD4-2 at logical error rate about $2\times 10^{-6}$ for the various quantum codes.
For the toric, surface, and twisted XZZX codes, ADOSD4 and OSD4-2 exhibit almost identical performances, while OSD4-2 performs better on the color codes and ADOSD4 performs better on the GHP code. This is reflected in the last column of Table IV.
To summarize, ADOSD4 is at least 30 times faster than OSD4-2, with a maximum performance loss of 4dB for topological codes at logical error rate $2\times 10^{-6}$. Additionally, for the [[882, 48]] GHP code, ADOSD4 also enhances performance.
VI-G2 Time Consumption of ADOSD4 vs. BP
We have demonstrated that ADOSD4 is generally faster than OSD4-2. Next, we compare the time consumption of ADOSD4 to MBP4, using surface codes.
TABLE V: Comparison of the time consumption for surface codes at logical error rate about $10^{-6}$.
code average time per MBP4 time per time per
distance iterations MBP4 failure ADOSD4 OSD4-2
iteration ($\mu$s) rate ($\mu$s) ($\mu$s)
11 22.3473 0.01464 21.46% 0.0272 0.667
13 37.1148 0.0212 35.97% 0.0469 1.342
15 52.6400 0.0287 51.47% 0.0506 2.155
: average over all samples, including MBP failures.
TABLE VI: Time Consumption of ADOSD4 vs. BP for the surface codes
code $\epsilon$ average iterations ADOSD4 time
distance when MBP4 succeeds in MBP4 iterations
11 1.7% 1.002 2
13 2.5% 1.79 2.21
15 3.3% 2.41 1.76
: average over only the samples when MBP4 succeeds.
Table V lists the time consumptions of MBP4, ADOSD4, and OSD4-2 for surface codes with distances of 11, 13, and 15. To achieve a logical error rate $10^{-6}$ using ADOSD4, the physical error rates are set to 0.017, 0.025, and 0.033 for codes with distances 11, 13, and 15, respectively. At least $10^{5}$ samples are collected for each case. For example, at $d=11$, the MBP4 failure rate is $21.46\%$ at physical error rate 1.7%. Since the maximum number of iterations is set to be $T=100$, this contributes to an average of $21.46$ iterations. Since the average iterations include MBP4 failures is 22.3473, the average iterations when MBP4
Table VI summarizes the time consumption of ADOSD4 in terms of MBP4 iterations. Specifically, the average time consumption per codeword for ADOSD4 is approximately equivalent to 1.76 to 2.21 MBP4 iterations. This shows that the time consumption for ADOSD4 is roughly equivalent to 1 to 2 successful MBP4 executions. Furthermore, the time usage for ADOSD4 is significantly smaller than OSD4-2.
VI-G3 HRSR and OSD-0 in ADOSD4
We have shown that ADOSD4 is efficient, thanks to Algorithm 3 and Corollary 10. These methods reduce the effective length using reliable subset reduction and allow us to perform only OSD4-0 by identifying degenerate error cosets. Now, we examine the effectiveness of these techniques on the surface code of distance 11, with similar results observed for distances 13 and 15.
Table VII enumerates the percentage of ADOSD4 trials that only perform OSD4-0 using Corollary 10. At $\epsilon=4.1\%$, 93.61% of ADOSD4 trials already perform only OSD4-0, increasing to 99.84% at $\epsilon=1.7\%$. Thus, at lower error rates, almost all trials execute OSD4-0.
Moreover, the HRSR algorithm effectively reduces the dimensions of the check matrix in both rows and columns. Table VII also provides the percentages of trials where the number of rows and columns in the check matrix is reduced to 20% or 30% of the original (denoted as $20\%$ and $30\%$ effective dimensions in the table). These percentages increase quickly as the error rate decreases. With fewer than 30% effective dimensions, the time consumption for Gaussian elimination reduces to roughly $0.3^{3}=2.7\%$, consistent with the results in Table IV.
TABLE VII: Percentage of OSD-0 in ADOSD4 on the surface code of distance $11$
depolarizing rate $\epsilon$ 1.7% 2.7% 4.1%
LER $10^{-6}$ $10^{-5}$ $10^{-4}$
OSD-0 99.84% 98.85% 93.61%
20% effective 44.13% 26.30% 5.13%
dimensions
30% effective 94.04% 84.46% 57.91%
dimensions
VII Conclusion
We proposed quaternary OSD with integrated hard- and soft-decision reliabilities to retain the $X/Z$ correlations. Our BP-OSD schemes have demonstrated highly effective decoding capabilities for various sparse quantum codes in the code capacity noise model, including both CSS and non-CSS codes. Additionally, when the logical error rate is around $10^{-6}$, ADOSD4 can further enhance the decoding process by offering accelerated decoding while maintaining comparable performance. As a result, ADOSD4 is applicable to small or medium-sized codes with $n\approx 1000$, significantly enhancing the performance of BP in the code capacity noise model. In our simulations, we restricted the number of error candidates to the size of order-2 OSD; better performance of ADOSD4 can be achieved by allowing for greater options.
We note that while all the techniques developed in this paper are presented in the context of quaternary error variables, they can naturally be adapted to binary OSD for handling CSS codes. This includes hard-decision reliability, highly reliable subset reduction, approximate degenerate decoding, and degenerate candidate pruning conditions, by restricting the approach to either $X$- or $Z$-stabilizers. Furthermore, these techniques can be integrated with other methods available in the literature.
Although the intention of [56] and [57] is quite similar to our HRSR algorithm, the underlying approaches differ: their methods focus on dividing the decoding problem into several smaller problems, whereas our approach involves simply removing the most reliable bits. Consequently, it may be possible to combine their methods with ours. For example, apply HRSR first, followed by either of the algorithms.
In our investigations, we primarily focused on MBP4+OSD4-2 and ADOSD4. However, MBP4+OSD4-0 has also exhibited excellent performance for most codes, aligning with the findings in [42]. In particular, we found that ADOSD4 mostly performs only MBP4+OSD4-0 at low error rates. On the other hand, at low error rates (e.g., below $10^{-3}$), there are small non-isolated errors, so the effective length for OSD is relatively small, allowing us to perform higher-order OSD. We have observed some instances of surface codes with order $w$ up to 13 in (11).
An interesting observation is that we can identify bit-flips in error candidates corresponding to multiplying a stabilizer operator, as stated in Lemma 8 and Corollary 10. This information can be used to improve the algorithm by flipping effective bits in approximate degenerate decoding. Additionally, another potential improvement is extending the candidate pruning condition by Fossorier et al[46] to quantum codes.
We chose the binary representation of quantum codes and syndrome correction for developing the OSD algorithm. A quantum code can be interpreted as a quaternary additive code with binary error syndromes [5]. Thus, one might consider formulating a system consisting of $n$ quaternary variables and $n-k$ additive constraints and employ a Gaussian elimination process for quaternary additive groups. However, for practical implementation, we chose to binarize this process.
There are various reliability metrics that consider the (weighted) accumulation of the log-likelihood ratio (LLR) signs for each bit during BP iterations [70, 71, 72, 73]. Our hard-decision reliability can be viewed as an extension of this idea, focusing on the length of the last run in the hard-decision vectors. The reliability order in Definition 4 can be adapted to employ alternative metrics, such as Entropy or Max. From our experiments, the integrated hard- and soft-decision reliabilities yielded the best results.
Finally, our BP-OSD scheme can be extended to handle both data and syndrome errors in the phenomenological noise model [74, 75], using techniques proposed in [76, 77, 78], or to the circuit-level noise model [79], as demonstrated in our recent work [78].
Currently, there is significant progress in circuit-level error decoding [80, 81, 55, 82]. Various decoders have been shown to work for circuit-level noise models, including BP-OSD-CS [43], guided decimal guessing (GDG)[54], ambiguity clustering[56], localized statistics decoding (LSD)[57], and ordered Tanner forest (OTF) [83]. These decoders are based on detector error models generated by STIM [84] and utilize binary BP combined with postprocessing.
However, there are several code capacity error decoding problems where BP-OSD is suboptimal, indicating potential for further improvement in the circuit-level noise model. Therefore, we aim to extend our BP+OSD method to address these more demanding scenarios effectively.

### v2

@@
 Reliable quantum communication is a critical area of research, essential for scaling quantum systems, enabling the exchange of quantum information across distances, and facilitating multiparty protocols.
-Quantum states are inherently fragile, requiring the implementation of quantum error correction to mitigate the effects of noise and decoherence [2, 3].
-In this work, we explore advanced decoding techniques for quantum error-correcting codes to enhance the reliability and efficiency of quantum communication.
+Quantum states are inherently fragile, requiring the implementation of quantum error correction to mitigate the effects of noise and decoherence [2, 3, 4].
 Quantum stabilizer codes, analogous to classical linear block codes, allow efficient encoding and binary syndrome decoding [5, 6, 7].
@@
 A recently developed modified BP4 algorithm, known as MBP4, incorporates additional memory effects and has demonstrated the ability to handle topological codes [32].
-The adaptive version of MBP4, termed AMBP4, further improves BP performance while maintaining nearly linear time complexity [31].
-Nonetheless, its practical application involves an unknown parameter search and may not be fully parallelized.
-We investigate ordered statistic decoding (OSD) in quantum coding theory as a complementary approach to BP.
-Fossorier and Lin [32] first proposed OSD in classical coding theory as an approach to approximate maximum-likelihood decoding [33, 34, 35, 36, 37, 38, 39, 40, 41].
-The OSD algorithm searches for possible solutions by sorting the reliability measures of symbols while maintaining controllable complexity [32, 42, 43, 44, 45].
-In cases where BP fails to find a valid error for a given syndrome, syndrome-based OSD can be utilized [46].
-Moreover, OSD has been demonstrated to be particularly useful for degenerate quantum codes by Panteleev and Kalachev [42].
-The core function of OSD is Gaussian elimination on the parity-check matrix of dimension $m\times n$, which identifies the independent unreliable bits and has a complexity of ${\cal O}(nm^{2})$.
-The remaining reliable bits can be used to regenerate the most probable error.
-In an order-$w$ OSD, the set of reliable bits can be flipped up to $w$ bits, generating a list of probable solutions.
-However, this technique incurs at least a cubic time complexity overhead from Gaussian elimination, along with additional costs for high-order examination, despite its effectiveness.
-Fossorier et al[46] identified specific candidate pruning conditions to reduce ineffective high-order OSD processing, as such operations do not yield improved error candidates.
-Many techniques have been proposed in classical coding theory to reduce the complexity of high-order OSD, by focusing on more probable candidates, thereby reducing the number of OSD candidates required [47, 48, 49, 50, 51, 52].
+We investigate ordered statistic decoding (OSD) as a complementary approach to BP in quantum coding theory.
+Originally proposed by Fossorier and Lin [33] in classical coding theory for approximating maximum-likelihood decoding [34, 35, 36, 37, 38, 39, 40, 41, 42], OSD sorts reliability measures of symbols to find probable solutions while maintaining controllable complexity [33, 43, 44, 45, 46].
+Syndrome-based OSD has been used when BP fails to identify a valid error [47] and is particularly effective for degenerate quantum codes [43].
+The core operation of OSD is Gaussian elimination on an $m\times n$ parity-check matrix, identifying unreliable bits with a complexity of $O(nm^{2})$.
+Order-$w$ OSD flips up to $w$ reliable bits to generate error candidates, but this incurs high computational costs.
+To mitigate complexity, Fossorier et al.[47] proposed candidate pruning, and various classical techniques further optimized high-order OSD by reducing the number of candidates examined[48, 49, 50, 51, 52, 53].
 […]
@@
 These methods primarily rely on binary BP to handle quaternary errors, neglecting error correlations, which suggests that binary decoders are inherently suboptimal.
-In contrast, the quaternary AMBP4 decoder in [31] has been shown to outperform all the aforementioned decoders without requiring postprocessing.
-For supporting examples, please refer to Figure 2, Figure 4, and Figure 7 in the simulation section.
-Since OSD acts as a complementary strategy to BP for cases where BP fails, BP-OSD is naturally expected to surpass the performance of AMBP4.
-In this paper, we propose a BP4-based OSD to achieve this improvement.
-Additionally, we introduce an efficient OSD algorithm tailored for quantum codes.
+In contrast, the quaternary AMBP4 decoder (the adaptive version of MBP4) from [32] has been shown to outperform all the aforementioned decoders across various quantum codes, including topological codes [25], BB codes [56], and generalized hypergraph-product (GHP) codes [43, 44], without requiring postprocessing.
+For supporting examples, please refer to Figure 7 in Section VI.
+OSD serves as a complementary strategy to BP, addressing cases where BP fails.
+As a result, BP-OSD is naturally expected to outperform AMBP4.
+In this paper, we propose a BP4-based OSD to achieve this improvement and introduce an efficient OSD algorithm for quantum codes.
 First, the sorting process plays a pivotal role in OSD [33].
@@
 Next, we consider approximate degenerate OSD in quantum codes.
-A typical OSD generates a list of error candidates for a given error syndrome and selects the one with the minimum weight, which is optimal for classical codes.
+A typical OSD generates a list of error candidates for a given error syndrome and selects the one with minimum weight, which is optimal for classical codes.
 For quantum codes, however, the optimal decoding criterion involves choosing the error coset with the highest probability, a criterion that is challenging to implement in practice.
-To address this, we consider approximate degenerate decoding in OSD by focusing only on relatively low-weight terms within an error coset.
+To address this, we consider approximate degenerate decoding in OSD by focusing only on relatively low-weight terms within a coset.
 Additionally, we introduce a degenerate error candidate pruning condition for stabilizer codes.
-We show that certain bit-flips in OSD correspond to multiplying stabilizers, resulting to degenerate error candidates.
+We show that certain bit-flips in OSD correspond to multiplying stabilizers, resulting in degenerate error candidates.
 These bit-flips may alter the error weight but do not change the error coset, allowing them to be ignored in approximate degenerate OSD.
-For improving OSD efficiency, using a minimum weight decoding criterion is beneficial.
-A key observation is that if all bit-flips correspond to multiplying a stabilizer, there is no need to implement high-order OSD.
-We will demonstrate that this is indeed the case at low error rates by simulations in Section VI-G.
-We note that a degeneracy condition is also utilized in ambiguity clustering to facilitate the resolution of subproblems [56].
+To improve OSD efficiency, the minimum weight decoding criterion is applied.
+A key observation is that if all bit-flips correspond to multiplying a stabilizer, high-order OSD is unnecessary.
+We note that a degeneracy condition is also utilized in ambiguity clustering to facilitate subproblem resolution [57].
 Integrating all the mentioned techniques, we propose an approximate degenerate OSD algorithm, abbreviated as ADOSD4.
-We conduct simulations of MBP4 and our OSD schemes OSD4 and ADOSD4 on several quantum codes under depolarizing errors in the code capacity noise model, including BB codes [55], rotated toric codes and rotated surface codes [24], (6.6.6) and (4.8.8) color codes [27], twisted XZZX codes on a torus [58], generalized hypergraph-product (GHP) codes [42, 43], and lift-connected surface (LCS) codes [59].
+We conduct simulations of MBP4 and our OSD schemes OSD4 and ADOSD4 on several quantum codes under depolarizing errors in the code capacity noise model, including BB codes [56], rotated toric codes and rotated surface codes [25], (6.6.6) and (4.8.8) color codes [28], twisted XZZX codes on a torus [59], GHP codes [43, 44], and lift-connected surface (LCS) codes [60].
 The results demonstrate that our proposed schemes outperform previous approaches in the literature, both in terms of error thresholds for topological codes and low error rate performance (logical error rate around $10^{-6}$).
-Both our MBP4+OSD4 and MBP4+ADOSD4 decoders outperform other binary BP-OSD schemes in these quantum codes.
-We achieve thresholds of approximately 17.5%–17.7% for the toric, surface, and twisted XZZX codes, and thresholds of 15%–15.42% for the (4.8.8) and (6.6.6) color codes, as summarized in Table II.
-Comparisons of our results with existing threshold values for the toric codes are provided in Table III.
-We also compare the performance of various reliability metrics on the GHP code, justifying our choice of reliability metric.
-Finally, we analyze the time efficiency of our OSD schemes for various topological codes of comparable distances to the $[[882,48,16]]$ GHP code using single-threaded CPU implementations.
-The simulation results, summarized in Table IV, show that the time consumption of ADOSD4 is approximately 1% to 3% of that for order-2 OSD4 with similar decoding performance.
-Moreover, Table VI presents the time consumption of ADOSD4 in terms of BP iterations for surface codes with distances from 11 to 15.
-Specifically, the average time consumption per codeword for ADOSD4 is approximately equivalent to 2 to 3 BP4 iterations, whereas a successful BP trial typically takes 2 to 3 iterations on average.
+Our MBP4+OSD4 and MBP4+ADOSD4 decoders outperform other binary BP-OSD schemes in these quantum codes.
+We achieve thresholds of approximately 17.5%–17.7% for the toric, surface, and twisted XZZX codes, and thresholds of 15%–15.42% for the (4.8.8) and (6.6.6) color codes, as summarized in Table III.
+They even surpass the minimum weight perfect matching (MWPM) decoder on surface and toric codes [69, 70].
+Finally, we analyze the time efficiency of our OSD schemes for these quantum codes using single-threaded CPU implementations.
+The simulation results, summarized in Table IV, show that the time consumption of ADOSD4 for surface codes with distances from 11 to 15 is approximately 2% to 4% of that for order-2 OSD4 with similar decoding performance at logical error rate around $10^{-6}$.
+Moreover, the average time consumption per codeword for ADOSD4 is approximately equivalent to 2 to 3 BP4 iterations, whereas a successful BP trial typically takes 2 to 3 iterations on average.
 This suggests that ADOSD4 is an excellent complement to BP, especially since ADOSD4 is used only when BP fails.
 This paper is organized as follows.
-We introduce quantum stabilizer codes in Section II.
+We introduce stabilizer codes in Section II.
 Section III discusses reliability metrics and the associated OSD algorithm.
@@
 In this section, we review the basic of stabilizer codes and define the relevant notation [5, 6, 7].
-Let $I=\begin{bmatrix}1&0\\ 0&1\end{bmatrix},$ $X=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},$ $Z=\begin{bmatrix}1&0\\ 0&-11\end{bmatrix},$ and $Y=iXZ$ denote the Pauli matrices.
-Consider the $n$-fold Pauli group
-$\mathcal{G}_{n}\triangleq\left\{cM_{1}\otimes\dots\otimes M_{n}:c\in\{\pm 1,\pm i\},M_{j}\in\{I,X,Y,Z\}\right\}.$
+Let $I=\begin{bmatrix}1&0\\ 0&1\end{bmatrix},$ $X=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},$ $Z=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix},$ and $Y=iXZ$ denote the Pauli matrices.
+The $n$-fold Pauli group, $\mathcal{G}_{n}$, is the group of $n$-qubit Pauli operators, defined as
+$\mathcal{G}_{n}\triangleq\left\{\omega M_{1}\otimes\dots\otimes M_{n}:\omega\in\{\pm 1,\pm i\},M_{j}\in\{I,X,Y,Z\}\right\}.$
 Every nonidentity Pauli operator in $\mathcal{G}_{n}$ has eigenvalues ±$\pm 1$.
 Any two Pauli operators either commute or anticommute with each other.
-The weight of a Pauli operator $E\in{\cal G}_{n}$, denoted $\mathrm{wt}\left(E\right)$, refers to the number of its nonidentity components.
-We consider Pauli errors in this paper, assuming independent depolarizing errors with rate $\epsilon$.
+The weight of a Pauli operator $g\in{\cal G}_{n}$, denoted $\mathrm{wt}\left(g\right)$, refers to the number of its nonidentity components.
+We consider Pauli errors, assuming independent depolarizing errors with rate $\epsilon$.
 Each qubit independently experiences an $X$, $Y$, or $Z$ error with probability $\epsilon/3$ and no error with probability $1-\epsilon$.
-Therefore, at a small error rate $\epsilon$, low-weight Pauli errors are more likely to occur.
+Therefore, at a small error rate $\epsilon$, low-weight Pauli errors are more likely to occur across $n$ qubits.
 A stabilizer group $\mathcal{S}$ is an Abelian subgroup in $\mathcal{G}_{n}$ such that $-I^{\otimes n}\not\in\mathcal{S}$ [7].
@@
 Then $\mathcal{S}$ defines an $[[n,k,d]]$ stabilizer code $\mathcal{C}(\mathcal{S})$ that encodes $k$ logical qubits into $n$ physical qubits:
-$\mathcal{C}(\mathcal{S})=\left\{\ket{\psi}\in\mathbb{C}^{2^{n}}:S\ket{\psi}=\ket{\psi},~{}\forall\,S\in\mathcal{S}\right\}.$
+$\mathcal{C}(\mathcal{S})=\left\{\ket{\psi}\in\mathbb{C}^{2^{n}}:g\ket{\psi}=\ket{\psi},~{}\forall\,g\in\mathcal{S}\right\}.$
 The elements in $\mathcal{S}$ are called stabilizers. The parameter $d$ represents the minimum distance of the code such that any Pauli error of weight less than $d$ is detectable.
-An error $E\in{\cal G}_{n}$ can be detected by ${\cal C}({\cal S})$ through stabilizer measurements if $E$ anticommutes with some of its stabilizers.
-Suppose that $\{S_{i}\}_{i=1}^{m}$, where $m\geq n-k$, is a set of stabilizers that generates ${\cal S}$.
+An error $e\in{\cal G}_{n}$ can be detected by ${\cal C}({\cal S})$ through stabilizer measurements if $e$ anticommutes with some of the stabilizers.
+Suppose that $\{g_{i}\}_{i=1}^{m}$, where $m\geq n-k$, is a set of stabilizers that generates ${\cal S}$.
 Their binary measurement outcomes are referred to as the error syndrome of $e$.
@@
 Therefore, the minimum distance of ${\cal C}({\cal S})$ is defined as the minimum weight of a nontrivial logical operator.
-It can be observed that $E$ and $ES$ for $S\in{\cal S}$ have the same effects on the codespace.
+It can be observed that $e\in{\cal G}_{n}$ and $eg$ for $g\in{\cal S}$ have the same effects on the codespace.
 They are called degenerate to each other.
@@
 A $\delta$-approximate degenerate decoding ($\delta$-ADD) criterion aims to find an error estimate $\hat{e}$ that matches the given syndrome, maximizing the dominant terms of its coset probability
-$\sum_{S\in{\cal S},\mathrm{wt}\left(S\right)\leq\delta}\Pr\big{\{}\hat{E}S\big{\}}.$ (1)
+$\sum_{g\in{\cal S},\mathrm{wt}\left(g\right)\leq\delta}\Pr\big{\{}\hat{e}g\big{\}}.$ (1)
 Note that 00-ADD reduces to minimum weight decoding, while $n$-ADD corresponds precisely to degenerate maximum likelihood decoding.
 II-A Binary representations
-Without loss of generality, we assume that each stabilizer in $\{S_{i}\}_{i=1}^{m}$ is of the form $S_{i}=S_{i1}\otimes S_{i2}\otimes\dots\otimes S_{in}$, where $S_{ij}\in\{I,X,Y,Z\}$.
-Then we can study the decoding problem in the binary vector space [4, 6], using a mapping $\varphi:$
+We can study the decoding problem in the binary vector space [5, 7], using a mapping $\varphi:{\cal G}_{1}\rightarrow\{0,1\}^{2}$
 $I\mapsto\left[\begin{array}[]{c|c}0&0\end{array}\right],X\mapsto\left[\begin{array}[]{c|c}1&0\end{array}\right],Z\mapsto\left[\begin{array}[]{c|c}0&1\end{array}\right],Y\mapsto\left[\begin{array}[]{c|c}1&1\end{array}\right].$
-A check matrix of the stabilizer code ${\cal C}({\cal S})$ is a binary matrix ${\mathbf{S}}=[{\mathbf{S}}_{ij}]\in\{0,1\}^{m\times 2n},$ where
-$\left[\begin{array}[]{c|c}{\mathbf{S}}_{ij}&{\mathbf{S}}_{i(j+n)}\end{array}\right]=\varphi(S_{ij}).$
-An $n$-fold Pauli operator $E\in{\cal G}_{n}$ can also be represented by $\varphi(E)\triangleq{\mathbf{E}}=\left[\begin{array}[]{c|c}{\mathbf{E}}^{X}&{\mathbf{E}}^{Z}\end{array}\right]\in\{0,1\}^{2n}$, where ${\mathbf{E}}^{X}$ and ${\mathbf{E}}^{Z}$ are the indicator vectors of $X$ and $Z$ components of $E$, respectively.
-Then the error syndrome of $E$ is
-${\mathbf{z}}={\mathbf{S}}\Lambda{\mathbf{E}}^{\top}\in\{0,1\}^{m\times 1},$ (2)
+We extend the $\varphi$ to ${\cal G}_{n}$ as follows: For $g=M_{1}\otimes M_{2}\otimes\dots\otimes M_{n}\in{\cal G}_{n}$,
+$\varphi(g)\triangleq{\mathbf{g}}=\left[\begin{array}[]{c|c}{\mathbf{g}}^{X}&{\mathbf{g}}^{Z}\end{array}\right]=\left[\begin{array}[]{c|c}{\mathbf{g}}^{X}_{1}\cdots{\mathbf{g}}^{X}_{n}&{\mathbf{g}}^{Z}_{1}\cdots{\mathbf{g}}^{Z}_{n}\end{array}\right],$
+where ${\mathbf{g}}^{X}\in\{0,1\}^{n}$ and ${\mathbf{g}}^{Z}\in\{0,1\}^{n}$ are the indicator vectors of $X$ and $Z$ components of $g$, respectively, such that $\left[\begin{array}[]{c|c}{\mathbf{g}}_{j}^{X}&{\mathbf{g}}_{j}^{Z}\end{array}\right]=\varphi(M_{j}).$
+A check matrix of ${\cal C}({\cal S})$ corresponding a set of stabilizer generators $\{g_{i}\}_{i=1}^{m}$ is an $m\times 2n$ binary matrix ${\mathbf{H}}$ with rows $\varphi(g_{1}),\dots,\varphi(g_{m})$.
+An error $e\in{\cal G}_{n}$ can be represented by ${\mathbf{e}}=\varphi(e)$ and its error syndrome is
+${\mathbf{s}}={\mathbf{H}}\Lambda{\mathbf{e}}^{\top}\in\{0,1\}^{m\times 1},$ (2)
 where ${\mathbf{e}}^{\top}$ denotes the transpose of ${\mathbf{e}}$, $\Lambda=\left[\begin{array}[]{c|c}\mathbf{0}&{\mathbf{I}}_{n}\\ {\mathbf{I}}_{n}&\mathbf{0}\end{array}\right],$ ${\mathbf{I}}_{n}$ is the $n\times n$ identity matrix, and $\mathbf{0}$ is the zero matrix of appropriate dimensions.
-Consequently, the decoding problem is to solve this system of linear equations with $2n$ binary variables ${\mathbf{E}}_{j}$ that are most probable.
-Note that ${\mathbf{S}}$ is of rank $n-k$ so the degree of freedom of this binary system is $n+k$.
+Suppose this code has $2k$ independent logical operators, $\bar{X}_{i}$ and $\bar{Z}_{j}$ for $i,j=1,\dots,k$, which, together with the stabilizers $\{g_{i}\}$, generate the normalizer group $N({\cal S})$.
+The $2k\times 2n$ binary logical matrix ${\mathbf{L}}$ is defined with rows $\varphi(\bar{X}_{1}),\dots,\varphi(\bar{X}_{k}),\varphi(\bar{Z}_{1}),\dots,\varphi(\bar{Z}_{k})$.
+Since stabilizers commute with logical operators, we have ${\mathbf{H}}\Lambda{\mathbf{L}}^{\top}=\mathbf{0}$.
+The decoding problem is to solve the system of linear equations (2) with $2n$ binary variables in ${\mathbf{e}}\in\{0,1\}^{2n}$ that are most probable.
+Note that ${\mathbf{H}}$ is of rank $n-k$ so the degree of freedom of this binary system is $n+k$.
+An error $e\in{\cal G}_{n}$ is classified as a logical error for a decoder if the decoder fails or outputs $\hat{e}\in{\cal G}_{n}$ such that
+${\mathbf{H}}\Lambda({\mathbf{e}}+\hat{{\mathbf{e}}})^{\top}=\mathbf{0}$ (3)
+but
+${\mathbf{L}}\Lambda({\mathbf{e}}+\hat{{\mathbf{e}}})^{\top}\neq\mathbf{0}.$ (4)
 II-B Belief Propagation (BP) Decoding
 We explain how a quaternary BP algorithm works in the following [13], which is computed in linear domain. A log-likelihood version of BP can be found in [14].
-Given a check matrix ${\mathbf{S}}$, a syndrome ${\mathbf{z}}$, and a depolarizing error rate $\epsilon$, BP performs the following steps:
+Given a check matrix ${\mathbf{H}}$, a syndrome ${\mathbf{s}}$, and a depolarizing error rate $\epsilon$, BP performs the following steps:
   1. 1.
@@
 If a valid error estimate is generated, it will be accepted as the solution; otherwise, go to the next iteration starting from step 2).
-Steps 2) and 3) will be iterated for a maximum number of $T$ iterations, where $T$ is chosen in advance. If no valid estimate is obtained after $T$ iterations, BP will claim a failure.
-Typically, the initial error distribution ${\mathbf{p}}_{i}$ is defined as $({1-\epsilon},\frac{\epsilon}{3},\frac{\epsilon}{3},\frac{\epsilon}{3})$.
-However, one may instead choose a fixed initialization by a given value $\epsilon_{0}\neq\epsilon$ for BP as long as BP performs well at this $\epsilon_{0}$.
-See more discussion on fixed initialization in [63, 13, 31].
+Steps 2) and 3) will be iterated for a maximum number of ${\mathrm{T}}$ iterations, where ${\mathrm{T}}$ is chosen in advance.
+If no valid estimate is obtained after ${\mathrm{T}}$ iterations, BP will claim a failure.
 III Order-statistic decoding based on quaternary belief propagation
-In the aforementioned syndrome decoding problem, we aim to solve a system of linear equations given by (2), for a given parity-check matrix ${\mathbf{S}}$ and an error syndrome ${\mathbf{z}}$ with $2n$ binary variables.
+In the aforementioned syndrome decoding problem, we aim to solve a system of linear equations given by (2), for a given parity-check matrix ${\mathbf{H}}$ and an error syndrome ${\mathbf{s}}$ with $2n$ binary variables in ${\mathbf{e}}$.
 The linear system has $n-k$ independent constraints, impliying that there are $n+k$ degrees of freedom in the vector ${\mathbf{e}}$.
 Given the values of these $n+k$ bits, one can reconstruct ${\mathbf{e}}$.
-If BP fails to provide a valid error estimate after $T$ iterations, OSD will be utilized.
+If BP fails to provide a valid error estimate after ${\mathrm{T}}$ iterations, OSD will be utilized.
 In the context of OSD, a critical step involves identifying and sorting the accurate and reliable coordinates to establish $n+k$ linearly independent bits.
@@
 To utilize the quaternary distribution generated by BP4 for each qubit error in OSD, we propose a method that employs the hard-decision history from all BP iterations and the output probabilities from the last BP iteration.
-Suppose that BP fails to provide a valid error after a maximum of $T$ iterations.
+Suppose that BP fails to provide a valid error after a maximum of ${\mathrm{T}}$ iterations.
 At this point, we have the output distribution ${\mathbf{q}}_{i}=(q_{i}^{I},q_{i}^{X},q_{i}^{Y},q_{i}^{Z})$, $i=1,2,\dots,n$ from the last BP iteration.
-Additionally, we have the hard-decision results for all $T$ iterations, denoted by ${\mathbf{h}}_{i}^{(j)}\in\{I,X,Y,Z\}$ for $i=1,\dots,n$ and $j=1,\dots,T$.
-For example, ${\mathbf{h}}_{i}^{(1)},{\mathbf{h}}_{i}^{(2)},\dots,{\mathbf{h}}_{i}^{(T)}$ represent the hard-decision result history of BP at qubit $i$ and ${\mathbf{h}}_{1}^{(j)},{\mathbf{h}}_{2}^{(j)},\dots,{\mathbf{h}}_{n}^{(j)}$ represent the hard-decision results of BP at the $j$-th iteration.
-Our objective is to assign reliability for each Pauli $X$ and $Z$ error. Following the reliability definitions in our previous paper [1], we define two types of measures in the following.
+Additionally, we have the hard-decision results for all ${\mathrm{T}}$ iterations, denoted by ${\mathbf{h}}_{i}^{(j)}\in\{I,X,Y,Z\}$ for $i=1,\dots,n$ and $j=1,\dots,{\mathrm{T}}$.
+For example, ${\mathbf{h}}_{i}^{(1)},{\mathbf{h}}_{i}^{(2)},\dots,{\mathbf{h}}_{i}^{({\mathrm{T}})}$ represent the hard-decision result history of BP at qubit $i$ and ${\mathbf{h}}_{1}^{(j)},{\mathbf{h}}_{2}^{(j)},\dots,{\mathbf{h}}_{n}^{(j)}$ represent the hard-decision results of BP at the $j$-th iteration.
+Our objective is to assign reliability for each Pauli $X$ and $Z$ error. Following the reliability definitions in [1, 43], we define two types of measures in the following.
 Definition 2.
 Define a hard-decision reliability vector $\bm{\eta}=(\bm{\eta}_{1},\dots,\bm{\eta}_{n})\in\mathbb{Z}^{n}$ such that $\bm{\eta}_{i}$ represents the number of consecutive iterations during which the error at qubit $i$ remains unchanged until the final hard decision is made.
-In other words, $\bm{\eta}_{i}$ is the length of the last run in the string ${\mathbf{h}}_{i}^{(1)},{\mathbf{h}}_{i}^{(2)},\dots,{\mathbf{h}}_{i}^{(T)}$.
+In other words, $\bm{\eta}_{i}$ is the length of the last run in the string ${\mathbf{h}}_{i}^{(1)},{\mathbf{h}}_{i}^{(2)},\dots,{\mathbf{h}}_{i}^{({\mathrm{T}})}$.
 For example, if the hard-decision outputs at the first qubit over the last five iterations are $X,Y,Z,Z,Z$, then the final hard decision output is $Z$ and $\bm{\eta}_{1}=3$.
 This means that the error at qubit 1 persisted unchanged as $Z$ for three iterations when BP stops.
-For implementation, storing the entire hard-decision history ${\mathbf{h}}$ is unnecessary.
-One can record only the hard-decision outputs at the latest iteration and uses the latest belief vector ${\mathbf{q}}_{i}$ to update the values of $\{\bm{\eta}_{i}:i=1,\dots,n\}$.
-Our modified BP4 algorithm, which incorporates hard-decision reliability measures, is presented in Algorithm 1.
+When BP fails, the belief distributions ${\mathbf{q}}_{i}=(q_{i}^{I},q_{i}^{X},q_{i}^{Y},q_{i}^{Z})$ will be utilized in the reliability measure.
+Definition 3.
+Define two soft reliability functions
+$\phi^{X}(i)$ $=\max\{q_{i}^{X}+q_{i}^{Y},q_{i}^{I}+q_{i}^{Z}\},$
+$\phi^{Z}(i)$ $=\max\{q_{i}^{Z}+q_{i}^{Y},q_{i}^{I}+q_{i}^{X}\},$
+for $i\in\{1,2,\dots,n\}$.
+Note that $\{q_{i}^{X}+q_{i}^{Y},q_{i}^{I}+q_{i}^{Z}\}$ is a binary distribution indicating the likelihood of an $X$ error occuring at the $i$-th qubit and similarly for $\{q_{i}^{X}+q_{i}^{Y},q_{i}^{I}+q_{i}^{Z}\}$.
+Our modified BP4 algorithm is presented in Algorithm 1.
 Algorithm 1 BP4
-Input: check matrix ${\mathbf{S}}\in\{0,1\}^{m\times 2n}$, syndrome ${\mathbf{z}}\in\{0,1\}^{m\times 1}$, depolarizing error rate $\epsilon$, maximum number of iterations $T$.
-Output: A valid error estimate $\hat{{\mathbf{E}}}\in\{0,1\}^{2n}$ and belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$.
+Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, error rate $\epsilon$, maximum number of iterations ${\mathrm{T}}$.
+Output: A valid error estimate $\hat{{\mathbf{e}}}\in\{0,1\}^{2n}$, belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$ and hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$.
 Initialization:
@@
 Steps:
-for $j=1$ to $T$ do
-     Update $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, using ${\mathbf{S}},{\mathbf{z}},$ and $\{{\mathbf{p}}_{i}\}_{i=1}^{n}$.
+for $j=1$ to ${\mathrm{T}}$ do
+     Update $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, using ${\mathbf{H}},{\mathbf{s}},$ and $\{{\mathbf{p}}_{i}\}_{i=1}^{n}$.
      for $i=1$ to $n$ do
@@
      end for
-     if  the hard-decision ${{\mathbf{h}}}$ matches the syndrome ${\mathbf{z}}$ then
-         return $\varphi({\mathbf{h}})$ and $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$.
+     if  the hard-decision ${{\mathbf{h}}}$ matches the syndrome ${\mathbf{s}}$ then
+         return “Success" and $\varphi({\mathbf{h}})$.
      end if
 end for
-return $\mathbf{0}$ and $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$. $\triangleright$ BP fails.
-When BP fails, we are left with belief distributions ${\mathbf{q}}_{i}=(q_{i}^{I},q_{i}^{X},q_{i}^{Y},q_{i}^{Z})$ for $i=1,2,\dots,n$, which will be utilized in the reliability measure.
-Definition 3.
-Define two soft reliability functions
-$\phi^{X}(i)$ $=\max\{q_{i}^{X}+q_{i}^{Y},q_{i}^{I}+q_{i}^{Z}\},$
-$\phi^{Z}(i)$ $=\max\{q_{i}^{Z}+q_{i}^{Y},q_{i}^{I}+q_{i}^{X}\},$
-for $i\in\{1,2,\dots,n\}$.
-Note that $\{q_{i}^{X}+q_{i}^{Y},q_{i}^{I}+q_{i}^{Z}\}$ is a binary distribution indicating the likelihood of an $X$ error occuring at the $i$-th qubit and similarly for $\{q_{i}^{X}+q_{i}^{Y},q_{i}^{I}+q_{i}^{Z}\}$.
+return “Failure", $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, and $\bm{\eta}$. $\triangleright$ BP fails.
 Now we define a reliability order for the bits in ${\mathbf{e}}=\left[\begin{array}[]{c|c}{\mathbf{e}}^{X}&{\mathbf{e}}^{Z}\end{array}\right]\in\{0,1\}^{2n}$ based on the outputs of BP4, using the aforementioned hard and soft reliability functions.
@@
 This definition means that the hard-decision history reliability is used to rank the error coordinates, with the soft reliability serving as a tie-breaker.
-Let ${\mathbf{E}}$ be the hard-decision output from BP4 after $T$ iterations.
-If ${\mathbf{E}}$ does not match the syndrome, OSD will be applied.
-We now present our OSD algorithm based on the reliability orders as in Algorithm 2, which is also referred to as OSD4-00.
-If the bits in the reliable subset are correct, the output of OSD4-00 is the required solution.
+Let $\tilde{{\mathbf{e}}}$ be the hard-decision output from BP4 after ${\mathrm{T}}$ iterations.
+If $\tilde{{\mathbf{e}}}$ does not match the syndrome, OSD will be applied.
+We now present our OSD algorithm based on the reliability orders as in Algorithm 2, which is referred to as OSD4-00.
 Algorithm 2 OSD4-0
-Input: check matrix ${\mathbf{S}}\in\{0,1\}^{m\times 2n}$ of rank $n-k$, syndrome ${\mathbf{z}}\in\{0,1\}^{m\times 1}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, hard-decision vector ${\mathbf{E}}\in\{0,1\}^{2n}$, and belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$.
+Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$ of rank $n-k$, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, hard-decision vector $\tilde{{\mathbf{e}}}\in\{0,1\}^{2n}$, and belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$.
 Output: A valid error estimate $\hat{{\mathbf{e}}}\in\{0,1\}^{2n}$.
 Steps:
-1:Sort the error bits in ascending order of reliability. Construct a column permutation function $\pi$ corresponding to the sorting order. Calculate $\pi({\mathbf{E}})$ and $\pi({\mathbf{S}}\Lambda)$.
-2:Perform Gaussian elimination on $\pi({\mathbf{S}}\Lambda)$ and let the output be denoted by $R(\pi({\mathbf{S}}\Lambda))$, where $R$ is the corresponding row operations in the Gaussian elimination.
+1:Sort the error bits in ascending order of reliability.
+Construct a column permutation function $\pi$ corresponding to the sorting order.
+Calculate $\pi(\tilde{{\mathbf{e}}})$ and $\pi({\mathbf{H}}\Lambda)$.
+2:Perform Gaussian elimination on $\pi({\mathbf{H}}\Lambda)$ and let the output be denoted by $R(\pi({\mathbf{H}}\Lambda))$, where $R$ represents the corresponding row operations in the Gaussian elimination.
 If necessary, apply a column permutation function $\mu$ to ensure that the first $n-k$ columns of the resulting matrix are linearly independent.
 Thus we have
-$\mu(R(\pi({\mathbf{S}}\Lambda)))=\left[\begin{array}[]{cc}{\mathbf{I}}_{n-k}&{\mathbf{A}}\\ \mathbf{0}&\mathbf{0}\end{array}\right],$ (5)
+$\mu(R(\pi({\mathbf{H}}\Lambda)))=\left[\begin{array}[]{cc}{\mathbf{I}}_{n-k}&{\mathbf{A}}\\ \mathbf{0}&\mathbf{0}\end{array}\right],$ (7)
 for some ${\mathbf{A}}\in\{0,1\}^{(n-k)\times(n+k)}$, where the last $m-(n-k)$ rows are all zeros.
-3:Let ${\mathbf{z}}^{\prime}=R({\mathbf{z}})$ so that the parity-check condition (2) becomes ${\mathbf{z}}^{\prime}=\mu(R(\pi({\mathbf{S}}\Lambda)))(\mu(\pi(\hat{{\mathbf{E}}})))^{\top}$.
-Remove the last $m-(n-k)$ entries of ${\mathbf{z}}^{\prime}$.
-4:Suppose $\mu(\pi({\mathbf{E}}))=\left[{\mathbf{E}}^{U}~{}~{}{\mathbf{E}}^{R}\right]$, where ${\mathbf{E}}^{U}$ represents the independent and unreliable subset of $n-k$ bits and ${\mathbf{E}}^{R}$ is the reliable subset of $n+k$ bits.
+3:Let ${\mathbf{s}}^{\prime}=R({\mathbf{s}})$ so that the parity-check condition (2) becomes
+${\mathbf{s}}^{\prime}=\mu(R(\pi({\mathbf{H}}\Lambda)))(\mu(\pi({\mathbf{e}})))^{\top}.$ (8)
+Remove the last $m-(n-k)$ entries of ${\mathbf{s}}^{\prime}$.
+4:Let $\mu(\pi(\tilde{{\mathbf{e}}}))=\left[\tilde{{\mathbf{e}}}^{U}~{}~{}\tilde{{\mathbf{e}}}^{R}\right]$, where $\tilde{{\mathbf{e}}}^{U}$ represents the independent and unreliable subset of $n-k$ bits and $\tilde{{\mathbf{e}}}^{R}$ is the reliable subset of $n+k$ bits.
+Use the reliable subset $\tilde{{\mathbf{e}}}^{R}$ and the modified parity-check condition 8 to select feasible unreliable bits
+$\tilde{{\mathbf{e}}}^{U}={{\mathbf{s}}^{\prime}}^{\top}\oplus{\tilde{{\mathbf{e}}}^{R}}{\mathbf{A}}^{\top}.$ (9)
 return
-$\hat{{\mathbf{E}}}\triangleq\pi^{-1}\left(\mu^{-1}\left(\left[{{\mathbf{z}}^{\prime}}^{\top}\oplus{{\mathbf{E}}^{R}}{\mathbf{A}}^{\top}\qquad{\mathbf{E}}^{R}\right]\right)\right).$ (6)
-For comparison with conventional OSD procedures that do not incorporate the history of BP hard-decision outputs across iterations, we also define a reliability order that accounts for only the soft reliability functions of BP4.
+$\hat{{\mathbf{e}}}\triangleq\pi^{-1}\left(\mu^{-1}\left(\left[{{\mathbf{s}}^{\prime}}^{\top}\oplus{\tilde{{\mathbf{e}}}^{R}}{\mathbf{A}}^{\top}\qquad\tilde{{\mathbf{e}}}^{R}\right]\right)\right).$ (10)
+We also define a reliability order that accounts for only the soft reliability functions of BP4.
 Definition 5.
 Error bit ${\mathbf{e}}_{i}^{a}$ is said to be more reliable than error bit ${\mathbf{e}}_{j}^{b}$ in the marginal distribution if $\phi^{a}(i)\geq\phi^{b}(j)$.
-An OSD algorithm that follows a reliability order depending only on the belief distributions from the last BP iteration is referred to as mOSD. The OSD based on Definition 5 is referred to as mOSD4.
-Note that the reliability metric in Definition 5 is distinct from those in [42, 43]. Other alternative metrics can also be considered, such as
+The OSD based on Definition 5 is referred to as mOSD4.
+Other alternative metrics can also be considered, such as
 […]
@@
 Lemma 6.
-Given an OSD4-0 output vector, flipping one of its reliable bits to generate an error candidate takes time ${\cal O}(n)$.
-Similarly, generating an error candidate corresponding to a child node from its parent node also takes ${\cal O}(n)$ time.
+Given an OSD4-0 output vector, flipping one of its reliable bits to generate an error candidate takes time $O(n)$.
+Similarly, generating an error candidate corresponding to a child node from its parent node also takes $O(n)$ time.
 Proof.
-Consider the vector $\Big{[}{\mathbf{z}}^{\prime\top}\oplus{\mathbf{E}}^{R}{\mathbf{A}}^{\top}\,\Big{|}{\mathbf{E}}_{1}^{R}\,{\mathbf{E}}_{2}^{R}\,\cdots\,{\mathbf{E}}_{n+k}^{R}\Big{]}$ generated by $\operatorname{OSD}_{4}$-0 in (6), where ${\mathbf{E}}_{j}^{R}$ are reliable bits.
+Consider the vector $\Big{[}{\mathbf{s}}^{\prime\top}\oplus{\mathbf{e}}^{R}{\mathbf{A}}^{\top}\,\Big{|}\tilde{{\mathbf{e}}}_{1}^{R}\,\tilde{{\mathbf{e}}}_{2}^{R}\,\cdots\,\tilde{{\mathbf{e}}}_{n+k}^{R}\Big{]}$ generated by $\operatorname{OSD}_{4}$-0 in (10), where $\tilde{{\mathbf{e}}}_{j}^{R}$ are reliable bits.
 Suppose that the matrix ${\mathbf{A}}$ in (7) has columns ${\mathbf{A}}_{1},{\mathbf{A}}_{2},\dots,{\mathbf{A}}_{n+k}\in\{0,1\}^{n-k}$.
-Then flipping the bit ${\mathbf{E}}_{i}^{R}$ generates the vector
-$\left[{\mathbf{z}}^{\prime\top}\oplus{\mathbf{E}}^{R}{\mathbf{A}}^{\top}\oplus{\mathbf{A}}_{i}^{\top}\middle|{\mathbf{E}}_{1}^{R}\,{\mathbf{E}}_{2}^{R}\,\cdots\Big{(}1\oplus{\mathbf{E}}_{i}^{R}\Big{)}\cdots\,{\mathbf{E}}_{n+k}^{R}\right],$ (9)
-which is a valid error candidate after permutations. It can be checked that this calculation takes time $\mathcal{O}(n)$. ∎
+Then according to (9), flipping the bit $\tilde{{\mathbf{e}}}_{i}^{R}$ corresponds to adding the OSD4-0 output by the vector
+$\begin{bmatrix}{\mathbf{A}}_{j}^{\top}&0\cdots 0&1&0\cdots 0\end{bmatrix}\in\{0,1\}^{2n},$ (13)
+where ${\mathbf{A}}_{j}$ is the $j$-th column of ${\mathbf{A}}$, and generating the vector
+$\left[{\mathbf{s}}^{\prime\top}\oplus\tilde{{\mathbf{e}}}^{R}{\mathbf{A}}^{\top}\oplus{\mathbf{A}}_{i}^{\top}\middle|\tilde{{\mathbf{e}}}_{1}^{R}\,\tilde{{\mathbf{e}}}_{2}^{R}\,\cdots\Big{(}1\oplus\tilde{{\mathbf{e}}}_{i}^{R}\Big{)}\cdots\,\tilde{{\mathbf{e}}}_{n+k}^{R}\right],$ (14)
+which is a valid error candidate after permutations. This calculation takes $O(n)$ time. ∎
 To efficiently navigate through these possibilities, we can utilize the technique described in Lemma 6 combining with the DFS algorithm.
 Then all $\sum_{i=0}^{w}\binom{n+k}{i}$ candidates can be recursively generated.
-We prioritize flipping the least reliable bits first.
-If an error with a lower Hamming weight is encountered, we update the solution accordingly.
-However, if the error has the same Hamming weight as the current one, we retain the previous solution since the bit flipped in the previous step is deemed less reliable.
 00001000110010101001010001100101001000110001
@@
 We analyze the complexity of OSD based on 00-ADD in the following.
-The complexity of OSD4-0 is dominated by the Gaussian elimination step, which is then ${\cal O}(2n\times m^{2})={\cal O}(n^{3})$, assuming $m={\cal O}(n)$.
-For OSD4-$w$ with $w>0$, the total number of possible flips is $\sum_{j=0}^{w}\binom{n+k}{j}$, which can be expressed as ${\cal O}((n+k)^{w})$ or simply ${\cal O}(n^{w})$.
-In Lemma 6, when we flip a bit in the reliable part, it requires only $\mathcal{O}(n)$ calculations to find an error candidate.
-Additionally, by running through $\sum_{j=0}^{w}\binom{n+k}{j}$ possibilities using the DFS algorithm, the OSD4-$w$ algorithm has a complexity of $\mathcal{O}(n^{3}+n^{w+1})$, which simplifies to $\mathcal{O}(n^{3})$ when $w\leq 2$.
-The above OSD complexity is only required when BP fails to converge.
-BP has a worst-case time complexity of $O(n\gamma T)$ using a sequential schedule, where $n$ is the number of error variables, $\gamma$ is the mean column weight of the check matrix, and $T$ is the maximum number of iterations.
-However, BP typically converges quickly, within $O(\log N)$ iterations, for sparse matrices.
-This is also the case for MBP4 as demonstrated in Table VI in the simulation section.
-Thus MBP4 has a nearly linear complexity in $n$ from simulations and has a high probability of convergence for many spare-graph quantum codes [12, 13].
+The complexity of OSD4-0 is dominated by the Gaussian elimination step, which is then $O(2nm^{2})=O(n^{3})$, assuming $m=O(n)$.
+For OSD4-$w$ with $w>0$, the total number of possible flips is $\sum_{j=0}^{w}\binom{n+k}{j}$, which can be expressed as $O((n+k)^{w})$ or simply $O(n^{w})$.
+In Lemma 6, when we flip a bit in the reliable part, it requires only $O(n)$ calculations to find an error candidate.
+Additionally, by running through $\sum_{j=0}^{w}\binom{n+k}{j}$ possibilities using the DFS algorithm, the OSD4-$w$ algorithm has a complexity of $O(n^{3}+n^{w+1})$, which simplifies to $O(n^{3})$ when $w\leq 2$.
 Algorithm 3 HRSR
-Input: check matrix ${\mathbf{S}}\in\{0,1\}^{m\times 2n}$ of rank $n-k$, syndrome ${\mathbf{z}}\in\{0,1\}^{m\times 1}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, hard-decision vector ${\mathbf{E}}\in\{0,1\}^{2n}$, belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, and soft reliability threshold $\theta$.
-Output: shortened variables $\tilde{{\mathbf{z}}}$, $\tilde{{\mathbf{S}}}$, ${\mathbf{E}}^{H}$, and $\sigma$.
+Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$ of rank $n-k$, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, hard-decision vector ${\mathbf{e}}\in\{0,1\}^{2n}$, belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, and soft reliability threshold $\theta$.
+Output: shortened variables $\tilde{{\mathbf{s}}}$, $\tilde{{\mathbf{H}}}$, ${\mathbf{e}}^{H}$, and $\sigma$.
 Steps:
@@
 Let $\sigma(\hat{{\mathbf{e}}})=\begin{bmatrix}\hat{{\mathbf{e}}}^{\prime}&{\mathbf{e}}^{H}\end{bmatrix}$ be a permuted error estimate, which contains the highly reliable bits in ${\mathbf{e}}^{H}$.
-3:Let $\tau$ be a row permutation such that $\tau(\sigma({\mathbf{S}}\Lambda))=\begin{bmatrix}\tilde{{\mathbf{S}}}&{\mathbf{B}}\\ \mathbf{0}&{\mathbf{C}}\\ \end{bmatrix}$ for $\tilde{{\mathbf{S}}}\in\{0,1\}^{m^{\prime}\times(2n-v)}$ with $m^{\prime}\leq m$.
+3:Let $\tau$ be a row permutation such that $\tau(\sigma({\mathbf{H}}\Lambda))=\begin{bmatrix}\tilde{{\mathbf{H}}}&{\mathbf{B}}\\ \mathbf{0}&{\mathbf{C}}\\ \end{bmatrix}$ for $\tilde{{\mathbf{H}}}\in\{0,1\}^{m^{\prime}\times(2n-v)}$ with $m^{\prime}\leq m$.
 The $m-m^{\prime}$ rows $\begin{bmatrix}\mathbf{0}&{\mathbf{C}}\end{bmatrix}$ correspond to parity checks related only to the highly reliable bits.
-Assume that $\tau({\mathbf{z}})=\begin{bmatrix}{\mathbf{z}}^{\prime\prime}\\ {\mathbf{z}}^{H}\end{bmatrix}$, where ${\mathbf{z}}^{\prime\prime}\in\{0,1\}^{m^{\prime}\times 1}$, and ${\mathbf{z}}^{H}\in\{0,1\}^{(m-m^{\prime})\times 1}$ are the syndrome bits corresponding to $\begin{bmatrix}\mathbf{0}&{\mathbf{C}}\end{bmatrix}$.
+Assume that $\tau({\mathbf{s}})=\begin{bmatrix}{\mathbf{s}}^{\prime\prime}\\ {\mathbf{s}}^{H}\end{bmatrix}$, where ${\mathbf{s}}^{\prime\prime}\in\{0,1\}^{m^{\prime}\times 1}$, and ${\mathbf{s}}^{H}\in\{0,1\}^{(m-m^{\prime})\times 1}$ are the syndrome bits corresponding to $\begin{bmatrix}\mathbf{0}&{\mathbf{C}}\end{bmatrix}$.
 4:Verify the parity check condition related to the highly reliable bits.
-if ${\mathbf{z}}^{H}\neq{\mathbf{C}}\big{(}{\mathbf{E}}^{H}\big{)}^{\top}$ then return failure;
+if ${\mathbf{s}}^{H}\neq{\mathbf{C}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}$ then return “Failure";
 else we have a new parity-check condition
-$\tilde{{\mathbf{z}}}=$ $\tilde{{\mathbf{S}}}\big{(}\hat{{\mathbf{E}}}^{\prime}\big{)}^{\top},$ (10)
-     where $\tilde{{\mathbf{z}}}={\mathbf{z}}^{\prime\prime}-{\mathbf{B}}\big{(}{\mathbf{E}}^{H}\big{)}^{\top}$.
-     if (10) is solvable then return $\tilde{{\mathbf{z}}}$, $\tilde{{\mathbf{S}}}$, ${\mathbf{E}}^{H}$, and $\sigma$;
-     else return failure.
+$\tilde{{\mathbf{s}}}=$ $\tilde{{\mathbf{H}}}\big{(}\hat{{\mathbf{e}}}^{\prime}\big{)}^{\top},$ (15)
+     where $\tilde{{\mathbf{s}}}={\mathbf{s}}^{\prime\prime}-{\mathbf{B}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}$.
+     if (15) is solvable then return $\tilde{{\mathbf{s}}}$, $\tilde{{\mathbf{H}}}$, ${\mathbf{e}}^{H}$, and $\sigma$;
+     else return “Failure".
      end if
 end if
-IV Efficient OSD with highly reliable subset reduction
+IV Highly reliable subset reduction
 As we increase the order $w$ of OSD, we can expect an improvement in the decoder’s performance.
@@
 The number of binary variables in the reduced linear system is referred to as the effective length for OSD, which corresponds to the length of the remaining binary error variable vector.
-We note that OSD is applied when BP fails to find a valid error that matches the syndrome.
-During the $T$ iterations of message passing in BP, some coordinates become relatively more reliable than others.
 We have the following definition of highly reliable coordinates.
 Definition 7.
-An error bit ${\mathbf{E}}_{i}^{a}$, where $i\in{1,\dots,n}$ and $a\in{X,Z}$, is considered highly reliable if its hard-decision reliability equals $T$ or $T+1$, and its soft reliability $\phi^{a}(i)$ satisfies $\phi^{a}(i)\geq\theta$, where $\theta$ is a predetermined reliability threshold.
-An error variable has hard reliability $T$ if BP has a constant output at all iterations; a hard reliability of $T+1$ means the corresponding error variable is the identity during the BP process after being initialized to the identity.
+An error bit ${\mathbf{e}}_{i}^{a}$, where $i\in{1,\dots,n}$ and $a\in{X,Z}$, is considered highly reliable if its hard-decision reliability equals ${\mathrm{T}}$ or ${\mathrm{T}}+1$, and its soft reliability $\phi^{a}(i)$ satisfies $\phi^{a}(i)\geq\theta$, where $\theta$ is a predetermined reliability threshold.
+An error variable has hard reliability ${\mathrm{T}}$ if BP has a constant output at all iterations; a hard reliability of ${\mathrm{T}}+1$ means the corresponding error variable is the identity during the BP process after being initialized to the identity.
 The soft reliability threshold $\theta$ is chosen close to one to ensure that a highly reliable coordinate not only exhibits an almost constant hard-decision history but also has a high soft reliability value.
 These highly reliable variables are crucial in the OSD process. Identifying these variable nodes allows us to greatly reduce the computation burden of OSD.
-In topological codes, such as toric or surface codes, an isolated Pauli error refers to a single-qubit Pauli error (including identity) with no neighboring nonidentity Pauli error within a radius of two qubits.
-These isolated Pauli errors can be decoded using BP4 with high confidence, contributing to the highly reliable parts in the output of BP.
 These highly reliable parts provide outcomes that can be trusted and will remain unchanged throughout the OSD4 procedure with high probability.
@@
 Our highly reliable subset reduction (HRSR) algorithm is given in Algorithm 3.
-Equation (10) follows from the parity-check condition (2) as
-$\begin{bmatrix}{\mathbf{z}}^{\prime\prime}\\ {\mathbf{z}}^{H}\end{bmatrix}=$ $\tau(\sigma({\mathbf{S}}\Lambda))\begin{bmatrix}\hat{{\mathbf{E}}}^{\prime}&{\mathbf{E}}^{H}\end{bmatrix}^{\top}$
-$=$ $\begin{bmatrix}\tilde{{\mathbf{S}}}\big{(}\hat{{\mathbf{E}}}^{\prime}\big{)}^{\top}+{\mathbf{B}}\big{(}{\mathbf{E}}^{H}\big{)}^{\top}&{\mathbf{C}}\big{(}{\mathbf{E}}^{H}\big{)}^{\top}\end{bmatrix}$
-and using the conditions ${\mathbf{z}}^{H}={\mathbf{C}}\big{(}{\mathbf{E}}^{H}\big{)}^{\top}$ and $\tilde{{\mathbf{z}}}={\mathbf{z}}^{\prime\prime}-{\mathbf{B}}\big{(}{\mathbf{E}}^{H}\big{)}^{\top}$.
+Equation (15) follows from the parity-check condition (2) with
+$\begin{bmatrix}{\mathbf{s}}^{\prime\prime}\\ {\mathbf{s}}^{H}\end{bmatrix}=$ $\tau(\sigma({\mathbf{H}}\Lambda))\begin{bmatrix}\hat{{\mathbf{e}}}^{\prime}&{\mathbf{e}}^{H}\end{bmatrix}^{\top}=$ $\begin{bmatrix}\tilde{{\mathbf{H}}}\big{(}\hat{{\mathbf{e}}}^{\prime}\big{)}^{\top}+{\mathbf{B}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}\\ {\mathbf{C}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}\end{bmatrix}$
+and the conditions ${\mathbf{s}}^{H}={\mathbf{C}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}$ and $\tilde{{\mathbf{s}}}={\mathbf{s}}^{\prime\prime}-{\mathbf{B}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}$.
 If HRSR returns a reduced linear system as in (15), we then use OSD4-$w$ to solve it with hard-decision BP output ${\mathbf{e}}^{\prime}$ and corresponding reliabilities $\bm{\eta}$, $\phi^{X}$, and $\phi^{Z}$.
@@
 Reducing the problem size may cause the reduced check matrix to lose the ability to represent all valid codewords if its dimension becomes too small.
-To address this, we set the soft reliability threshold $\theta$ very close to one to ensure that highly reliable bits are accurate.
-Furthermore, in Step 4 of Algorithm 3, we verify that the removed highly reliable bits align with the corresponding syndrome bits.
+To address this, in Step 4 of Algorithm 3, we verify that the removed highly reliable bits align with the corresponding syndrome bits.
 If they do not, the algorithm reports a failure.
@@
 As shown in Lemma 6, a new candidate can be generated from an error candidate by flipping a bit in the reliable subset and adding a certain vector to the unreliable subset.
-This new candidate must belong to the same error coset, and thus the bit flipping in the reliable subset corresponds to an operator in the normalizer group.
-We have the following degenerate error candidate pruning conditions.
-Lemma 8.
-Consider an $[[n,k,d]]$ stabilizer code with a check matrix given by $\begin{bmatrix}{\mathbf{I}}_{n-k}&{\mathbf{A}}\end{bmatrix}$ after Gaussian elimination and a column permutation as in Equation (5), where ${\mathbf{A}}\in\{0,1\}^{(n-k)\times(n+k)}$ are associated with the reliable bits of OSD4-0 output.
-1) When using $n$-ADD, only the reliable bits corresponding to nontrivial logical operators need to be considered in OSD4-$w$.
-2) If each error corresponding to flipping a reliable bit of OSD4-0 output is a stabilizer, then OSD4-$w$ is equivalent to OSD4-0 for any $w$.
+This corresponds to adding a logical operator in the normalizer group to the error candidate.
+By examining these logical operators, we have the following degenerate error candidate pruning conditions.
+Theorem 8.
+Consider an $[[n,k,d]]$ stabilizer code with a check matrix given by $\begin{bmatrix}{\mathbf{I}}_{n-k}&{\mathbf{A}}\end{bmatrix}$ after Gaussian elimination and a column permutation as in (7), where ${\mathbf{A}}\in\{0,1\}^{(n-k)\times(n+k)}$ are associated with the reliable bits of OSD4-0 output.
+  1. 1.
+    When using $n$-ADD, only the reliable bits corresponding to nontrivial logical operators need to be considered in OSD4-$w$.
+  2. 2.
+    If each logical operator induced by flipping a reliable bit of OSD4-0 output is a stabilizer, then OSD4-$w$ is equivalent to OSD4-0 for any $w$.
 Proof.
-1)
-In optimal degenerate decoding ($n$-ADD), different stabilizer cosets are compared. Therefore, multiplying a stabilizer by an error candidate does not create a new nontrivial candidate.
+1) In $n$-ADD, different stabilizer cosets are compared. Therefore, multiplying a stabilizer by an error candidate does not create a new nontrivial candidate.
 2) The statement is straightforward, as flipping these reliable bits results in only degenerate errors to the OSD4-0 output.
-Exploiting Lemma 8-1) with $n$-ADD makes OSD4-$w$ more effective.
-By incorporating with the HRSR algorithm, we can work on a shortened check matrix $\tilde{{\mathbf{S}}}$ and higher order of OSD can be executed.
+Exploiting Theorem 8-1) with $n$-ADD makes OSD4-$w$ more effective.
+By incorporating with the HRSR algorithm, we can work on a shortened check matrix $\tilde{{\mathbf{H}}}$ and higher order of OSD can be executed.
 Consequently, (17) can be improved to $w=\operatorname*{arg\,max}_{x}\sum_{i=0}^{x}\binom{u}{i}\leq\Gamma,$ where $u\leq n+k-v+(m-m^{\prime})$ is the number of columns in the shortened check matrix corresponding to nontrivial logical operators.
-For checking whether flipping a bit is nontrivial can be done by checking whether its induced error anticommutes with the permuted logical operators.
-If an error commutes with all logical operators, it belongs to the stabilizer group.
-This aligns with the standard procedure in QEC simulations for calculating the logical error rate.
-We note that the computational complexity of verifying such a bit-flipping event is $O((n-k)k)$.
-Consequently, checking all the columns of $A$ has a computational complexity of roughly the same order as Gaussian elimination.
+Proposition 9.
+Checking whether a logical operator induced by flipping a reliable bit of the OSD4-0 output is a stabilizer can be done in $O((n-k)k)$ time.
+Proof.
+Consider an $[[n,k,d]]$ stabilizer code with a $m\times{2n}$ check matrix ${\mathbf{H}}$ and a $2k\times 2n$ binary logical matrix ${\mathbf{L}}$.
+During the process of OSD4-0, ${\mathbf{H}}$ leads to $\begin{bmatrix}{\mathbf{I}}_{n-k}&{\mathbf{A}}\end{bmatrix}$ as in (7) using column permutations $\pi,\mu$ and row operations $R$, where ${\mathbf{A}}\in\{0,1\}^{(n-k)\times(n+k)}$ are associated with the reliable bits of OSD4-0 output.
+By (13), flipping the $j$-th reliable bit corresponds to adding the vector ${\mathbf{g}}=\begin{bmatrix}{\mathbf{A}}_{j}^{\top}&0\cdots 0&1&0\cdots 0\end{bmatrix}\in\{0,1\}^{2n}$, where ${\mathbf{A}}_{j}$ is the $j$-th column of ${\mathbf{A}}$.
+The vector ${\mathbf{g}}$ corresponds to a stabilizer if $\mu(\pi({\mathbf{L}}\Lambda)){\mathbf{g}}^{\top}=\mathbf{0}$, and otherwise, it represents a nontrivial logical operator.
+Since ${\mathbf{A}}_{j}$ is of dimension $n-k$, this computation takes $O((n-k)k)$ time.
+The computational complexity of verifying such a bit-flipping event is $O((n-k)k)$, which implies that checking all the columns of $A$ has a computational complexity of roughly the same order as Gaussian elimination.
 This approach is both effective and relatively efficient for implementing high-order OSD.
 However, the complexity of $n$-ADD is not practically feasible.
-We propose to use $\delta$-ADD by choosing $\delta$ as the highest weight of a set of low-weight stabilizer generators.
+We may use $\delta$-ADD by choosing $\delta$ as the highest weight of a set of low-weight stabilizer generators.
 For the case of a rotated surface code [25], the stabilizer generators are of weight $2$ or $4$, and the next lowest stabilizer weight is $6$.
@@
 Then, $4$-ADD outputs:
-$\hat{{\mathbf{E}}}=\operatorname*{arg\,max}_{{\mathbf{E}}}\sum_{i=0}^{n-k}\Pr\big{\{}{\mathbf{E}}+{\mathbf{S}}_{i}\big{\}},$ (13)
+$\hat{{\mathbf{e}}}=\operatorname*{arg\,max}_{{\mathbf{e}}}\sum_{i=0}^{n-k}\Pr\big{\{}{\mathbf{e}}+{\mathbf{H}}_{i}\big{\}},$ (18)
 where $\mathbf{S}_{0}=I$ and $\{\mathbf{S}_{i}\}_{i=1}^{n-k}$ is a set of stabilizer generators for the rotated surface code.
@@
 To achieve optimal execution complexity, 00-ADD is often used.
-Lemma 8 suggests focusing on the columns of $A$ corresponding to nontrivial logical operators when $n$-ADD is used.
+Theorem 8 suggests focusing on the columns of $A$ corresponding to nontrivial logical operators when $n$-ADD is used.
 However, in the scenario of 00-ADD, a stabilizer may influence the output of $\text{OSD}_{4}$-$w$ since multiplying a stabilizer can reduce the weight of the error, potentially providing a better candidate in 00-ADD, even though the error coset remains unchanged.
-Thus, $\text{OSD}_{4}$-$w$ cannot simply flip bits corresponding to those nontrivial logical operators as in the case of Lemma 8-1).
-The following provides a sufficient condition to determine whether flipping a reliable bit corresponds to a stabilizer.
-Lemma 9.
-Consider an $[[n,k,d]]$ stabilizer code with a check matrix given by $\begin{bmatrix}{\mathbf{I}}_{n-k}&{\mathbf{A}}\end{bmatrix}$ after Gaussian elimination and a column permutation as in Equation (5), where ${\mathbf{A}}\in\{0,1\}^{(n-k)\times(n+k)}$ are associated with the reliable bits of OSD4-0 output.
-If the weight of the $j$-th column of ${\mathbf{A}}$ is less than $d-1$, then flipping the $j$-th reliable bit of an error candidate corresponds to multiplying the error candidate by a stabilizer.
-Proof.
-Suppose that ${\mathbf{E}}=\begin{bmatrix}{\mathbf{F}}\ \Big{|}&{\mathbf{E}}_{1}^{R}&{\mathbf{E}}_{2}^{R}&\cdots&{\mathbf{E}}_{n+k}^{R}\end{bmatrix}$ is a valid error candidate, where ${\mathbf{E}}_{i}\in\{0,1\}$ for $i=1,\dots,n$ and ${\mathbf{F}}\in\{0,1\}^{n-k}$.
-By (9), flipping the bit ${\mathbf{E}}_{j}^{T}$ corresponds to adding ${\mathbf{E}}$ by the vector $\begin{bmatrix}{\mathbf{A}}_{j}^{\top}&0\cdots 0&1&0\cdots 0\end{bmatrix}\in\{0,1\}^{2n}$, where ${\mathbf{A}}_{j}$ is the $j$-th column of ${\mathbf{A}}$.
-This vector corresponds to a Pauli operator of weight at most $d-1$ and, therefore, is a stabilizer since the minimum distance of the code is $d$.
-By Lemma 8-2) and Corollary 9, we establish the following operational degenerate error candidate pruning condition.
+Thus, $\text{OSD}_{4}$-$w$ cannot simply flip bits corresponding to those nontrivial logical operators as in the case of Theorem 8-1).
+However, Theorem 8-2) still works and we may use Proposition 9 to check this condition.
+In the following we provide a weaker sufficient but operationally more efficient condition, for degenerate error candidate pruning.
 Corollary 10.
 Let $\tilde{S}\in\{0,1\}^{m^{\prime}\times(2n-v)}$ be a shortened check matrix by the HRSR algorithm.
-Suppose that $\tilde{S}$ can be transformed into $\begin{bmatrix}{\mathbf{I}}&\tilde{A}\end{bmatrix}$ after Gaussian elimination and a column permutation, where $\tilde{A}$ is associated with the reliable bits.
+Suppose that $\tilde{S}$ can be transformed into $\begin{bmatrix}{\mathbf{I}}&\tilde{{\mathbf{A}}}\end{bmatrix}$ after Gaussian elimination and a column permutation, where $\tilde{{\mathbf{A}}}$ is associated with the reliable bits.
 If all the columns of $\tilde{{\mathbf{A}}}$ are of weight less than $d-1$, then OSD4-$w$ on the corresponding reduced linear system is equivalent to OSD4-0 for any $w$.
@@
 We combine all the aforementioned techniques into the approximate degenerate OSD (ADOSD) algorithm, as detailed in Algorithm 4, using 00-ADD.
-It is important to note that if decoding performance is of higher priority, $\delta$-ADD can be utilized, and Lemma 9 can be integrated into Algorithm 4.
+Note that if decoding performance is of higher priority, $\delta$-ADD can be utilized, and Proposition 9 can be integrated into Algorithm 4.
 Algorithm 4 ADOSD4
-Input: check matrix ${\mathbf{S}}\in\{0,1\}^{m\times 2n}$ of rank $n-k$, syndrome ${\mathbf{z}}\in\{0,1\}^{m\times 1}$, BP output distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, soft reliability threshold $\theta$, hard-decision vector ${\mathbf{E}}\in\{0,1\}^{2n}$, maximum number of OSD flips $\Gamma$, and backup OSD order $w$.
+Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$ of rank $n-k$, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, BP output distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, soft reliability threshold $\theta$, hard-decision vector ${\mathbf{e}}\in\{0,1\}^{2n}$, maximum number of OSD flips $\Gamma$, and backup OSD order $w$.
 Output: A valid error estimate $\hat{{\mathbf{e}}}\in\{0,1\}^{2n}$.
@@
      use OSD4-$w$ to find an error estimate $\hat{{\mathbf{e}}}\in\{0,1\}^{2n}$ for the original problem, and return $\hat{{\mathbf{e}}}$;
-else HRSR returns ${\mathbf{z}}^{\prime}$, $\tilde{{\mathbf{S}}}$, ${\mathbf{E}}^{H}$, and $\sigma$.
-     if all the columns of $\tilde{{\mathbf{S}}}$ in reduced row echelon form have weight less than $d-1$then $w\leftarrow 0$;
+else HRSR returns ${\mathbf{s}}^{\prime}$, $\tilde{{\mathbf{H}}}$, ${\mathbf{e}}^{H}$, and $\sigma$.
+     if all the columns of $\tilde{{\mathbf{H}}}$ in reduced row echelon form have weight less than $d-1$then $w\leftarrow 0$;
      else determine $w$ by (17).
      end if
-     Use OSD4-$w$ on ${\mathbf{z}}^{\prime}$, $\tilde{S}$, and the given data to find an error estimate $\hat{{\mathbf{E}}}^{\prime}\in\{0,1\}^{2n-v}$, and return $\sigma^{-1}\left(\begin{bmatrix}\hat{{\mathbf{E}}}^{\prime}&{\mathbf{E}}^{H}\end{bmatrix}\right)$.
+     Use OSD4-$w$ on ${\mathbf{s}}^{\prime}$, $\tilde{S}$, and the given data to find an error estimate $\hat{{\mathbf{e}}}^{\prime}\in\{0,1\}^{2n-v}$, and return $\sigma^{-1}\left(\begin{bmatrix}\hat{{\mathbf{e}}}^{\prime}&{\mathbf{e}}^{H}\end{bmatrix}\right)$.
 end if
-Remark 11.
-We can incorporate the idea of checking the commutation relations with permuted logical operators for column-induced errors of weight at least $d$ into Algorithm 4.
-For instance, we first identify columns with weights less than $d-1$ and then verify the commutation relations between the permuted logical operators and the errors induced by the remaining columns.
-As will be shown in Table VII, using the strategy outlined in Corollary 10, OSD-00 is applied in the majority of trials when the error rate is low.
-decoder source note
-MBP4+ADOSD4 this paper based on Algorithm 4
-MBP4+OSD4-$w$ this paper order-$w$ OSD4 with reliability metric from Definition 4
-MBP4+mOSD4-$w$ this paper order-$w$ OSD4 with reliability metric from Definition 5
-AMBP4 [31] No postprocessing. Can act as a predecoder for ADOSD4 and OSD4
-BP+mOSD-($w$,$\lambda$) [42] binary BP+ order-w OSD4 with additional flips on $\lambda$ positions
-BP-OSD-CS($\lambda$) [43] binary BP-OSD-0 with a greed search on (${\lambda\choose 2}$ positions
-BP+GDG [54] binary BP with guided decimal guessing based on a decision tree
-TABLE I: Various decoders compared in this paper.
 VI Simulation results
-We simulate the performance of various BP+OSD schemes on LCS codes [59], GHP codes [42, 43], BB codes [55], and various 2D topological codes, including toric, rotated surface, color codes and twisted XZZX codes [22, 23, 24, 25, 26, 58, 28] (see [64, Table I]).
-The decoding performance of a quantum code ${\cal C}({\cal S})$ and its decoder is assessed through the logical error rate (LER).
-An error $E\in{\cal G}_{n}$ is classified as a logical error if it is a nontrivial logical operator.
-The simulation parameters are as follows: The maximum number of BP iterations, $T$, is set to $100$ unless otherwise specified.
-In most cases, a much smaller value of $T$ suffices for good decoding performance due to the rapid convergence of MBP4[31], which typically completes in just a few iterations for codes with lengths under one thousand.
-For instance, as shown in Table V, When MBP4 successfully decodes the $[[225,1,15]]$ rotated surface code, it requires an average of only $2.41$ iterations at a depolarizing rate of $\epsilon=3.3\%$.
+We simulate the performance of the MBP4+ADOSD4 scheme on LCS codes [60], GHP codes [43, 44], BB codes [56], and various 2D topological codes, including toric, rotated surface, color codes and twisted XZZX codes [23, 24, 25, 26, 27, 59, 29] (see [64, Table I]).
+We use the MBP4 decoder introduced in [32] as a predecoder for OSD.
+MBP4 features a parameter $\alpha$ that controls the step size of message updates during BP iterations: when $\alpha<1$, the step size is enlarged; when $\alpha>1$, the step size is reduced.
+For $\alpha=1$, MBP4 is equivalent to the refined BP4 in [13].
+We will specify the value of $\alpha$ only when $\alpha\neq 1$ is used.
+The optimization of $\alpha$ is discussed in Section VI-A.
+The maximum number of BP iterations, ${\mathrm{T}}$, is set to $100$ unless otherwise specified.
 Furthermore, our MBP4+ADOSD4 scheme demonstrates robustness to variations in the maximum number of iterations, as discussed in Section VI-B.
-For each data point in the plots, at least 100 logical error events are collected.
-In Sections VI-D and VI-E, we first demonstrate the performance of the proposed BP+OSD schemes.
-Then, we analyze the execution time for these schemes in Section VI-G, illustrating the efficiency of ADOSD.
-In the simulations of ADOSD4, $\Gamma$ is chosen according to (11) for the purpose of comparison with OSD-2.
-The soft reliability threshold is set to $\theta=0.999995$.
-Based on our simulations, selecting $\theta$ within the range of $0.999$ to $0.999995$ results in similar performance.
-For comparison, we consider the following decoders: (1) the AMBP4 decoder [31] using a random serial schedule.
-(2) mOSD-($w$, $\lambda$), a partial order-$w$ OSD approach introduced in [42], which selects $\sum_{j=0}^{w}\binom{\lambda}{j}$ candidates for $\lambda$ less reliable bits based on the reliability metric Max.
-(3) the BP-OSD-CS($\lambda$) decoder in [43].
-(4) BP-GDG algorithm in [54].
-These decoders are summarized in Table I.
-We observe that both parallel and serial schedules in BP yield similar results for toric codes and surface codes.
+The soft reliability threshold is set to $\theta=0.999995$ and its selection is discussed in Section VI-C.
+In ADOSD4 simulations, $\Gamma$ is chosen according to (16) to enable comparison with OSD-2.
+In Section VI-D, we compare the proposed BP+OSD schemes with various reliability measures using GHP codes.
+We then evaluate decoding performance on BB codes and topological codes in Sections VI-E, and VI-F.
+Finally, Section VI-G analyzes the execution time of these schemes, demonstrating the efficiency of ADOSD.
+The decoding performance of a quantum code ${\cal C}({\cal S})$ and a decoder is assessed through the logical error rate (LER).
+For a code with distance $d$, errors up to $t=\lfloor\frac{d-1}{2}\rfloor$ are correctable, while some errors of weight $t+1$ become uncorrectable.
+Consequently, at low physical error rates, the performance curve of a decoder follows a scaling behavior of $O(\epsilon^{t+1})$.
+This implies that in a log-domain plot, the decoder’s performance curve should exhibit an error floor with a slope of $t+1$.
+A good decoder will have a low error floor.
+However, if a decoder fails to correct all errors of weight up to $t$, the error floor will have a slope lower than $t+1$.
+In the worst case, a flatter error floor may appear earlier than expected.
+Each data point in the plots is based on at least 100 logical error events in the low error rate regime.
+At LER $10^{-6}$, this represents an error bar less than $2\times 10^{-7}$, which is roughly the size of the markers for the data points.
+Thus the error bars are neglected for clarity.
+We observe that both parallel and serial schedules in MBP4+ADOSD4 yield similar results for toric codes and surface codes.
 However, the parallel schedule outperforms the serial schedule on the [[882, 48, 16]] GHP code.
 We believe that this is because that the parallel schedule typically results in unsettled oscillations in BP, which can be resolved by OSD, whereas the serial schedule may make premature decisions on some bits, leading to incorrect convergence in the decoding process.
-Therefore, we consider the parallel schedule for the following simulations.
+Therefore, we consider the parallel schedule for the following simulations of MBP4+ADOSD4.
 Similar observations are also made in [55, 65].
+decoder source note
+MBP4+ADOSD4 based on Algorithm 4
+MBP4+OSD4-$w$ this paper reliability metric based on Definition 4
+MBP4+mOSD4-$w$ reliability metric based on Definition 5
+AMBP4 [32] No postprocessing. Can act as
+a predecoder for ADOSD4 and OSD4
+BP+mOSD($w$,$\lambda$) [43] binary BP+ order-w OSD4
+with additional flips on $\lambda$ positions
+[…]
+a greedy search on (${\lambda\choose 2}$ positions
+BP+GDG [55] binary BP with guided decimation
+guessing based on a decision tree
+TABLE I: Various decoders compared in this paper.
+For comparison, we consider several decoders listed in Table I. Except for AMBP, we do not implement these decoders ourselves but instead compare our results with those reported in the literature.
+In [66], it was shown that a random order schedule improves BP decoding performance.
+Here, we implement the AMBP decoder with a random order serial schedule, achieving better performance than the results reported in [32].
+Note that AMBP incorporates various techniques for performance enhancement, and we do not focus on optimizing its performance in the following simulations.
 VI-A Optimization of the parameter $\alpha$ in MBP4 for ADOSD.
-In this subsection, we demonstrate how to optimize the parameters of our BP+OSD schemes using LCS codes [59].
-LCS codes are a specialized type of quantum code that combines the features of surface codes and lift-product constructions.
-Decoding LCS codes presents the most challenging problem among all the quantum codes considered in this paper.
-These codes cannot be effectively handled by our previous AMBP4 decoder [31], even though AMBP4 outperforms binary BP-OSD decoders [42, 43] on all other codes.
-In the following, we consider the MBP4 decoder introduced in [31] as a predecoder for OSD.
-MBP4 includes a parameter $\alpha$ that controls the step size of message updates during BP iterations: when $\alpha<1$, the step size is enlarged; when $\alpha>1$, the step size is reduced.
-For $\alpha=1$, MBP4 is equivalent to the refined BP4 in [12].
-We will specify the value of $\alpha$ only when $\alpha\neq 1$ is used.
-Thus, optimizing the choice of $\alpha$ is crucial.
+In this subsection, we demonstrate how to optimize the parameter $\alpha$ of our MBP4+ADOSD4 scheme using LCS codes [60] with ${\mathrm{T}}=100$ and $\theta=0.999995$.
 For most nondegenerate codes, $\alpha=1$ is sufficient to achieve good decoding performance.
 However, for degenerate codes, selecting $\alpha>1$ can result in more stable message passing, leading to improved OSD performance.
-This optimization should be performed for each specific code, although the same value of $\alpha$ typically applies across an entire code family.
-Figure 2 demonstrates the performance of MBP4+ADOSD4444 on the $[[175,7,7]]$ LCS code with various values of $\alpha$.
-The performance of MBP4+ADOSD4444 improves significantly and stabilizes as $\alpha$ increases from $1$ to approximately $1.6$.
-For reference, we provide a curve $y=ax^{4}$ for some $a$ if errors of weight up to half of the code distance $\lfloor\frac{d-1}{2}\rfloor$ can be corrected.
-Therefore, MBP4+ADOSD4444 with $\alpha=1.6$ matches this prediction around LER $10^{-7}$.
-On the other hand, AMBP meets an earlier error floor around LER $10^{-3}$.
-Figure 2 also includes a curve labeled MLE (independent X/Z), which is an estimate of the maximum likelihood binary decoding performance using integer programming, as reported in [59].
-Since the decoding of X and Z errors is performed independently, this MLE curve represents the optimal performance only under the assumption that X and Z errors are uncorrelated.
-In [59], the performance of the BP-OSD-CS decoding algorithm from [43] was shown to approximately match the MLE curve.
-However, it is evident that MBP4+ADOSD4 significantly outperforms this MLE curve.
-We observe a similar performance of MBP4+ADOSD4 for $[[15,3,3]]$, $[[65,5,5]]$ and $[[369,9.9]]$ LCS codes as well, while AMBP has an even higher error floor around LER $10^{-2}$ for the $[[369.9,9]]$ LCS code and a lower error floor around $10^{-5}$ for the $[[65,5,5]]$ LCS code.
+This optimization should be conducted for each code and potentially for each error rate, though a single value of $\alpha$ often applies across an entire code family.
+LCS codes are a specialized type of quantum code that combines the features of surface codes and lift-product constructions.
+Decoding LCS codes presents the most challenging problem for AMBP4 among all the quantum codes considered in this paper.
+We find that the AMBP4 decoder [32] is ineffective for LCS codes, despite its superior performance over binary BP-OSD decoders [43, 44] on all other codes.
 Figure 2: MBP4+ADOSD4 decoding with various values of $\alpha$ for the $[[175,7,7]]$ LCS code.
@@
 The notation $\alpha\in(1.6,-0.01,0.5)$ means that $\alpha$ is tested in the sequence $1.60,1.59,1.58,\dots,0.51,0.50$.
+Figure 2 compares various decoders on the $[[175,7,7]]$ LCS code, including an MLE (independent X/Z) reference curve from [60].
+This MLE curve serves as a benchmark, as a decoder capable of correcting errors of weight up to $3$ should exhibit a performance curve parallel to it at low error rates.
+The performance of MBP4+ADOSD4 on the $[[175,7,7]]$ LCS code, with varying values of $\alpha$, exhibits fluctuations at lower values of $\alpha$ but stabilizes as $\alpha$ increases to 1.6.
+With $\alpha=1.6$, its performance curve indicates that it can correct nearly all weight-3 errors at LER $10^{-7}$.
+In [60], the BP-OSD-CS60 decoder was shown to closely match the MLE curve on this code.
+Since the decoding of $X$ and $Z$ errors is performed independently, the MLE curve represents optimal performance under the assumption that $X$ and $Z$ errors are uncorrelated.
+As MBP4+ADOSD4 leverages the $X/Z$ correlations, it significantly outperforms both the MLE curve and BP-OSD-CS60 by more than 15 times.
+In contrast, AMBP4 encounters an error floor for $\epsilon\leq 0.03$.
+We observe a similar performance of MBP4+ADOSD4 for $[[15,3,3]]$, $[[65,5,5]]$ and $[[369,9.9]]$ LCS codes as well, while AMBP has an even higher error floor around LER $10^{-2}$ for the $[[369.9,9]]$ LCS code and a lower error floor around $10^{-5}$ for the $[[65,5,5]]$ LCS code.
+For improved performance, $\alpha$ can be adaptively chosen for each $\epsilon$, as demonstrated in Section VI-D.
 VI-B Robustness of the maximum iterations in MBP4+ADOSD4
-Figure 3: MBP4+ADOSD4 decoding with various maximum number of iterations $T$ for the $[[175,7,7]]$ LCS code.
+Figure 3: MBP4+ADOSD4 decoding with various maximum number of iterations ${\mathrm{T}}$ for the $[[175,7,7]]$ LCS code.
 Next, we demonstrate that the performance of ADOSD4 remains robust even with variations in the maximum number of iterations in MBP4.
@@
 However, the failures of the HRSR algorithm can impact the simulation time, as OSD-$w$ is significantly more time-consuming.
-Figure 3 illustrates the performance of MBP4+ADOSD4 decoding with varying maximum numbers of iterations, $T=1,2,3,4,5,10,20,50,100,200$.
-The curve for $T=4$ is already good enough and smooth.
+Figure 3 illustrates the performance of MBP4+ADOSD4 decoding for the [[175,7,7]] LCS code with varying maximum numbers of iterations, ${\mathrm{T}}=1,2,3,4,5,100,200$.
+The curve for ${\mathrm{T}}=4$ is already good enough and smooth.
 While increasing the number of iterations results in slightly improved performance, the curves converge closely around an LER of $2\times 10^{-7}$ to $3\times 10^{-7}$.
-In the extreme case of $T=1$, there are few highly reliable bits or valid highly reliable bits, causing the decoder to essentially revert to performing ordinary OSD4 with increased complexity.
-We have demonstrated that the performance of ADOSD4 remains robust even with variations in the maximum number of iterations in MBP4.
-To balance accuracy and efficiency, we typically select $T$ within the range of $50$ to $100$.
-VI-C Bivariate bicycle codes
-A class of QLDPC codes, referred to as BB codes, was proposed in [55], demonstrating a high error threshold and a significantly higher code rate compared to topological codes.
-Several attempts at decoding this code family have been presented in [53, 54] under the code capacity noise model.
-However, the BP+CB decoder in [53] does not perform as well as BP-OSD-0.
-The BP+GDG decoder in [54] shows improved performance, surpassing BP-OSD-0 and being comparable to BP2-OSD-CS10 over independent X errors, as shown in [54, Figure 4].
-Figure 4 presents the simulation results of MBP4+ADOSD4 and AMBP4 decoding for BB codes, both of which demonstrate significantly better performance than GDG decoding.
-For instance, consider the [[144,12,12]] BB code.
-At an $X$ error rate of $0.02$ (corresponding to a depolarizing rate of about $0.03$), BP+GDG achieves an LER of approximately $10^{-4}$.
-In contrast, both MBP4+ADOSD4 and AMBP4 achieve an LER below $4\times 10^{-6}$ at the same depolarizing rate.
-Moreover, MBP4+ADOSD4 outperforms AMBP4.
-For the [[144,12,12]] code, we also include the decoding performance of MBP4+OSD4-2, which is nearly identical to that of MBP4+ADOSD4.
-Figure 4: AMBP4 and MBP4+ADOSD4 decoding performance for various BB codes.
+$\epsilon$ 0.1 0.056 0.032 0.018 0.01 0.007 0.0056
+Avg. iter. 194.46 97.67 23.51 6.31 2.44 1.70 1.23
+TABLE II: The average number of iterations for MBP4 decoding on the $[[175,7,7]]$ LCS code with ${\mathrm{T}}=200$.
+Table II presents the average number of iterations for ${\mathrm{T}}=200$.
+It can be observed that the average number of iterations rapidly decreases to one or two when $\epsilon<0.01$.
+Since we are targeting lower error rates, to balance accuracy and efficiency, we typically choose ${\mathrm{T}}=100$.
+VI-C Selection of Reliability Threshold
+We now discuss the selection of the reliability threshold. Figure 4 presents an example for the $[[175,7,7]]$ LCS code, which is representative of our simulation results.
+Based on our experience, at high error rates, the reliability threshold has little impact on performance.
+However, at low error rates, where higher accuracy is required, the choice of $\theta$ depends on both ${\mathrm{T}}$ and the specific code.
+For instance, MBP4+ADOSD4 achieves excellent performance on the $[[175,7,7]]$ LCS code with ${\mathrm{T}}=2$ and $\theta=0.5$ but performs poorly on the (4.8.8) color code with $d=5$ under the same setting.
+With ${\mathrm{T}}=100$ and $\theta=0.999995$, we generally observe good and stable decoding performance across all codes considered, as shown in Figure 4.
+Therefore, we adopt these values for our analysis in the following subsections.
+Figure 4: MBP4+ADOSD4 decoding with various values of reliability threshold $\theta$ for the $[[175,7,7]]$ LCS code.
 VI-D Generalized hypergraph-product codes
-The $[[882,48,16]]$ GHP code [42] is highly-degenerate, which causes conventional BP to perform poorly on this code.
-With minimum distance of 16, we can gauge the effectiveness of a good decoding algorithm on this GHP code.
-The performance curves of different combinations of MBP and OSD are shown in Figure 5, where $\alpha$ is either fixed at 1.6 or chosen by $\alpha(\epsilon)$.
-It is suggested in [31, Fig.
-3] (arXiv version) that $\alpha$ and $\epsilon$ follow the relationship $\alpha(\epsilon)=-0.16\log_{10}(\epsilon)-0.48$, with the coefficients $-0.16$ and $-0.48$ determined through pre-simulations on the [[882, 48]] code.
-For comparison, we also plot the performance of AMBP4 [31] and the BP+mOSD-($w$, $\lambda$[42].
-Figure 5: Comparison of various BP-OSD schemes on the $[[882,48]]$ GHP code. The curve BP+mOSD-(15,15) is from [42].
-The results are summarized as follows.
-  1. 1.
-    MBP4 can be greatly enhanced by OSD.
-  2. 2.
-    MBP4+OSD4-00 outperforms MBP4+mOSD4-00 (without hard-decision reliability) by roughly half an order of performance.
-This suggests that the hard-decision history of BP is crucial in determining the reliability order.
-  3. 3.
-    MBP4+ADOSD4 performs slightly better than MBP4+OSD4-2, while also reducing the time consumption, as shown in Table IV in Section VI-G.
-  4. 4.
-    Both MBP4+ADOSD4 and MBP+OSD4-2 improve upon BP+mOSD-($15$, $15$[42] by more than an order of magnitude at $\epsilon=0.05$.
-Note that mOSD-($15$, $15$) essentially enumerates all the 15 less reliable bits.
-Figure 6: Comparison of various metrics for reliability order in MBP4-OSD4 on the [[882, 48, 16]] GHP code.
-In this paper, we propose using integrated hard-decision and soft reliabilities as a reliability metric (Definition 4).
+The $[[882,48,16]]$ GHP code [43] is highly degenerate, which causes conventional BP to perform poorly.
+Figure 5 presents the performance curves of different MBP and OSD combinations, where $\alpha$ is either fixed at 1.6 or adaptively chosen as $\alpha(\epsilon)$.
+For comparison, we also include the performance of AMBP4 [32] and BP+mOSD($w$, $\lambda$[43].
+BP+mOSD($w$, $\lambda$) exhibits an early error floor around $\epsilon=0.09$.
+Both MBP4+ADOSD4 and MBP4+OSD4-2 with $\alpha=1.6$ outperform BP+mOSD($15$, $15$[43] by more than an order of magnitude at $\epsilon=0.05$.
+[…]
+Figure 5: Comparison of various decoders on the $[[882,48,16]]$ GHP code. The curve BP+mOSD(15,15) is from [43].
+Figure 6: Comparison of various metrics for reliability order in MBP4-OSD4 on the [[882, 48, 16]] GHP code with $\alpha=1.6$.
+In ADOSD4, we propose using integrated hard-decision and soft reliabilities as a reliability metric (Definition 4).
 Herein we also simulate the performances of MBP4+mOSD based on the reliability metrics of Entropy or Max as shown in Figure 6, including mOSD4  as well.
@@
     The performances of mOSD based on Definition 5, Entropy, or Max are mostly comparable; mOSD-$w$ based on Definition 5 is slightly better than the other two metrics at lower error rate for $w=2$.
-We have conducted additional simulations on a [[625,25,8]] GHP code [43], as shown in Figure 7.
-In these examples, MBP4+ADOSD4 slightly outperforms both AMBP4 and MBP4+OSD4-2, with all three decoders performing significantly better than BP-OSD-CS60 in [43].
-Figure 7: Decoding performance of the [[625,25,8]]] GHP code.
-Note that the performance curve for BP-OSD-CS60 is taken from [43, Figure 5], with its independent X/Z error rate $p$ rescaled by a factor of $1.5$ to approximate depolarizing error performance.
-VI-E 2D topological codes
-Next, we consider the decoding performance of our BP+OSD schemes on various 2D topological codes.
-Since AMBP4 is also capable of decoding various families of topological codes, it serves as a useful reference for evaluating the performance of the BP+OSD approach.
-It is important to examine the decoding performance when the logical error rate is around $10^{-6}$, as this can reveal whether an iterative decoder exhibits an error floor behavior.
-Figure 8 illustrates the performance curves of MBP4+ADOSD4 and AMBP4 on the (4.8.8) color codes with distances ranging from 3 to 17.
-The results show that MBP4+ADOSD4 using $\alpha=1.6$ improves AMBP4 starting from $d=5$.
-The advantages of our BP+OSD schemes over AMBP4 are also evident in rotated toric, rotated surface, XZZX, and (6.6.6) color codes.
-Additionally, MBP4+OSD4-2 and MBP4+ADOSD4exhibit nearly identical performance on these codes.
-Figure 8: AMBP4 vs. MBP4+OSD4-2 on the (4.8.8) color codes.
-To sum up, ADOSD4 and OSD4-2 exhibit nearly identical performances on toric, surface, and twisted XZZX codes, with OSD4-2 performing slightly better on the (6.6.6) and (4.8.8) color codes.
-The threshold analysis for these codes in the next subsection will reflect these results as well.
-We remark, however, the time consumption for ADOSD4 is significantly lower than that for OSD4-2, as will be shown in Section VI-G.
-VI-F Threshold analysis for the topological codes
-The decoding performance of a family of 2D topological codes is typically evaluated by the its simulated error threshold, which can be estimated using the scaling ansatz method [66, 67], which is explained as follows.
-Suppose that the threshold value of a code family is $\epsilon_{\mathrm{th}}$. For a code in this family of distance $d$, if the logical error rate is $p_{L}$ at physical error rate $\epsilon$, then
-$p_{L}=f(d^{\frac{1}{\nu}}(\epsilon-\epsilon_{\mathrm{th}})),$ (14)
-where $f$ is a low-degree polynomial and $\nu$ is the critical exponent. The scaling ansatz method involves fitting this equation with the best values of $\epsilon_{\mathrm{th}}$, $\nu$, and $f$.
-Figure 9: The threshold of BP4+OSD4-2 on the surface codes is about $17.67\%$ from the simulations. The dashed line stands for no error correction.
-Figure 10: The accuracy of the threshold estimation about MBP4+OSD4-2 on the rotated surface codes can be demonstrated using the critical scaling approach.
-The estimation process begins with $d=7$ and continues up to $d=17$, ensuring that no biased information influences the results.
-The threshold analysis for rotated surface codes is presented in Figure 9 and Figure 10.
+VI-E Bivariate bicycle codes
+BB codes are a class of QLDPC codes proposed in [56], featuring a high error threshold and a significantly higher code rate than topological codes.
+Several attempts at decoding this code family have been presented in [54, 55] under the code capacity noise model.
+However, the BP+CB decoder in [54] does not perform as well as BP-OSD-0.
+The BP+GDG decoder in [55] shows improved performance, surpassing BP-OSD-0 and being comparable to BP2-OSD-CS10 over independent X errors, as shown in [55, Figure 4].
+Figure 7 presents the simulation results of MBP4+ADOSD4 and AMBP4 decoding for BB codes, both of which demonstrate significantly better performance than GDG decoding.
+For instance, consider the [[144,12,12]] BB code.
+At an $X$ error rate of $0.02$ (corresponding to a depolarizing rate of about $0.03$), BP+GDG achieves an LER of approximately $10^{-4}$.
+In contrast, both MBP4+ADOSD4 and AMBP4 achieve an LER below $4\times 10^{-6}$ at the same depolarizing rate.
+Moreover, MBP4+ADOSD4 outperforms AMBP4.
+For reference, we include dotted $O(\epsilon^{t+1})$ curves for each code.
+The $[[72,12,6]]$ and $[[90,8,10]]$ codes closely follow these reference curves, while the other three codes exhibit lower error floors below LER $10^{-7}$.
+Figure 7: AMBP4 and MBP4+ADOSD4 decoding performance for various BB codes.
+VI-F 2D topological codes
+Figure 8: MBP4+OSD4-2, MBP4+ADOSD4, and AMBP4 decoding performance for various $[[d^{2},1,d]]$ rotated surface codes.
+Figure 9: (a) The threshold of MBP4+OSD4-2 on the $[[d^{2},1,d]]$ rotated surface codes with $\alpha=1.0$ is about $17.67\%$ from the simulations.
 For each data point, we collected at least 10,000 logical error events.
-The analysis reveals a threshold of approximately 17.67% for BP4+OSD4-2 on surface codes.
-Similar analyses are conducted for toric, twisted XZZX, (6.6.6), and (4.8.8) color codes, using both BP4+OSD4-2 and BP4+ADOSD4.
-In the simulations, the data is fitted using a low-degree polynomial.
-Note that MBP4 with $\alpha=1$ is used in the threshold analyses. Optimizing the value of $\alpha(\epsilon)$ for MBP can enhance overall performance.
-The summarized results, including thresholds obtained with AMBP4, are shown in Table II.
-These results are consistent with the simulations in the previous subsection.
-Notably, BP4+OSD4-2 shows a significant enhancement over AMBP4 for surface and color codes.
-TABLE II: Thresholds for 2D topological codes subjecte to depolarizing errors. The parameter for MBP4 is set to $\alpha=1.0$.
+The dashed line stands for no error correction.
+(b) The accuracy of the threshold estimation about MBP4+OSD4-2 on the rotated surface codes can be demonstrated using the critical scaling approach.
+The estimation is based on data from $d=7$ to $d=17$, ensuring that the results are not influenced by biased information.
+Here, we evaluate the decoding performance of our BP+OSD schemes on various 2D topological codes.
+Examining LER around $10^{-6}$ is important, as it helps identify whether an iterative decoder exhibits error floor behavior.
+Figure 8 presents the performance curves of MBP4+ADOSD4 and MBP4+OSD4-2 on $[[d^{2},1,d]]$ rotated surface codes with distances ranging from 5 to 17.
+They exhibit excellent performance, closely following the dotted reference curves.
+The threshold analysis for $[[d^{2},1,d]]$ rotated surface codes, conducted using the scaling ansatz method [67, 68], is shown in Figure 9.
+The results show a threshold of approximately 17.67% for BP4+OSD4-2, with MBP4+ADOSD4 exhibiting nearly identical performance.
+Notably, they outperform AMBP4, particularly for larger $d$, resulting in a higher threshold compared to the  16% threshold of AMBP4.
+In summary, our decoding schemes demonstrate superior performance on surface codes compared to the MWPM decoder [69, 70], both in terms of LER performance and error threshold (see, e.g., [71]).
+Similar analyses are performed for toric, twisted XZZX, (6.6.6), and (4.8.8) color codes using both BP4+OSD4-2 and BP4+ADOSD4.
+The summarized threshold results are presented in Table III.
+ADOSD4 and OSD4-2 exhibit nearly identical performances on toric, surface, and twisted XZZX codes, with OSD4-2 performing slightly better on the (6.6.6) and (4.8.8) color codes.
+Figure 10: AMBP4 and MBP4+ADOSD4 decoding performance for various (4.8.8) color codes.
+TABLE III: Thresholds for 2D topological codes subjecte to depolarizing errors. The parameter for MBP4 is set to $\alpha=1.0$.
 […]
-We also compare the thresholds of our decoder with those of other known decoders in Table III, including minimum-weight perfect matching (MWPM) [68, 69].
-In [43], BP2+mOSD-(2,60) achieved a threshold of $p=9.9\%$ on toric codes for bit-flip errors.
-We estimate that it has a threshold of $\epsilon_{\mathrm{th}}=9.9\%\times\frac{3}{2}=14.85\%$ for depolarizing errors [1].
-TABLE III: Thresholds of several decoders over depolarizing errors.
-[…]
-: estimated from the threshold of 9.9% in the bit-flip channel [43].
+The thresholds for the color codes are lower than those of toric and surface codes.
+This is evident from the performance curves of the (4.8.8) color codes as shown in Figure 10.
+While the slopes of these curves increase as the error rate decreases, they do not closely match the slopes of the dotted reference curves.
+This suggests that some low-weight errors, even those with weight less than $t+1$, may remain uncorrected by these decoders.
+Consequently, additional techniques or higher-order OSD may be necessary.
+TABLE IV: Comparison of the time consumption for surface codes at LER about $10^{-6}$.
+code $\epsilon$ MBP4 time per MBP4 time per time per average iterations ADOSD4 time
+distance failure rate iteration ($\mu$s) ADOSD4 ($\mu$s) OSD4-2 ($\mu$s) when MBP4 succeeds in MBP4 iterations
+11 1.7% 21.46% 0.01464 0.0272 0.667 1.002 2
+13 2.5% 35.97% 0.0212 0.0469 1.342 1.79 2.21
+15 3.3% 51.47% 0.0287 0.0506 2.155 2.41 1.76
+${}^{\dagger}:$ the time per ADOSD4 operation divided by the time per MBP4 iteration.
 VI-G Time consumption
-TABLE IV: Comparison of the time consumption between ADOSD4 and OSD4-2 at logical error rate about $2\times 10^{-6}$.
-Code BP failure ADOSD4 OSD4-2 time LER
-rate time ($\mu$s) time ($\mu$s) ratio ratio
-toric 27.76% 88.63 4712.9 1.88% $\approx 0$dB
-surface 67.69% 71.97 3549 2.82% $\approx 0$dB
-(6.6.6) 10.32% 35.42 3812 0.93% -2.3dB
-(4.8.8) 7.17% 26.50 1265 2.09% -3.9dB
-XZZX 11.93% 22.29 1001.3 2.23% $\approx$ 0dB
-GHP 0.86% 1188.7 45910.3 2.59% 2.49dB
-: the column ‘time ratio’ represents the ratio of MBP4+ADOSD4 time to OSD4-2 time.
-: the column ‘LER ratio’ represents the ratio of the logical error rate using ADOSD4 to that using OSD4-2. If the LER ratio is greater than 0dB, it means that ADOSD4 is better.
-In the previous subsections, we have demonstrated the effectiveness of our BP+OSD schemes. In this subsection we analyze the time efficiency of our OSD schemes.
-VI-G1 Time Consumption of ADOSD4 vs. OSD4-2
-First, we compare the time consumption of OSD4-2 and ADOSD4 for various topological codes of comparable distances with the [[882, 48,16]] GHP code, including the [[324,2,18]] rotated toric code, the [[289,1,17]] rotated surface code, the [[[217, 1,17]] (6.6.6) color code, the [[161, 1,17]] (4.8.8) color code, and the [[145, 1,17]] twisted XZZX code.
-For each code, the simulated physical error rate is chosen so that the logical error rate of MBP4+ADOSD4 is approximately $2\times 10^{-6}$.
-We fix $\alpha=1$ in this subsection and the decoding performance can be further improved.
-In the simulations, when BP fails, the BP outputs will be handled by ADOSD4 and OSD4-2, respectively, We note that ADOSD4 is invoked only when MBP4 fails.
-and their time consumption statistics will be counted separately in the following tables.
-For example, ADOSD4 time only counts the time to run Algorithm 4 when MBP4 fails.
-The simulation results are summarized in Table IV and the unit of time is microseconds ($\mu$s) measured using a single-thread desktop computer.
-It can be observed that the time consumption of ADOSD4 is roughly $1\%$ to $3\%$ of the time consumption for OSD4-2 at logical error rate about $2\times 10^{-6}$ for the various quantum codes.
-For the toric, surface, and twisted XZZX codes, ADOSD4 and OSD4-2 exhibit almost identical performances, while OSD4-2 performs better on the color codes and ADOSD4 performs better on the GHP code.
-This is reflected in the last column of Table IV.
-To summarize, ADOSD4 is at least 30 times faster than OSD4-2, with a maximum performance loss of 4dB for topological codes at logical error rate $2\times 10^{-6}$.
-Additionally, for the [[882, 48]] GHP code, ADOSD4 also enhances performance.
-VI-G2 Time Consumption of ADOSD4 vs. BP
-We have demonstrated that ADOSD4 is generally faster than OSD4-2. Next, we compare the time consumption of ADOSD4 to MBP4, using surface codes.
-TABLE V: Comparison of the time consumption for surface codes at logical error rate about $10^{-6}$.
-code average time per MBP4 time per time per
-distance iterations MBP4 failure ADOSD4 OSD4-2
-iteration ($\mu$s) rate ($\mu$s) ($\mu$s)
-11 22.3473 0.01464 21.46% 0.0272 0.667
-13 37.1148 0.0212 35.97% 0.0469 1.342
-15 52.6400 0.0287 51.47% 0.0506 2.155
-: average over all samples, including MBP failures.
-TABLE VI: Time Consumption of ADOSD4 vs. BP for the surface codes
-code $\epsilon$ average iterations ADOSD4 time
-distance when MBP4 succeeds in MBP4 iterations
-11 1.7% 1.002 2
-13 2.5% 1.79 2.21
-15 3.3% 2.41 1.76
-: average over only the samples when MBP4 succeeds.
-Table V lists the time consumptions of MBP4, ADOSD4, and OSD4-2 for surface codes with distances of 11, 13, and 15.
-To achieve a logical error rate $10^{-6}$ using ADOSD4, the physical error rates are set to 0.017, 0.025, and 0.033 for codes with distances 11, 13, and 15, respectively.
+In the previous subsections, we have demonstrated the effectiveness of our BP+OSD schemes. In this subsection we analyze the time efficiency of our OSD schemes. We fix $\alpha=1$ in MBP4.
+VI-G1 Time Consumption of ADOSD4 vs. OSD4-2 and BP
+Table IV lists the average time consumptions of MBP4, ADOSD4, and OSD4-2 for surface codes with distances of 11, 13, and 15 at LER arouond $10^{-6}$.
+The unit of time is microseconds ($\mu$s) measured using a single-thread desktop computer.
+To achieve LER $10^{-6}$ using MBP4+ADOSD4, the physical error rates are set to 0.017, 0.025, and 0.033 for codes with distances 11, 13, and 15, respectively.
 At least $10^{5}$ samples are collected for each case.
-For example, at $d=11$, the MBP4 failure rate is $21.46\%$ at physical error rate 1.7%.
-Since the maximum number of iterations is set to be $T=100$, this contributes to an average of $21.46$ iterations.
-Since the average iterations include MBP4 failures is 22.3473, the average iterations when MBP4
-Table VI summarizes the time consumption of ADOSD4 in terms of MBP4 iterations.
+It can be observed that the time consumption of ADOSD4 is approximately 2% to 4% of that of OSD4-2, with similar results observed for other quantum codes in this study.
+In summary, ADOSD4 is at least 25 times faster than OSD4-2 at an LER of around $10^{-6}$.
+Table IV also lists the time consumption of ADOSD4 in terms of MBP4 iterations.
 Specifically, the average time consumption per codeword for ADOSD4 is approximately equivalent to 1.76 to 2.21 MBP4 iterations.
 This shows that the time consumption for ADOSD4 is roughly equivalent to 1 to 2 successful MBP4 executions.
-Furthermore, the time usage for ADOSD4 is significantly smaller than OSD4-2.
-VI-G3 HRSR and OSD-0 in ADOSD4
+VI-G2 HRSR and OSD-0 in ADOSD4
 We have shown that ADOSD4 is efficient, thanks to Algorithm 3 and Corollary 10.
@@
 Now, we examine the effectiveness of these techniques on the surface code of distance 11, with similar results observed for distances 13 and 15.
-Table VII enumerates the percentage of ADOSD4 trials that only perform OSD4-0 using Corollary 10.
+Table V enumerates the percentage of ADOSD4 trials that only perform OSD4-0 using Corollary 10.
 At $\epsilon=4.1\%$, 93.61% of ADOSD4 trials already perform only OSD4-0, increasing to 99.84% at $\epsilon=1.7\%$.
 Thus, at lower error rates, almost all trials execute OSD4-0.
-Moreover, the HRSR algorithm effectively reduces the dimensions of the check matrix in both rows and columns.
-Table VII also provides the percentages of trials where the number of rows and columns in the check matrix is reduced to 20% or 30% of the original (denoted as $20\%$ and $30\%$ effective dimensions in the table).
+Moreover, the HRSR algorithm reduces the dimensions of the check matrix in both rows and columns.
+Table V also provides the percentages of trials where the number of rows and columns in the check matrix is reduced to 20% or 30% of the original (denoted as $20\%$ and $30\%$ effective dimensions in the table).
 These percentages increase quickly as the error rate decreases.
-With fewer than 30% effective dimensions, the time consumption for Gaussian elimination reduces to roughly $0.3^{3}=2.7\%$, consistent with the results in Table IV.
-TABLE VII: Percentage of OSD-0 in ADOSD4 on the surface code of distance $11$
+With fewer than 30% effective dimensions, the time consumption for Gaussian elimination reduces to roughly $0.3^{3}=2.7\%$, consistent with the previous results.
+TABLE V: Percentage of OSD-0 in ADOSD4 on the surface code of distance $11$
 depolarizing rate $\epsilon$ 1.7% 2.7% 4.1%
@@
 Furthermore, these techniques can be integrated with other methods available in the literature.
-Although the intention of [56] and [57] is quite similar to our HRSR algorithm, the underlying approaches differ: their methods focus on dividing the decoding problem into several smaller problems, whereas our approach involves simply removing the most reliable bits.
-Consequently, it may be possible to combine their methods with ours.
-For example, apply HRSR first, followed by either of the algorithms.
 In our investigations, we primarily focused on MBP4+OSD4-2 and ADOSD4.
@@
 We have observed some instances of surface codes with order $w$ up to 13 in (16).
-An interesting observation is that we can identify bit-flips in error candidates corresponding to multiplying a stabilizer operator, as stated in Lemma 8 and Corollary 10.
+An interesting observation is that we can identify bit-flips in error candidates corresponding to multiplying a stabilizer operator, as stated in Theorem 8 and Corollary 10.
 This information can be used to improve the algorithm by flipping effective bits in approximate degenerate decoding.
@@
 Our hard-decision reliability can be viewed as an extension of this idea, focusing on the length of the last run in the hard-decision vectors.
-The reliability order in Definition 4 can be adapted to employ alternative metrics, such as Entropy or Max.
+The reliability order in Definition 4 can be adapted to employ alternative metrics, such as Entropy or Max [43].
 From our experiments, the integrated hard- and soft-decision reliabilities yielded the best results.
-Finally, our BP-OSD scheme can be extended to handle both data and syndrome errors in the phenomenological noise model [74, 75], using techniques proposed in [76, 77, 78], or to the circuit-level noise model [79], as demonstrated in our recent work [78].
-Currently, there is significant progress in circuit-level error decoding [80, 81, 55, 82].
-Various decoders have been shown to work for circuit-level noise models, including BP-OSD-CS [43], guided decimal guessing (GDG)[54], ambiguity clustering[56], localized statistics decoding (LSD)[57], and ordered Tanner forest (OTF) [83].
-These decoders are based on detector error models generated by STIM [84] and utilize binary BP combined with postprocessing.
-However, there are several code capacity error decoding problems where BP-OSD is suboptimal, indicating potential for further improvement in the circuit-level noise model.
-Therefore, we aim to extend our BP+OSD method to address these more demanding scenarios effectively.
+Finally, there is significant progress in circuit-level error decoding [76, 77, 78, 56, 79].
+Various decoders have been shown to work for circuit-level noise models, including BP-OSD-CS, GDG, ambiguity clustering[57], LSD [58], and ordered Tanner forest [80].
+Our BP-OSD scheme can be extended to handle both data and syndrome errors in the phenomenological noise model [81, 82], using techniques proposed in [83, 84, 85], as demonstrated in our recent work [86].

### v3

@@
 Abstract
-Efficient decoding of quantum codes is crucial for achieving high-performance quantum error correction.
-In this paper, we introduce the concept of approximate degenerate decoding and integrate it with ordered statistics decoding (OSD).
-Previously, we proposed a reliability metric that leverages both hard and soft decisions from the output of belief propagation (BP), which is particularly useful for identifying highly reliable subsets of variables.
-Using the approach of reliable subset reduction, we reduce the effective problem size.
+Efficient and scalable decoding of quantum codes is essential for high-performance quantum error correction.
+In this work, we introduce Reliable Subset Reduction (RSR), a reliability-driven preprocessing framework that leverages belief propagation (BP) statistics to identify and remove highly reliable qubits, substantially reducing the effective problem size.
 Additionally, we identify a degeneracy condition that allows high-order OSD to be simplified to order-0 OSD.
-By integrating these techniques, we present an ADOSD algorithm that significantly improves OSD efficiency in the code capacity noise model.
-We demonstrate the effectiveness of our BP+ADOSD approach through extensive simulations on a varity of quantum codes, including generalized hypergraph-product codes, topological codes, lift-connected surface codes, and bivariate bicycle codes.
-The results indicate that the BP+ADOSD decoder outperforms existing methods, achieving higher error thresholds and enhanced performance at low error rates.
-Additionally, we validate the efficiency of our approach in terms of computational time, demonstrating that ADOSD requires, on average, the same amount of time as two to three BP iterations on surface codes at a depolarizing error rate of around $1\%$.
-All the proposed algorithms are compared using single-threaded CPU implementations.
+By integrating these techniques, we present an ADOSD algorithm that significantly improves OSD efficiency.
+Our BP+RSR+ADOSD framework extends naturally to circuit-level noise and can handle large-scale codes with more than $10^{4}$ error variables.
+Through extensive simulations, we demonstrate improved performance over MWPM and Localized Statistics Decoding for a variety of CSS and non-CSS codes under the code-capacity noise model, and for rotated surface codes under realistic circuit-level noise.
+At low physical error rates, RSR reduces the effective problem size to less than 5%, enabling higher-order OSD with accelerated runtime.
+These results highlight the practical efficiency and broad applicability of the BP+ADOSD framework for both theoretical and realistic quantum error correction scenarios.
 I Introduction
@@
 Consequently, the complexity of BP4 can be reduced by restricting message passing in the Tanner graph [13] to scalar messages, which encode the likelihood ratios associated with Pauli commutation relations.
-The decoding problem for quantum codes with high degeneracy is particularly challenging for BP, as seen in topological codes [23, 24, 25, 26, 27, 28, 29].
-To address these challenges, several remedies for the BP algorithm have been introduced, including techniques such as random perturbation [9], enhanced feedback [10, 11], check-matrix augmentation [12], message normalization and offset [13, 14], and trapping set analysis [30, 31].
-These methods aim to improve the decoding accuracy and convergence of the BP algorithm.
-A recently developed modified BP4 algorithm, known as MBP4, incorporates additional memory effects and has demonstrated the ability to handle topological codes [32].
-We investigate ordered statistic decoding (OSD) as a complementary approach to BP in quantum coding theory.
-Originally proposed by Fossorier and Lin [33] in classical coding theory for approximating maximum-likelihood decoding [34, 35, 36, 37, 38, 39, 40, 41, 42], OSD sorts reliability measures of symbols to find probable solutions while maintaining controllable complexity [33, 43, 44, 45, 46].
-Syndrome-based OSD has been used when BP fails to identify a valid error [47] and is particularly effective for degenerate quantum codes [43].
-The core operation of OSD is Gaussian elimination on an $m\times n$ parity-check matrix, identifying unreliable bits with a complexity of $O(nm^{2})$.
-Order-$w$ OSD flips up to $w$ reliable bits to generate error candidates, but this incurs high computational costs.
-To mitigate complexity, Fossorier et al.[47] proposed candidate pruning, and various classical techniques further optimized high-order OSD by reducing the number of candidates examined[48, 49, 50, 51, 52, 53].
-[…]
-However, there is a great deal of room for improvement over the current postprocessing techniques in the literature under the code capacity noise model.
-These methods primarily rely on binary BP to handle quaternary errors, neglecting error correlations, which suggests that binary decoders are inherently suboptimal.
-In contrast, the quaternary AMBP4 decoder (the adaptive version of MBP4) from [32] has been shown to outperform all the aforementioned decoders across various quantum codes, including topological codes [25], BB codes [56], and generalized hypergraph-product (GHP) codes [43, 44], without requiring postprocessing.
-For supporting examples, please refer to Figure 7 in Section VI.
-OSD serves as a complementary strategy to BP, addressing cases where BP fails.
-As a result, BP-OSD is naturally expected to outperform AMBP4.
-In this paper, we propose a BP4-based OSD to achieve this improvement and introduce an efficient OSD algorithm for quantum codes.
-First, the sorting process plays a pivotal role in OSD [33].
-For efficient implementation, we consider binary representations of an $n$-qubit quantum code and thus an $n$-qubit Pauli error variable can be represented by $2n$ binary variables.
-Previously we proposed a reliability metric that takes into account the hard-decision history obtained during BP4 iterations and incorporates soft messages from the last iteration [1].
-The $2n$ binary error variables of an error vector are then sorted according to the reliability metric.
-Specifically, the length of the last run in a bit’s hard-decision history is the primary term in our metric.
-Through simulations, we demonstrate that incorporating the hard-decision history to determine the reliability order for OSD enhances decoding performance.
-Our OSD algorithm based on this reliability metric will be denoted as OSD4.
-To develop an efficient OSD algorithm, it is crucial to reduce the problem size in OSD and minimize the computational burden of Gaussian elimination.
-To address this challenge, we propose an approach called reliable subset reduction.
-In the code capacity error decoding problem of an $[[n,k]]$ quantum code, we need to solve a linear system of equations with $2n$ binary variables and $n-k$ independent constraints.
-The core idea is that if the solution for a subset of the variables can be determined with confidence, the linear system can be reduced, thereby lowering the computation burden of both Gaussian elimination and the OSD process.
-So the primary task is to identify a highly reliable subset.
-An excellent choice are the bits that exhibit nearly constant hard decision history vectors during BP4 iterations and show high soft reliability in the final iteration.
-At low error rates, there is a large portion of highly reliable bits.
-Eliminating these bits would significantly accelerate the OSD process.
-Moreover, resources can be allocated to the less reliable parts, similar to the approaches in [43, 44].
-We will demonstrate that in our simulations of surface codes with distances from 11 to 15, the problem sizes are reduced to less than 30% of the original in over 90% of the trials at an error rate of around 2%.
-We note that the ambiguity clustering decoder [57] and LSD [58] operate at a different stage from our approach, and these algorithms could be applied after reliable subset reduction.
-Next, we consider approximate degenerate OSD in quantum codes.
-A typical OSD generates a list of error candidates for a given error syndrome and selects the one with minimum weight, which is optimal for classical codes.
-For quantum codes, however, the optimal decoding criterion involves choosing the error coset with the highest probability, a criterion that is challenging to implement in practice.
-To address this, we consider approximate degenerate decoding in OSD by focusing only on relatively low-weight terms within a coset.
-Additionally, we introduce a degenerate error candidate pruning condition for stabilizer codes.
-We show that certain bit-flips in OSD correspond to multiplying stabilizers, resulting in degenerate error candidates.
-These bit-flips may alter the error weight but do not change the error coset, allowing them to be ignored in approximate degenerate OSD.
-To improve OSD efficiency, the minimum weight decoding criterion is applied.
-A key observation is that if all bit-flips correspond to multiplying a stabilizer, high-order OSD is unnecessary.
-We note that a degeneracy condition is also utilized in ambiguity clustering to facilitate subproblem resolution [57].
-Integrating all the mentioned techniques, we propose an approximate degenerate OSD algorithm, abbreviated as ADOSD4.
-We conduct simulations of MBP4 and our OSD schemes OSD4 and ADOSD4 on several quantum codes under depolarizing errors in the code capacity noise model, including BB codes [56], rotated toric codes and rotated surface codes [25], (6.6.6) and (4.8.8) color codes [28], twisted XZZX codes on a torus [59], GHP codes [43, 44], and lift-connected surface (LCS) codes [60].
+The decoding problem for highly degenerate quantum codes, such as topological codes [23, 24, 25, 26, 27, 28, 29], remains particularly challenging for belief propagation (BP).
+A variety of remedies have been proposed to improve BP convergence and accuracy, including random perturbation [9], enhanced feedback [10, 11], check-matrix augmentation [12], message normalization and offsets [13, 14], and trapping-set analysis [30, 31].
+A notable advance is MBP4 [32], which introduces memory effects and enables BP to decode many topological codes.
+When BP alone is insufficient, postprocessing decoders such as ordered statistic decoding (OSD) are commonly employed.
+OSD was originally developed in classical coding theory as a controllable-complexity approximation to maximum-likelihood decoding [33, 34, 35, 36, 37, 38, 39, 40, 41, 42], and later adapted to syndrome-based and quantum decoding [43, 44].
+Its core operation is Gaussian elimination on an $m\times n$ parity-check matrix, followed by flipping unreliable bits, but high-order OSD incurs significant computational cost [45, 46, 47, 48, 49, 50].
+In the quantum regime, several BP–OSD hybrids and variants have been proposed [51, 52, 53], along with divide-and-conquer approaches such as ambiguity clustering (AC) [54] and localized statistics decoding (LSD) [55].
+These methods attempt to reduce the Gaussian elimination burden by partitioning the problem.
+Despite these developments, most existing postprocessing schemes under the code-capacity noise model rely on binary BP applied to quaternary quantum errors, thereby discarding $X/Z$ correlations and leading to intrinsic suboptimality.
+In contrast, the quaternary AMBP4 decoder [32] directly exploits these correlations and has been shown to outperform all of the above decoders on a wide range of quantum codes, including topological codes [25], BB codes [56], and generalized hypergraph-product codes [44, 51].
+However, AMBP4 still fails on certain important code families, such as lift-connected surface (LCS) codes [57], motivating the need for a more powerful postprocessing framework.
+In this paper, we investigate postprocessing techniques for enhancing the performance of BP.
+When BP fails to converge to a valid global solution, it still provides rich statistical information about the error configuration.
+In particular, motivated by density evolution arguments for sparse LDPC codes [58, 59], most error variables are typically resolved with high confidence, except for a small number of unsettled variables associated with trapping sets [30].
+Consequently, a large fraction of variables can be regarded as already determined and effectively part of the solution.
+Based on this observation, we propose reliable subset reduction (RSR).
+In the code-capacity decoding problem of an $[[n,k]]$ quantum code, decoding amounts to solving a linear system with $2n$ binary variables subject to nkn-k independent constraints.
+The central idea of RSR is that if a subset of variables can be identified as highly reliable, the linear system can be reduced by fixing those variables, thereby substantially lowering the computational burden of subsequent postprocessing steps, such as Gaussian elimination.
+To identify such a reliable subset, we exploit both the hard-decision history and the final soft information produced by BP4 [1].
+Variables whose hard decisions remain nearly constant across BP iterations and whose final beliefs have high confidence are classified as reliable.
+At low physical error rates, a large fraction of variables satisfy these criteria, so eliminating them dramatically reduces the effective problem size and accelerates downstream postprocessing decoders.
+We apply OSD after RSR, using the reliabilities produced by RSR to order the remaining unreliable variables. The resulting OSD variant, which is driven by BP4 statistics, is denoted as OSD4.
+We next consider approximate degenerate OSD for quantum codes.
+In typical OSD, a list of error candidates consistent with a given syndrome is generated and the minimum-weight candidate is selected.
+For quantum codes, however, the optimal criterion is to choose the most probable error coset, which is generally intractable.
+To address this, we adopt an approximate degenerate decoding strategy in which only low-weight representatives within each coset are considered.
+When RSR is applied, the resulting reduced problem makes such approximate degenerate decoding feasible, in a manner analogous to the ambiguity clustering (AC) approach [54].
+For computational efficiency, we nevertheless use minimum-weight selection among the retained candidates in our simulations.
+Additionally, we introduce a degeneracy-aware pruning condition for stabilizer codes.
+We show that certain OSD bit flips correspond to multiplying stabilizers and therefore generate degenerate error candidates: although their weights may differ, they belong to the same error coset.
+Such flips can be ignored in approximate degenerate OSD.
+A key observation is that when all candidate flips correspond to the trivial logical coset (i.e., zero logical syndrome), high-order OSD is unnecessary.
+By combining these ideas, we obtain an approximate degenerate OSD algorithm, which we denote as ADOSD4.
+Finally, there has been substantial recent progress in circuit-level decoding [60, 61, 62, 56, 63].
+A variety of decoders have been developed for circuit-level noise models, including BP-OSD-CS, GDG, AC [54], LSD [55], and ordered Tanner forest [64].
+Our BP–OSD framework can also be extended to handle both data and syndrome errors under the phenomenological noise model [65, 66], using techniques proposed in [67, 68, 69], as demonstrated in our recent work [70].
+In this paper, we further show that our approach can be directly applied to the full circuit-level noise model through the detector error model (DEM) generated by STIM [71].
+The resulting binary DEM decoding problem can be handled using the proposed BP–RSR–ADOSD procedure.
+We conduct simulations of MBP4 and our OSD schemes OSD4 and ADOSD4 on several quantum codes under depolarizing errors in the code capacity noise model, including BB codes [56], rotated toric codes and rotated surface codes [25], (6.6.6) and (4.8.8) color codes [28], twisted XZZX codes on a torus [72], GHP codes [44, 51], and LCS codes [57].
 The results demonstrate that our proposed schemes outperform previous approaches in the literature, both in terms of error thresholds for topological codes and low error rate performance (logical error rate around $10^{-6}$).
-Our MBP4+OSD4 and MBP4+ADOSD4 decoders outperform other binary BP-OSD schemes in these quantum codes.
-We achieve thresholds of approximately 17.5%–17.7% for the toric, surface, and twisted XZZX codes, and thresholds of 15%–15.42% for the (4.8.8) and (6.6.6) color codes, as summarized in Table III.
-They even surpass the minimum weight perfect matching (MWPM) decoder on surface and toric codes [69, 70].
-Finally, we analyze the time efficiency of our OSD schemes for these quantum codes using single-threaded CPU implementations.
-The simulation results, summarized in Table IV, show that the time consumption of ADOSD4 for surface codes with distances from 11 to 15 is approximately 2% to 4% of that for order-2 OSD4 with similar decoding performance at logical error rate around $10^{-6}$.
-Moreover, the average time consumption per codeword for ADOSD4 is approximately equivalent to 2 to 3 BP4 iterations, whereas a successful BP trial typically takes 2 to 3 iterations on average.
-This suggests that ADOSD4 is an excellent complement to BP, especially since ADOSD4 is used only when BP fails.
+In DEM circuit-level setting for rotated surface codes, we further show that MBP4+ADOSD4 consistently outperforms LSD.
+When both $X$- and $Z$-type detector variables are used, MBP4+ADOSD4 also surpasses the minimum-weight perfect-matching (MWPM) decoder [86, 87] and achieves an improved threshold of approximate $0.76\%$.
+Beyond improved decoding performance, a central outcome of our simulations is the strong reduction effect produced by reliable subset reduction (RSR): at low physical error rates, RSR typically reduces the effective decoding problem size by one to two orders of magnitude (e.g., below $5\%$ of the original size for $\epsilon\leq 0.005$, and below $1\%$ for $\epsilon=0.001$ in surface-code DEM problems).
+This dramatic dimensionality reduction enables higher-order OSD to be applied efficiently even for problems with more than $10^{4}$ error variables, a regime that is far beyond the reach of conventional OSD-based postprocessing.
+As a result, our MBP–RSR–ADOSD approach achieves strong threshold performance for topological codes and substantially improved logical error rates in the low-error regime.
 This paper is organized as follows.
 We introduce stabilizer codes in Section II.
-Section III discusses reliability metrics and the associated OSD algorithm.
-The reliable subset reduction method is presented in Section IV.
-We then cover approximate degenerate decoding in Section V, and provide simulation results in Section VI.
-Finally, we conclude in Sec.
-VII.
+The reliable subset reduction method and the associated reliability metrics are presented in Section III.
+We then introduce the associated OSD algorithm in Section IV.
+Approximate degenerate decoding is covered in Section V.
+The MBP-ADOSD procedure is extended to circuit-level decoding via DEM in Section VI, and simulation results are provided in Section VII.
+Related work is discussed in Section VIII.
+Finally, we conclude in Section IX.
 II Quantum stabilizer codes
@@
 $\mathcal{G}_{n}\triangleq\left\{\omega M_{1}\otimes\dots\otimes M_{n}:\omega\in\{\pm 1,\pm i\},M_{j}\in\{I,X,Y,Z\}\right\}.$
-Every nonidentity Pauli operator in $\mathcal{G}_{n}$ has eigenvalues ±$\pm 1$.
+Every nonidentity Pauli operator in $\mathcal{G}_{n}$ has eigenvalues $\pm 1$.
 Any two Pauli operators either commute or anticommute with each other.
@@
 Therefore, at a small error rate $\epsilon$, low-weight Pauli errors are more likely to occur across $n$ qubits.
-A stabilizer group $\mathcal{S}$ is an Abelian subgroup in $\mathcal{G}_{n}$ such that $-I^{\otimes n}\not\in\mathcal{S}$ [7].
-Suppose that $\mathcal{S}$ is generated by $n-k$ independent generators.
+A stabilizer group $\mathcal{S}$ is an Abelian subgroup in $\mathcal{G}_{n}$ such that $I^{\otimes n}\not\in\mathcal{S}$ [7].
+Suppose that $\mathcal{S}$ is generated by nkn-k independent generators.
 Then $\mathcal{S}$ defines an $[[n,k,d]]$ stabilizer code $\mathcal{C}(\mathcal{S})$ that encodes $k$ logical qubits into $n$ physical qubits:
-$\mathcal{C}(\mathcal{S})=\left\{\ket{\psi}\in\mathbb{C}^{2^{n}}:g\ket{\psi}=\ket{\psi},~{}\forall\,g\in\mathcal{S}\right\}.$
+[…]
 The elements in $\mathcal{S}$ are called stabilizers. The parameter $d$ represents the minimum distance of the code such that any Pauli error of weight less than $d$ is detectable.
@@
 We say that a quantum code is highly degenerate if it has many stabilizers with low weight relative to its minimum distance.
-A syndrome decoding problem can be stated as follows: given an error syndrome of an unknown Pauli error $e$, find the most probable $\hat{e}\in{\cal G}_{n}$ that matches the syndrome.
-An error estimate $\hat{e}$ is considered valid if it matches the syndrome or $\hat{e}e\in{\cal S}$.
+A syndrome decoding problem is defined as follows: given a measured syndrome corresponding to an unknown Pauli error $e$, a decoder outputs an estimate $\hat{e}\in\mathcal{G}_{n}$.
+An error estimate $\hat{e}$ is considered valid if it reproduces the measured syndrome.
+The decoder is said to succeed if the estimate is degenerate with the true error, i.e.,
+$\hat{e}e\in{\cal S},$
+or equivalently, $\hat{e}e$ commutes with all elements of the normalizer group $N({\cal S})$. Otherwise, the decoder produces a logical error.
 Two decoding criteria are typically considered: minimum weight decoding and degenerate maximum likelihood decoding [74, 75, 76].
@@
 $\sum_{g\in{\cal S},\mathrm{wt}\left(g\right)\leq\delta}\Pr\big\{\hat{e}g\big\}.$ (1)
-Note that 00-ADD reduces to minimum weight decoding, while $n$-ADD corresponds precisely to degenerate maximum likelihood decoding.
+Note that 0-ADD reduces to minimum weight decoding, while $n$-ADD corresponds precisely to degenerate maximum likelihood decoding.
 II-A Binary representations
@@
 We extend the $\varphi$ to ${\cal G}_{n}$ as follows: For $g=M_{1}\otimes M_{2}\otimes\dots\otimes M_{n}\in{\cal G}_{n}$,
-$\varphi(g)\triangleq{\mathbf{g}}=\left[\begin{array}[]{c|c}{\mathbf{g}}^{X}&{\mathbf{g}}^{Z}\end{array}\right]=\left[\begin{array}[]{c|c}{\mathbf{g}}^{X}_{1}\cdots{\mathbf{g}}^{X}_{n}&{\mathbf{g}}^{Z}_{1}\cdots{\mathbf{g}}^{Z}_{n}\end{array}\right],$
-where ${\mathbf{g}}^{X}\in\{0,1\}^{n}$ and ${\mathbf{g}}^{Z}\in\{0,1\}^{n}$ are the indicator vectors of $X$ and $Z$ components of $g$, respectively, such that $\left[\begin{array}[]{c|c}{\mathbf{g}}_{j}^{X}&{\mathbf{g}}_{j}^{Z}\end{array}\right]=\varphi(M_{j}).$
-A check matrix of ${\cal C}({\cal S})$ corresponding a set of stabilizer generators $\{g_{i}\}_{i=1}^{m}$ is an $m\times 2n$ binary matrix ${\mathbf{H}}$ with rows $\varphi(g_{1}),\dots,\varphi(g_{m})$.
-An error $e\in{\cal G}_{n}$ can be represented by ${\mathbf{e}}=\varphi(e)$ and its error syndrome is
+$\varphi(g)\triangleq{\mathbf{g}}=\left[\begin{array}[]{c|c}{\mathbf{g}}^{X}&{\mathbf{g}}^{Z}\end{array}\right]=\left[\begin{array}[]{c|c}a_{1}\cdots a_{n}&b_{1}\cdots b_{n}\end{array}\right]\in\{0,1\}^{2n},$
+where ${\mathbf{g}}^{X}=(a_{1},\dots,a_{n})\in\{0,1\}^{n}$ and ${\mathbf{g}}^{Z}=(b_{1},\dots,b_{n})\in\{0,1\}^{n}$ are the indicator vectors of $X$ and $Z$ components of $g$, respectively, such that $\varphi(M_{j})=\left[\begin{array}[]{c|c}a_{j}&b_{j}\end{array}\right].$
+A check matrix of ${\cal C}({\cal S})$ corresponding to a set of stabilizer generators $\{g_{i}\}_{i=1}^{m}$ is an $m\times 2n$ binary matrix $\mathbf{H}$} with rows $\varphi(g_{1}),\dots,\varphi(g_{m})$.
+An error $e\in{\cal G}_{n}$ can be represented by ${\mathbf{e}}=\varphi(e)\in\{0,1\}^{2n}$ and its error syndrome is
 ${\mathbf{s}}={\mathbf{H}}\Lambda{\mathbf{e}}^{\top}\in\{0,1\}^{m\times 1},$ (2)
@@
 The decoding problem is to solve the system of linear equations (2) with $2n$ binary variables in ${\mathbf{e}}\in\{0,1\}^{2n}$ that are most probable.
-Note that ${\mathbf{H}}$ is of rank $n-k$ so the degree of freedom of this binary system is $n+k$.
-An error $e\in{\cal G}_{n}$ is classified as a logical error for a decoder if the decoder fails or outputs $\hat{e}\in{\cal G}_{n}$ such that
-${\mathbf{H}}\Lambda({\mathbf{e}}+\hat{{\mathbf{e}}})^{\top}=\mathbf{0}$ (3)
+Note that $\mathbf{H}$} is of rank nkn-k so the degree of freedom of this binary system is $n+k$.
+An error $e\in{\cal G}_{n}$ is classified as a logical error for a decoder if the decoder fails to produce a valid estimate, or outputs an estimate $\hat{e}\in{\cal G}_{n}$ such that
+${\mathbf{H}}\Lambda{\mathbf{e}}^{\top}={\mathbf{H}}\Lambda\hat{{\mathbf{e}}}^{\top}$ (3)
 but
-${\mathbf{L}}\Lambda({\mathbf{e}}+\hat{{\mathbf{e}}})^{\top}\neq\mathbf{0}.$ (4)
+${\mathbf{L}}\Lambda{\mathbf{e}}^{\top}\neq{\mathbf{L}}\Lambda\hat{{\mathbf{e}}}^{\top}.$ (4)
+We define the logical syndrome of an error $\mathbf{e}$} as
+${\mathbf{L}}\Lambda{\mathbf{e}}^{\top}\in\{0,1\}^{2k\times 1}.$ (5)
 II-B Belief Propagation (BP) Decoding
 We explain how a quaternary BP algorithm works in the following [13], which is computed in linear domain. A log-likelihood version of BP can be found in [14].
-Given a check matrix ${\mathbf{H}}$, a syndrome ${\mathbf{s}}$, and a depolarizing error rate $\epsilon$, BP performs the following steps:
+Given a check matrix $\mathbf{H}$}, a syndrome $\mathbf{s}$}, and the prior error distributions ${\mathbf{p}}_{i}\in\mathbb{R}^{4}$ for qubit $i\in\{1,2,\dots,n\}$, BP performs the following steps:
   1. 1.
-    An initial error distribution ${\mathbf{p}}_{i}\in\mathbb{R}^{4}$ for qubit $i\in\{1,2,\dots,n\}$ is chosen for BP.
+    The belief distribution vector ${\mathbf{q}}_{i}=(q_{i}^{I},q_{i}^{X},q_{i}^{Y},q_{i}^{Z})$ is updated by BP, using the parity-check messages and the prior ${\mathbf{p}}_{i}$, for $i=1,2,\ldots,n$.
   2. 2.
-    The belief distribution vector ${\mathbf{q}}_{i}=(q_{i}^{I},q_{i}^{X},q_{i}^{Y},q_{i}^{Z})$ will be updated by BP, using the messages of parity-check satisfaction and the channel statistics ${\mathbf{p}}_{i}$ for $i=1,2,\ldots,n$.
+[…]
+These two steps are iterated for at most $\mathrm{T}$} iterations, where $\mathrm{T}$} is chosen in advance.
+If no valid estimate is obtained after $\mathrm{T}$} iterations, BP declares a decoding failure.
+III The Reliable Subset Reduction Method
+In practice, a BP decoder typically either converges within a small number of iterations, on the order of $O(\log\log n)$ on average, or remains unsettled until the maximum number of iterations is reached.
+In the latter case, although BP fails to converge globally, most error variables have already been decided with high confidence.
+From the BP output, a large fraction of variables are therefore very reliable and can be regarded as part of the solution.
+The remaining unsettled variables are usually associated with trapping sets [30], and additional post-processing techniques are required to resolve them.
+In this section, we estimate the expected size of these unsettled error variables using density evolution [58, 59].
+III-A Density-Evolution-Guided Reliability Heuristic
+We consider a family of regular LDPC codes over a quaternary symmetric channel and use the syndrome-based Gallager-A algorithm [15] as a heuristic approximation of quantum stabilizer decoding under depolarizing noise.
+Density evolution is then employed to characterize the behavior of this iterative decoder.
+Gallager-A is a hard-decision message-passing algorithm defined on a Tanner graph.
+In our setting, variable nodes $v_{j}$ represent Pauli errors on physical qubits, while check nodes $c_{i}$ represent stabilizer generators.
+Since a Pauli operator either commutes or anticommutes with a stabilizer, all messages are binary: $1$ indicates anti-commutation and 0 indicates commutation.
+We consider a syndrome-based quaternary–binary variant of Gallager-A suitable for CSS-type stabilizer codes, where each stabilizer contains only $X$ or $Z$ operators.
+Given a check matrix $\mathbf{H}$} and syndrome ${\mathbf{s}}=(s_{1},\dots,s_{m})$, the decoding proceeds as follows.
+  1. 1.
+    Initialization: Set estimated error $\hat{E}_{j}=I$ for all qubits and initialize all variable-to-check messages
+    $m_{v_{j}\to c_{i}}^{(0)}=0.$
+  2. 2.
+    Check update: At iteration $\ell$, each check node computes a recommendation message
+    $m_{c_{i}\to v_{j}}^{(\ell)}=s_{i}\oplus\bigoplus_{k\in{\cal I}(c_{i})\setminus\{j\}}m_{v_{k}\to c_{i}}^{(\ell)}.$
   3. 3.
-    At each iteration, an error estimate $\hat{{\mathbf{e}}}\in\{0,1\}^{2n}$ is made according to the hard decision on ${\mathbf{q}}_{i}$.
-If a valid error estimate is generated, it will be accepted as the solution; otherwise, go to the next iteration starting from step 2).
-Steps 2) and 3) will be iterated for a maximum number of ${\mathrm{T}}$ iterations, where ${\mathrm{T}}$ is chosen in advance.
-If no valid estimate is obtained after ${\mathrm{T}}$ iterations, BP will claim a failure.
-III Order-statistic decoding based on quaternary belief propagation
-In the aforementioned syndrome decoding problem, we aim to solve a system of linear equations given by (2), for a given parity-check matrix ${\mathbf{H}}$ and an error syndrome ${\mathbf{s}}$ with $2n$ binary variables in ${\mathbf{e}}$.
-The linear system has $n-k$ independent constraints, impliying that there are $n+k$ degrees of freedom in the vector ${\mathbf{e}}$.
-Given the values of these $n+k$ bits, one can reconstruct ${\mathbf{e}}$.
-If BP fails to provide a valid error estimate after ${\mathrm{T}}$ iterations, OSD will be utilized.
-In the context of OSD, a critical step involves identifying and sorting the accurate and reliable coordinates to establish $n+k$ linearly independent bits.
-Then, we can use these coordinates to generate a valid error that matches the syndrome.
-To accomplish this, it is necessary to establish the concept of reliability order as a means of prioritizing the coordinates based on their reliability.
-III-A Reliability order based on BP4
-To utilize the quaternary distribution generated by BP4 for each qubit error in OSD, we propose a method that employs the hard-decision history from all BP iterations and the output probabilities from the last BP iteration.
-Suppose that BP fails to provide a valid error after a maximum of ${\mathrm{T}}$ iterations.
-At this point, we have the output distribution ${\mathbf{q}}_{i}=(q_{i}^{I},q_{i}^{X},q_{i}^{Y},q_{i}^{Z})$, $i=1,2,\dots,n$ from the last BP iteration.
-Additionally, we have the hard-decision results for all ${\mathrm{T}}$ iterations, denoted by ${\mathbf{h}}_{i}^{(j)}\in\{I,X,Y,Z\}$ for $i=1,\dots,n$ and $j=1,\dots,{\mathrm{T}}$.
-For example, ${\mathbf{h}}_{i}^{(1)},{\mathbf{h}}_{i}^{(2)},\dots,{\mathbf{h}}_{i}^{({\mathrm{T}})}$ represent the hard-decision result history of BP at qubit $i$ and ${\mathbf{h}}_{1}^{(j)},{\mathbf{h}}_{2}^{(j)},\dots,{\mathbf{h}}_{n}^{(j)}$ represent the hard-decision results of BP at the $j$-th iteration.
-Our objective is to assign reliability for each Pauli $X$ and $Z$ error. Following the reliability definitions in [1, 43], we define two types of measures in the following.
+    Variable update: Each variable node $v_{j}$ updates its estimate according to the unanimous recommendations of its neighboring checks:
+    • $X$-flip: if all incoming messages from $Z$-type checks are $1$, set $\hat{E}_{j}\leftarrow\hat{E}_{j}X$;
+    • $Z$-flip: if all incoming messages from $X$-type checks are $1$, set $\hat{E}_{j}\leftarrow\hat{E}_{j}Z$.
+    (If both updates occur, a $Y$ correction is effectively applied up to a global phase.)
+  4. 4.
+    Message update: Set $m_{v_{j}\to c_{i}}^{(\ell+1)}=1$ if $\hat{E}_{j}$ anticommutes with the stabilizer of $c_{i}$, and 0 otherwise.
+  5. 5.
+    Termination: The algorithm stops if the current estimate $\hat{\mathbf{E}}$ is valid.
+For a CSS-type stabilizer code, $X$-type and $Z$-type errors are processed independently by the modified Gallager-A algorithm.
+The correlations between $X$ and $Z$ errors induced by the depolarizing channel are therefore neglected at this stage.
+Accordingly, we focus on the analysis of $X$-error decoding, which reduces to a classical binary decoding problem: estimating an $n$-bit vector ${\mathbf{e}}=(e_{1},\dots,e_{n})\in\{0,1\}^{n}$ from its syndrome.
+Suppose that the $Z$-type stabilizers define a $(d_{c},d_{v})$-regular LDPC code, where each check node is connected to $d_{c}$ variable nodes and each variable node is connected to $d_{v}$ check nodes.
+Proposition 2.
+[Heuristic Density Evolution Analysis] Consider a quantum code of length $n$ under depolarizing noise with rate $\epsilon$.
+Assume that the Tanner graph induced by the $Z$ stabilizers is locally tree-like and that decoding is performed using the modified Gallager-A algorithm.
+Let $x^{(\ell)}$ denote the probability that a variable-to-check message is incorrect at iteration $\ell$, and let $y^{(\ell)}$ denote the probability that a check-to-variable message is incorrect.
+Then the message error probabilities evolve according to
+$y^{(\ell)}$ $=\frac{1-(1-2x^{(\ell)})^{d_{c}-1}}{2},$ (6)
+$x^{(\ell+1)}$ $=p_{a}\big[1-(1-y^{(\ell)})^{d_{v}-1}\big]+(1-p_{a})(y^{(\ell)})^{d_{v}-1},$ (7)
+where $p_{a}=\frac{2\epsilon}{3}$ is the probability that a Pauli error anticommutes with a $Z$-type stabilizer, and $x^{(0)}=p_{a}$.
+The recursion (6)–(7) follows from the standard density-evolution analysis of Gallager-A decoding for regular LDPC codes over the binary symmetric channel [15, 58, 59].
+At error rates below the heuristic threshold, density evolution predicts that most variable nodes converge rapidly, leaving a fraction proportional to the channel error probability $\epsilon$ unresolved.
+The expected number of erroneous variable nodes at iteration $\ell$ is $n\cdot x^{(\ell)}$, and after an average convergence time $\ell_{\mathrm{succ}}$ the residual error set has size roughly $n\cdot x^{(\ell_{\mathrm{succ}})}$.
+For finite-length codes, cycles in the Tanner graph cause deviations from the infinite-length density evolution prediction.
+As $\epsilon$ decreases, $x^{(\ell_{\mathrm{succ}})}$ rapidly vanishes, leaving extremely sparse residual errors.
+Unlike classical LDPC ensembles, stabilizer Tanner graphs inherently contain many short cycles due to commutation constraints.
+As a result, the above density-evolution analysis should be interpreted as a heuristic guideline rather than a rigorous asymptotic prediction.
+Although one could in principle extend DE to full belief propagation with real-valued LLRs, the Gallager-A approximation already captures the main qualitative behavior of the decoding process.
+Nonetheless, our heuristic results serve as a useful reference for BP algorithms that exploit memory effects to mitigate the impact of short cycles [32, 77].
+III-B Reliable subset reduction for the syndrome decoding problem
+In the syndrome decoding problem, we seek a binary vector ${\mathbf{e}}\in\{0,1\}^{2n}$ satisfying
+${\mathbf{H}}\Lambda{\mathbf{e}}^{\top}={\mathbf{s}},$
+where $\mathbf{H}$} has rank nkn-k and hence the solution space has dimension $n+k$.
+When BP fails to converge to a valid solution within $\mathrm{T}$} iterations, a postprocessing algorithm such as OSD is typically required.
+Motivated by density evolution, we expect that after a few BP iterations, most variables have already converged to highly reliable values, while only a small subset remains ambiguous.
+We denote by $R\subset\{e_{1},\dots,e_{2n}\}$ a subset of reliable variables, whose values are assumed to be correct and fixed.
+The remaining variables are treated as unknowns.
+By eliminating the reliable variables from the parity-check equations, the original syndrome decoding problem can be reduced to a much smaller linear system.
+We refer to this procedure as reliable subset reduction (RSR).
+The number of remaining variables after reduction is called the effective length.
+Algorithm 1 RSR
+Input: parity-check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$, syndrome $\mathbf{s}$}, reliable subset $R$.
+Output: reduced matrix $\tilde{{\mathbf{H}}}$, reduced syndrome $\tilde{{\mathbf{s}}}$, and permutation $\sigma$.
+Steps:
+1:Let $\sigma$ be a column permutation such that $\sigma({\mathbf{e}})=\begin{bmatrix}{\mathbf{e}}^{\prime}&{\mathbf{e}}^{R}\end{bmatrix}$, where ${\mathbf{e}}^{R}$ is the reliable subset of $v$ bits, and $\mathbf{e}$$^{\prime}$ is the remaining 2nv2n-v bits.
+Let $\sigma(\hat{{\mathbf{e}}})=\begin{bmatrix}\hat{{\mathbf{e}}}^{\prime}&{\mathbf{e}}^{R}\end{bmatrix}$ be a permuted error estimate, which contains the reliable bits in ${\mathbf{e}}^{R}$.
+2:Let $\tau$ be a row permutation such that $\tau(\sigma({\mathbf{H}}\Lambda))=\begin{bmatrix}\tilde{{\mathbf{H}}}&{\mathbf{B}}\\ \mathbf{0}&{\mathbf{C}}\\ \end{bmatrix}$ for $\tilde{{\mathbf{H}}}\in\{0,1\}^{m^{\prime}\times(2n-v)}$ with $m^{\prime}\leq m$.
+The $m-m^{\prime}$ rows $\begin{bmatrix}\mathbf{0}&{\mathbf{C}}\end{bmatrix}$ correspond to parity checks related only to the reliable bits.
+Assume that $\tau({\mathbf{s}})=\begin{bmatrix}{\mathbf{s}}^{\prime\prime}\\ {\mathbf{s}}^{R}\end{bmatrix}$, where ${\mathbf{s}}^{\prime\prime}\in\{0,1\}^{m^{\prime}\times 1}$, and ${\mathbf{s}}^{R}\in\{0,1\}^{(m-m^{\prime})\times 1}$ are the syndrome bits corresponding to $\begin{bmatrix}\mathbf{0}&{\mathbf{C}}\end{bmatrix}$.
+3:Verify the parity check condition related to the highly reliable bits.
+if ${\mathbf{s}}^{R}\neq{\mathbf{C}}\big({\mathbf{e}}^{R}\big)^{\top}$ then return “Failure";
+else we have a new parity-check condition
+$\tilde{{\mathbf{s}}}$= $\tilde{{\mathbf{H}}}\big(\hat{{\mathbf{e}}}^{\prime}\big)^{\top},$ (8)
+  where $\tilde{{\mathbf{s}}}={\mathbf{s}}^{\prime\prime}-{\mathbf{B}}\big({\mathbf{e}}^{R}\big)^{\top}$.
+  if (8) is solvable then return $\tilde{{\mathbf{s}}}$, $\tilde{{\mathbf{H}}}$, and $\sigma$;
+  else return “Failure".
+  end if
+end if
+Equation (8) follows directly from the block decomposition
+$\tau(\sigma({\mathbf{H}}))=\begin{bmatrix}\tilde{{\mathbf{H}}}&{\mathbf{B}}\\ \mathbf{0}&{\mathbf{C}}\end{bmatrix}$,
+which separates the contributions of the unknown variables $\mathbf{e}$$^{\prime}$ and the fixed reliable variables ${\mathbf{e}}^{R}$.
+There are two possible failure modes in RSR:
+  1. (i)
+    Stage-1 failure: the reliable bits are inconsistent with the syndrome, i.e., ${\mathbf{s}}^{R}\neq{\mathbf{C}}({\mathbf{e}}^{R})^{\top}$.
+  2. (ii)
+    Stage-2 failure: the reduced linear system $\tilde{{\mathbf{H}}}(\hat{{\mathbf{e}}}^{\prime})^{\top}=\tilde{{\mathbf{s}}}$ is rank-deficient or otherwise unsolvable.
+Both conditions are explicitly verified in Step 3 of Algorithm 1. The stage-1 failure corresponds to checking the sparse consistency condition
+${\mathbf{s}}^{R}={\mathbf{C}}({\mathbf{e}}^{R})^{\top},$
+which only requires a sparse matrix–vector multiplication and can be performed efficiently.
+The stage-2 failure concerns the rank and pivot structure of $\tilde{{\mathbf{H}}}$.
+In many applications, this information must be extracted in order to solve the reduced linear system, for example through Gaussian elimination or related procedures.
+Therefore, this check can often be performed with little or no additional overhead when RSR is used together with a downstream decoder that requires such a step (e.g., OSD [44], LSD [55]).
+If RSR succeeds, the reduced system
+$\tilde{{\mathbf{H}}}(\hat{{\mathbf{e}}}^{\prime})^{\top}=\tilde{{\mathbf{s}}}$
+is passed to a downstream decoder to obtain an estimate $\hat{{\mathbf{e}}}^{\prime}$. The full error estimate is recovered as
+$\hat{{\mathbf{e}}}=\sigma^{-1}\big(\hat{{\mathbf{e}}}^{\prime},{\mathbf{e}}^{R}\big).$
+III-C Reliability order based on BP4
+To quantify the reliability of each error-variable estimate, we combine two sources of information: (i) the hard-decision trajectory of BP across all iterations, and (ii) the belief distribution produced at the final BP iteration.
+Suppose that BP terminates without finding a valid solution after a maximum of $\mathrm{T}$} iterations.
+At this point, BP provides a belief distribution ${\mathbf{q}}_{i}=(q_{i}^{I},q_{i}^{X},q_{i}^{Y},q_{i}^{Z})$ for each qubit $i=1,\dots,n$, as well as a sequence of hard decisions ${\mathbf{h}}_{i}^{(j)}\in\{I,X,Y,Z\}$ for $j=1,\dots,{\mathrm{T}}$.
+Here ${\mathbf{h}}_{i}^{(j)}$ denotes the Pauli error assigned to qubit $i$ at iteration $j$.
+Our goal is to assign a reliability score separately for the $X$ and $Z$ components of each qubit. Following [1, 44], we introduce two complementary measures.
 Definition 3.
@@
 This means that the error at qubit 1 persisted unchanged as $Z$ for three iterations when BP stops.
+A larger $\eta_{i}$ indicates that the estimate at qubit $i$ has stabilized over a longer period and is therefore more reliable.
 When BP fails, the belief distributions ${\mathbf{q}}_{i}=(q_{i}^{I},q_{i}^{X},q_{i}^{Y},q_{i}^{Z})$ will be utilized in the reliability measure.
@@
 for $i\in\{1,2,\dots,n\}$.
-Note that $\{q_{i}^{X}+q_{i}^{Y},q_{i}^{I}+q_{i}^{Z}\}$ is a binary distribution indicating the likelihood of an $X$ error occuring at the $i$-th qubit and similarly for $\{q_{i}^{X}+q_{i}^{Y},q_{i}^{I}+q_{i}^{Z}\}$.
+Here $\{q_{i}^{X}+q_{i}^{Y},\;q_{i}^{I}+q_{i}^{Z}\}$ forms a binary distribution corresponding to whether an $X$ error is present on qubit $i$, and similarly $\{q_{i}^{Z}+q_{i}^{Y},\;q_{i}^{I}+q_{i}^{X}\}$ corresponds to a $Z$ error.
+Definition 5.
+An error bit ${\mathbf{e}}_{i}^{a}$, where $i\in\{1,\dots,n\}$ and $a\in\{X,Z\}$, is said to be highly reliable if its hard-decision reliability is either $\mathrm{T}$} or ${\mathrm{T}}+1$, and its soft reliability $\phi^{a}(i)$ satisfies
+$\phi^{a}(i)\geq\theta,$
+where $\theta\in(0,1)$ is a predefined reliability threshold.
+An error variable has hard reliability $\mathrm{T}$} if its hard decision remains constant throughout all BP iterations.
+A value of ${\mathrm{T}}+1$ occurs when the variable stays equal to the identity operator $I$ from initialization through all iterations.
+The threshold $\theta$ is chosen close to one to ensure that a highly reliable coordinate is not only stable in its hard-decision history but also strongly supported by the final soft information.
+Now we define a reliability order for the bits in ${\mathbf{e}}=\left[\begin{array}[]{c|c}{\mathbf{e}}^{X}&{\mathbf{e}}^{Z}\end{array}\right]\in\{0,1\}^{2n}$ based on the outputs of BP4, using the aforementioned hard and soft reliability functions.
+Definition 6.
+For $1\leq i,j\leq n$ and $a,b\in\{X,Z\}$, the error bit ${\mathbf{e}}_{i}^{a}$ is said to be more reliable than ${\mathbf{e}}_{j}^{b}$ if either
+$\eta_{i}>\eta_{j},$
+or
+$\eta_{i}=\eta_{j}\quad\text{and}\quad\phi^{a}(i)\geq\phi^{b}(j).$
+That is, the stability of the hard-decision trajectory at the qubit level provides the primary ranking, while the final soft reliability acts as a tie-breaker between individual $X$ and $Z$ components.
 Our modified BP4 algorithm is presented in Algorithm 2.
+There exist various reliability metrics based on the (weighted) accumulation of log-likelihood ratio (LLR) signs across BP iterations [78, 79, 80, 81].
+Our hard-decision reliability can be viewed as a natural extension of this idea, focusing on the length of the most recent run in the hard-decision sequence.
+The reliability order in Definition 6 can also be adapted to incorporate alternative metrics, such as Entropy or Max [44].
+In these variants, the reliability of qubit $i$ is quantified either by the quaternary entropy of its belief distribution $\mathbf{q}_{i}$ (Entropy) or by the maximum probability within the distribution (Max).
 Algorithm 2 BP4
-Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, error rate $\epsilon$, maximum number of iterations ${\mathrm{T}}$.
-Output: A valid error estimate $\hat{{\mathbf{e}}}\in\{0,1\}^{2n}$, belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$ and hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$.
+Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, maximum number of iterations $\mathrm{T}$}, and initial error distributions ${\mathbf{p}}_{i}\in\mathbb{R}^{4}$ for $i=1,\dots,n$.
+[…]
 Initialization:
-Choose initial error distributions ${\mathbf{p}}_{i}\in\mathbb{R}^{4}$ for $i=1,\dots,n$.
-Let $\bm{\eta}_{i}=1$ for $i=1,\dots,n$.
-Let ${\mathbf{h}}_{i}=I$ for $i=1,\dots,n$.
-Let ${\mathbf{q}}_{i}=\mathbf{0}$ for $i=1,\dots,n$.
+Let $\bm{\eta}_{i}=1$ for $i=1,\dots,n$. Let ${\mathbf{h}}_{i}=I$ for $i=1,\dots,n$ be the initial hard decisions. Let ${\mathbf{q}}_{i}=\mathbf{0}$ for $i=1,\dots,n$.
 Steps:
@@
 return “Failure", $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, and $\bm{\eta}$. $\triangleright$ BP fails.
-Now we define a reliability order for the bits in ${\mathbf{e}}=\left[\begin{array}[]{c|c}{\mathbf{e}}^{X}&{\mathbf{e}}^{Z}\end{array}\right]\in\{0,1\}^{2n}$ based on the outputs of BP4, using the aforementioned hard and soft reliability functions.
-Definition 4.
-For $1\leq i,j\leq n$ and $a,b\in\{X,Z\}$, error bit ${\mathbf{e}}_{i}^{a}$ is said to be more reliable than error bit ${\mathbf{e}}_{j}^{b}$ if $\bm{\eta}_{i}>\bm{\eta}_{j}$, or if $\phi^{a}(i)\geq\phi^{b}(j)$ when $\bm{\eta}_{i}=\bm{\eta}_{j}$.
-This definition means that the hard-decision history reliability is used to rank the error coordinates, with the soft reliability serving as a tie-breaker.
+IV Order-statistic decoding based on quaternary belief propagation
+Once BP fails to produce a valid error estimate, combinatorial post-processing techniques are required.
+These techniques operate on the residual errors left by BP and can include methods such as OSD, LSD, or other syndrome-based decoders.
+In this section, we first review the basic OSD procedure, which relies on the reliability order defined in Section III-C.
+When combined with the RSR method introduced in Section III-B, OSD can operate on a reduced system of variables, thereby lowering computational complexity.
+In the context of OSD, a critical step involves identifying and sorting the accurate and reliable coordinates to establish $n+k$ linearly independent bits.
+Then, we can use these coordinates to generate a valid error that matches the syndrome.
+To accomplish this, we use the reliability orders defined in Section III-C.
 Let $\tilde{{\mathbf{e}}}$ be the hard-decision output from BP4 after $\mathrm{T}$} iterations.
 If $\tilde{{\mathbf{e}}}$ does not match the syndrome, OSD will be applied.
-We now present our OSD algorithm based on the reliability orders as in Algorithm 2, which is referred to as OSD4-00.
+We now present our OSD algorithm based on the reliability orders as in Algorithm 3, which is referred to as OSD4-0.
 Algorithm 3 OSD4-0
-Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$ of rank $n-k$, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, hard-decision vector $\tilde{{\mathbf{e}}}\in\{0,1\}^{2n}$, and belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$.
-Output: A valid error estimate $\hat{{\mathbf{e}}}\in\{0,1\}^{2n}$.
+Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$ of rank nkn-k, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, hard-decision vector $\tilde{{\mathbf{e}}}\in\{0,1\}^{2n}$, and belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$.
+[…]
 Steps:
-1:Sort the error bits in ascending order of reliability.
-Construct a column permutation function $\pi$ corresponding to the sorting order.
-Calculate $\pi(\tilde{{\mathbf{e}}})$ and $\pi({\mathbf{H}}\Lambda)$.
-2:Perform Gaussian elimination on $\pi({\mathbf{H}}\Lambda)$ and let the output be denoted by $R(\pi({\mathbf{H}}\Lambda))$, where $R$ represents the corresponding row operations in the Gaussian elimination.
-If necessary, apply a column permutation function $\mu$ to ensure that the first $n-k$ columns of the resulting matrix are linearly independent.
+[…]
+2:Perform Gaussian elimination on $\pi({\mathbf{H}}\Lambda)$ and let the output be denoted by $\rho(\pi({\mathbf{H}}\Lambda))$, where $\rho$ represents the corresponding row operations in the Gaussian elimination.
+If necessary, apply a column permutation function $\mu$ to ensure that the first nkn-k columns of the resulting matrix are linearly independent.
 Thus we have
-$\mu(R(\pi({\mathbf{H}}\Lambda)))=\left[\begin{array}[]{cc}{\mathbf{I}}_{n-k}&{\mathbf{A}}\\ \mathbf{0}&\mathbf{0}\end{array}\right],$ (7)
-for some ${\mathbf{A}}\in\{0,1\}^{(n-k)\times(n+k)}$, where the last $m-(n-k)$ rows are all zeros.
-3:Let ${\mathbf{s}}^{\prime}=R({\mathbf{s}})$ so that the parity-check condition (2) becomes
-${\mathbf{s}}^{\prime}=\mu(R(\pi({\mathbf{H}}\Lambda)))(\mu(\pi({\mathbf{e}})))^{\top}.$ (8)
-Remove the last $m-(n-k)$ entries of ${\mathbf{s}}^{\prime}$.
-4:Let $\mu(\pi(\tilde{{\mathbf{e}}}))=\left[\tilde{{\mathbf{e}}}^{U}~{}~{}\tilde{{\mathbf{e}}}^{R}\right]$, where $\tilde{{\mathbf{e}}}^{U}$ represents the independent and unreliable subset of $n-k$ bits and $\tilde{{\mathbf{e}}}^{R}$ is the reliable subset of $n+k$ bits.
-Use the reliable subset $\tilde{{\mathbf{e}}}^{R}$ and the modified parity-check condition 8 to select feasible unreliable bits
+$\mu(\rho(\pi({\mathbf{H}}\Lambda)))=\left[\begin{array}[]{cc}{\mathbf{I}}_{n-k}&{\mathbf{A}}\\ \mathbf{0}&\mathbf{0}\end{array}\right],$ (11)
+for some ${\mathbf{A}}\in\{0,1\}^{(n-k)\times(n+k)}$, where the last m(nk)m-(n-k) rows are all zeros.
+3:Let ${\mathbf{s}}^{\prime}=\rho({\mathbf{s}})$ so that the parity-check condition (2) becomes
+${\mathbf{s}}^{\prime}=\mu(\rho(\pi({\mathbf{H}}\Lambda)))(\mu(\pi({\mathbf{e}})))^{\top}.$ (12)
+Remove the last m(nk)m-(n-k) entries of $\mathbf{s}$$^{\prime}$.
+4:Let $\mu(\pi(\tilde{{\mathbf{e}}}))=\left[\tilde{{\mathbf{e}}}^{U}~~\tilde{{\mathbf{e}}}^{R}\right]$, where $\tilde{{\mathbf{e}}}^{U}$ represents the independent and unreliable subset of nkn-k bits and $\tilde{{\mathbf{e}}}^{R}$ is the reliable subset of $n+k$ bits.
+Use the reliable subset $\tilde{{\mathbf{e}}}^{R}$ and the modified parity-check condition 12 to select feasible unreliable bits
 $\tilde{{\mathbf{e}}}^{U}={{\mathbf{s}}^{\prime}}^{\top}\oplus{\tilde{{\mathbf{e}}}^{R}}{\mathbf{A}}^{\top}.$ (13)
@@
 $\hat{{\mathbf{e}}}\triangleq\pi^{-1}\left(\mu^{-1}\left(\left[{{\mathbf{s}}^{\prime}}^{\top}\oplus{\tilde{{\mathbf{e}}}^{R}}{\mathbf{A}}^{\top}\qquad\tilde{{\mathbf{e}}}^{R}\right]\right)\right).$ (14)
-We also define a reliability order that accounts for only the soft reliability functions of BP4.
-Definition 5.
-Error bit ${\mathbf{e}}_{i}^{a}$ is said to be more reliable than error bit ${\mathbf{e}}_{j}^{b}$ in the marginal distribution if $\phi^{a}(i)\geq\phi^{b}(j)$.
-The OSD based on Definition 5 is referred to as mOSD4.
-Other alternative metrics can also be considered, such as
-[…]
-[…]
-The reliability of qubit $i$ is assessed based on the quaternary entropy of its belief distribution $\mathbf{q}_{i}$ (Entropy) or based on the maximum probability of the distribution (Max).
-For our purpose, the hard-decision output of BP is first sorted in the Pauli basis based on the reliability metric (either Entropy or Max) and then transformed into a binary vector for OSD [43].
-We will compare these reliability metrics in Section VI-D.
-III-B OSD4-$w$
+IV-A OSD4-$w$
 If some bits in the reliable subset are incorrect, additional processing is necessary.
@@
 Figure 1: A tree with nodes consisting of all bit strings of length 4 and with at most weight 2.
-The root node is $0$.
+The root node is $00$.
 A child node is obtained by flipping one of the zero bits from its parent node that come after the last bit that is one.
-III-C Complexity of OSD4
-We analyze the complexity of OSD based on 00-ADD in the following.
+IV-B Complexity of OSD4
+We analyze the complexity of OSD based on 0-ADD in the following.
 The complexity of OSD4-0 is dominated by the Gaussian elimination step, which is then $O(2nm^{2})=O(n^{3})$, assuming $m=O(n)$.
@@
 Additionally, by running through $\sum_{j=0}^{w}\binom{n+k}{j}$ possibilities using the DFS algorithm, the OSD4-$w$ algorithm has a complexity of $O(n^{3}+n^{w+1})$, which simplifies to $O(n^{3})$ when $w\leq 2$.
-Algorithm 3 HRSR
-Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$ of rank $n-k$, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, hard-decision vector ${\mathbf{e}}\in\{0,1\}^{2n}$, belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, and soft reliability threshold $\theta$.
-Output: shortened variables $\tilde{{\mathbf{s}}}$, $\tilde{{\mathbf{H}}}$, ${\mathbf{e}}^{H}$, and $\sigma$.
-Steps:
-1:Use $\bm{\eta}$, ${{\mathbf{q}}_{i}}_{i=1}^{n}$, and $\theta$ to identify the highly reliable bits in ${\mathbf{e}}$ based on Definition 7. Assume that there are $v$ such highly reliable bits.
-2:Let $\sigma$ be a column permutation such that $\sigma({\mathbf{e}})=\begin{bmatrix}{\mathbf{e}}^{\prime}&{\mathbf{e}}^{H}\end{bmatrix}$, where ${\mathbf{e}}^{H}$ is the highly reliable subset of $v$ bits, and ${\mathbf{e}}^{\prime}$ is the remaining $2n-v$ bits.
-Let $\sigma(\hat{{\mathbf{e}}})=\begin{bmatrix}\hat{{\mathbf{e}}}^{\prime}&{\mathbf{e}}^{H}\end{bmatrix}$ be a permuted error estimate, which contains the highly reliable bits in ${\mathbf{e}}^{H}$.
-3:Let $\tau$ be a row permutation such that $\tau(\sigma({\mathbf{H}}\Lambda))=\begin{bmatrix}\tilde{{\mathbf{H}}}&{\mathbf{B}}\\ \mathbf{0}&{\mathbf{C}}\\ \end{bmatrix}$ for $\tilde{{\mathbf{H}}}\in\{0,1\}^{m^{\prime}\times(2n-v)}$ with $m^{\prime}\leq m$.
-The $m-m^{\prime}$ rows $\begin{bmatrix}\mathbf{0}&{\mathbf{C}}\end{bmatrix}$ correspond to parity checks related only to the highly reliable bits.
-Assume that $\tau({\mathbf{s}})=\begin{bmatrix}{\mathbf{s}}^{\prime\prime}\\ {\mathbf{s}}^{H}\end{bmatrix}$, where ${\mathbf{s}}^{\prime\prime}\in\{0,1\}^{m^{\prime}\times 1}$, and ${\mathbf{s}}^{H}\in\{0,1\}^{(m-m^{\prime})\times 1}$ are the syndrome bits corresponding to $\begin{bmatrix}\mathbf{0}&{\mathbf{C}}\end{bmatrix}$.
-4:Verify the parity check condition related to the highly reliable bits.
-if ${\mathbf{s}}^{H}\neq{\mathbf{C}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}$ then return “Failure";
-else we have a new parity-check condition
-$\tilde{{\mathbf{s}}}=$ $\tilde{{\mathbf{H}}}\big{(}\hat{{\mathbf{e}}}^{\prime}\big{)}^{\top},$ (15)
-     where $\tilde{{\mathbf{s}}}={\mathbf{s}}^{\prime\prime}-{\mathbf{B}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}$.
-     if (15) is solvable then return $\tilde{{\mathbf{s}}}$, $\tilde{{\mathbf{H}}}$, ${\mathbf{e}}^{H}$, and $\sigma$;
-     else return “Failure".
-     end if
-end if
-IV Highly reliable subset reduction
-As we increase the order $w$ of OSD, we can expect an improvement in the decoder’s performance.
-In particular, when $w=n+k$ and $\delta=n$, OSD4-$(n+k)$ using $n$-ADD becomes a maximum likelihood decoder.
-However, increasing the order $w$ also leads to an increase in time complexity.
-To address this issue, we propose an approach to reduce the problem size for OSD.
-Recall that OSD is used to solve a linear system with $2n$ binary variables and $n-k$ constraints.
-If a subset of binary variables is known, the problem size–and consequently the computational complexity–can be reduced.
-The number of binary variables in the reduced linear system is referred to as the effective length for OSD, which corresponds to the length of the remaining binary error variable vector.
-We have the following definition of highly reliable coordinates.
-Definition 7.
-An error bit ${\mathbf{e}}_{i}^{a}$, where $i\in{1,\dots,n}$ and $a\in{X,Z}$, is considered highly reliable if its hard-decision reliability equals ${\mathrm{T}}$ or ${\mathrm{T}}+1$, and its soft reliability $\phi^{a}(i)$ satisfies $\phi^{a}(i)\geq\theta$, where $\theta$ is a predetermined reliability threshold.
-An error variable has hard reliability ${\mathrm{T}}$ if BP has a constant output at all iterations; a hard reliability of ${\mathrm{T}}+1$ means the corresponding error variable is the identity during the BP process after being initialized to the identity.
-The soft reliability threshold $\theta$ is chosen close to one to ensure that a highly reliable coordinate not only exhibits an almost constant hard-decision history but also has a high soft reliability value.
-These highly reliable variables are crucial in the OSD process. Identifying these variable nodes allows us to greatly reduce the computation burden of OSD.
-These highly reliable parts provide outcomes that can be trusted and will remain unchanged throughout the OSD4 procedure with high probability.
-Consequently, we propose to exclude these highly reliable parts from the parity check relations.
-By doing so, we can avoid unnecessary calculations and prioritize our focus on the unreliable part.
-Our highly reliable subset reduction (HRSR) algorithm is given in Algorithm 3.
-Equation (15) follows from the parity-check condition (2) with
-$\begin{bmatrix}{\mathbf{s}}^{\prime\prime}\\ {\mathbf{s}}^{H}\end{bmatrix}=$ $\tau(\sigma({\mathbf{H}}\Lambda))\begin{bmatrix}\hat{{\mathbf{e}}}^{\prime}&{\mathbf{e}}^{H}\end{bmatrix}^{\top}=$ $\begin{bmatrix}\tilde{{\mathbf{H}}}\big{(}\hat{{\mathbf{e}}}^{\prime}\big{)}^{\top}+{\mathbf{B}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}\\ {\mathbf{C}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}\end{bmatrix}$
-and the conditions ${\mathbf{s}}^{H}={\mathbf{C}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}$ and $\tilde{{\mathbf{s}}}={\mathbf{s}}^{\prime\prime}-{\mathbf{B}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}$.
-If HRSR returns a reduced linear system as in (15), we then use OSD4-$w$ to solve it with hard-decision BP output ${\mathbf{e}}^{\prime}$ and corresponding reliabilities $\bm{\eta}$, $\phi^{X}$, and $\phi^{Z}$.
-Suppose that OSD4-$w$ returns $\hat{{\mathbf{e}}}^{\prime}$.
-Then the error estimate is $\sigma^{-1}\left(\begin{bmatrix}\hat{{\mathbf{e}}}^{\prime}&{\mathbf{e}}^{H}\end{bmatrix}\right)$.
-Reducing the problem size may cause the reduced check matrix to lose the ability to represent all valid codewords if its dimension becomes too small.
-To address this, in Step 4 of Algorithm 3, we verify that the removed highly reliable bits align with the corresponding syndrome bits.
-If they do not, the algorithm reports a failure.
-This method effectively reduces the overhead in OSD, especially at low error rates as will be shown in the simulation section.
-Furthermore, as the effective length in OSD is reduced, we can apply higher order OSD if only a fixed amount of computation is allowed.
-Consider, for instance, the case where the number of error candidates to be tested in OSD is $\Gamma$.
-For the original decoding problem, OSD4-$2$ has
+IV-C Reliable Subset Reduction combined with OSD
+As introduced in Section III-B, the RSR procedure removes highly reliable error variables from the parity-check constraints, producing a reduced linear system.
+This reduced system can then be solved using OSD or any other suitable combinatorial decoder to estimate the remaining unreliable variables.
+Increasing the OSD order $w$ generally improves decoder performance.
+In the extreme case, when $w=n+k$ and $\delta=n$, OSD4-$(n+k)$ with $n$-ADD achieves maximum likelihood decoding.
+However, larger $w$ comes at the cost of increased computational complexity.
+By reducing the problem size through RSR, we can lower the computational overhead in OSD, particularly at low error rates, as demonstrated in Section VII.
+Moreover, the reduced effective length allows the use of higher-order OSD under a fixed candidate budget.
+For example, consider a candidate limit of $\Gamma$.
+For the original problem, OSD4-$2$ tests
 $\Gamma=\binom{n+k}{0}+\binom{n+k}{1}+\binom{n+k}{2}$. (17)
-Using the HRSR algorithm for $v$ highly reliable bits, we have a reduced system of linear equations with $2n-v$ binary variables.
-Assume the rank of the reduced system is $n-k-(m-m^{\prime})$.
-Then we can apply OSD4-$w$ under the constraint of $\Gamma$ candidates, with $w$ given by
-$w=\operatorname*{arg\,max}_{x}\sum_{i=0}^{x}\binom{n+k-v+(m-m^{\prime})}{i}\leq\Gamma.$ (17)
-Note that even though the number of tested candidates is fixed, the complexity of OSD following the HRSR algorithm is lower due to the reduced effective length.
+If $v$ highly reliable bits are removed via RSR, the reduced system has 2nv2n-v variables and rank $n-k-(m-m^{\prime})$. The maximum OSD order $w$ that respects the same candidate budget is then
+$w=\max\left\{x:\sum_{i=0}^{x}\binom{n+k-v+(m-m^{\prime})}{i}\leq\Gamma\right\}.$ (18)
+Even with the same number of candidates, the complexity is lower due to the reduced effective length.
 V Approximate Degenerate OSD
-In this section, we discuss how degeneracy in quantum codes helps reduce the computation of OSD.
-Recall that in OSD-$w$, $\sum_{i=0}^{w}\binom{n+k}{i}$ candidates are generated, and the one that is optimal with respect to $\delta$-ADD is chosen as the output.
-As shown in Lemma 6, a new candidate can be generated from an error candidate by flipping a bit in the reliable subset and adding a certain vector to the unreliable subset.
-This corresponds to adding a logical operator in the normalizer group to the error candidate.
-By examining these logical operators, we have the following degenerate error candidate pruning conditions.
+In this section, we show how code degeneracy can be exploited to prune the candidate set in OSD.
+Recall that in OSD4-$w$, a list of $\sum_{i=0}^{w}\binom{n+k}{i}$ candidates is generated by flipping subsets of reliable bits of the OSD4-0 solution and adjusting the unreliable bits accordingly, and the best candidate is selected according to the $\delta$-ADD metric.
+In quantum stabilizer codes, however, many of these candidates correspond to degenerate errors, i.e., they differ only by stabilizers and therefore have identical logical effects.
+Testing such degenerate candidates is unnecessary and can be avoided by tracking their logical syndromes.
+Let $\mathbf{L}$} denote a binary logical matrix satisfying ${\mathbf{H}}\Lambda{\mathbf{L}}^{\top}=\mathbf{0},$ so that the logical syndrome of an error $\mathbf{e}$} is given by ${\mathbf{L}}\Lambda{\mathbf{e}}^{\top}.$
+Let $\hat{{\mathbf{e}}}$ be the OSD4-0 solution.
+As in (11), after applying the same column permutation $\pi$, row operations $\rho$, and column permutation $\mu$ used in Gaussian elimination, the logical matrix becomes
+${\mathbf{L}}^{\prime}=\mu\!\big(\rho(\pi({\mathbf{L}}\Lambda))\big).$ (19)
+Flipping a single reliable bit corresponds to adding the associated column of $\mathbf{L}$$^{\prime}$ to the logical syndrome.
+Therefore, a reliable bit contributes to a nontrivial logical change if and only if the corresponding column of $\mathbf{L}$$^{\prime}$ is nonzero.
 Theorem 8.
-Consider an $[[n,k,d]]$ stabilizer code with a check matrix given by $\begin{bmatrix}{\mathbf{I}}_{n-k}&{\mathbf{A}}\end{bmatrix}$ after Gaussian elimination and a column permutation as in (7), where ${\mathbf{A}}\in\{0,1\}^{(n-k)\times(n+k)}$ are associated with the reliable bits of OSD4-0 output.
+[Degenerate candidate pruning] Consider an $[[n,k,d]]$ stabilizer code whose parity-check matrix has been transformed into the form $\mu(\rho(\pi({\mathbf{H}}\Lambda)))=\begin{bmatrix}{\mathbf{I}}_{n-k}&{\mathbf{A}}\end{bmatrix}$ as in (11), where the last $n+k$ columns correspond to the reliable bits in OSD4-0.
+Let $\mathbf{L}$$^{\prime}$ be the transformed logical matrix defined above.
   1. 1.
-    When using $n$-ADD, only the reliable bits corresponding to nontrivial logical operators need to be considered in OSD4-$w$.
+    In $n$-ADD OSD, only reliable bits whose corresponding columns in $\mathbf{L}$$^{\prime}$ are nonzero need to be flipped when generating OSD4-$w$ candidates.
   2. 2.
-    If each logical operator induced by flipping a reliable bit of OSD4-0 output is a stabilizer, then OSD4-$w$ is equivalent to OSD4-0 for any $w$.
+    If all columns of $\mathbf{L}$$^{\prime}$ corresponding to the reliable bits are zero, then all OSD4-$w$ candidates are degenerate with the OSD4-0 solution, and OSD4-$w$ reduces to OSD4-0 for any $w$.
 Proof.
-1) In $n$-ADD, different stabilizer cosets are compared. Therefore, multiplying a stabilizer by an error candidate does not create a new nontrivial candidate.
-2) The statement is straightforward, as flipping these reliable bits results in only degenerate errors to the OSD4-0 output.
+Flipping a reliable bit produces a new error candidate whose logical syndrome differs from that of $\hat{{\mathbf{e}}}$ by the corresponding column of $\mathbf{L}$$^{\prime}$.
+In $n$-ADD, candidates that differ only by stabilizers are equivalent and need not be tested.
+1) Therefore, only reliable bits whose columns in $\mathbf{L}$$^{\prime}$ are nonzero can generate candidates in different logical cosets, and only those flips must be considered.
+2) If all such columns are zero, then flipping any reliable bit leaves the logical syndrome unchanged, so all OSD4-$w$ candidates are degenerate with $\hat{{\mathbf{e}}}$, and no higher-order OSD can improve upon OSD4-0.
+∎
 Exploiting Theorem 8-1) with $n$-ADD makes OSD4-$w$ more effective.
-By incorporating with the HRSR algorithm, we can work on a shortened check matrix $\tilde{{\mathbf{H}}}$ and higher order of OSD can be executed.
+By incorporating with the RSR algorithm, we can work on a shortened check matrix $\tilde{{\mathbf{H}}}$ and higher order of OSD can be executed.
 Consequently, (18) can be improved to $w=\operatorname*{arg\,max}_{x}\sum_{i=0}^{x}\binom{u}{i}\leq\Gamma,$ where $u\leq n+k-v+(m-m^{\prime})$ is the number of columns in the shortened check matrix corresponding to nontrivial logical operators.
-Proposition 9.
-Checking whether a logical operator induced by flipping a reliable bit of the OSD4-0 output is a stabilizer can be done in $O((n-k)k)$ time.
-Proof.
-Consider an $[[n,k,d]]$ stabilizer code with a $m\times{2n}$ check matrix ${\mathbf{H}}$ and a $2k\times 2n$ binary logical matrix ${\mathbf{L}}$.
-During the process of OSD4-0, ${\mathbf{H}}$ leads to $\begin{bmatrix}{\mathbf{I}}_{n-k}&{\mathbf{A}}\end{bmatrix}$ as in (7) using column permutations $\pi,\mu$ and row operations $R$, where ${\mathbf{A}}\in\{0,1\}^{(n-k)\times(n+k)}$ are associated with the reliable bits of OSD4-0 output.
-By (13), flipping the $j$-th reliable bit corresponds to adding the vector ${\mathbf{g}}=\begin{bmatrix}{\mathbf{A}}_{j}^{\top}&0\cdots 0&1&0\cdots 0\end{bmatrix}\in\{0,1\}^{2n}$, where ${\mathbf{A}}_{j}$ is the $j$-th column of ${\mathbf{A}}$.
-The vector ${\mathbf{g}}$ corresponds to a stabilizer if $\mu(\pi({\mathbf{L}}\Lambda)){\mathbf{g}}^{\top}=\mathbf{0}$, and otherwise, it represents a nontrivial logical operator.
-Since ${\mathbf{A}}_{j}$ is of dimension $n-k$, this computation takes $O((n-k)k)$ time.
-The computational complexity of verifying such a bit-flipping event is $O((n-k)k)$, which implies that checking all the columns of $A$ has a computational complexity of roughly the same order as Gaussian elimination.
-This approach is both effective and relatively efficient for implementing high-order OSD.
 However, the complexity of $n$-ADD is not practically feasible.
@@
 Thus, $n$-ADD can be well approximated by $4$-ADD at error rates lower than $1\%$.
-Assume that OSD4444-0 generates a list of error candidates $\mathcal{E}$.
+Assume that OSD4-0 generates a list of error candidates $\mathcal{E}$.
 Then, $4$-ADD outputs:
@@
 For codes with algebraic structures, the first few lowest-weight terms can be derived, but determining the complete weight enumerator for a general code is an NP-hard problem.
-As an alternative, we propose selecting a set of $m\geq n-k$ low-weight stabilizer generators and calculating the weight distribution of (${m\choose c}$ stabilizers, where $c$ is a constant independent of $n$.
+As an alternative, we propose selecting a set of $m\geq n-k$ low-weight stabilizer generators and calculating the weight distribution of $m\choose c$} stabilizers, where $c$ is a constant independent of $n$.
 This method provides a good approximation of the dominant terms for practical purposes.
-To maintain an overall complexity of $O(n^{3})$, the value of (${m\choose c}$ should be chosen accordingly.
-To achieve optimal execution complexity, 00-ADD is often used.
-Theorem 8 suggests focusing on the columns of $A$ corresponding to nontrivial logical operators when $n$-ADD is used.
-However, in the scenario of 00-ADD, a stabilizer may influence the output of $\text{OSD}_{4}$-$w$ since multiplying a stabilizer can reduce the weight of the error, potentially providing a better candidate in 00-ADD, even though the error coset remains unchanged.
-Thus, $\text{OSD}_{4}$-$w$ cannot simply flip bits corresponding to those nontrivial logical operators as in the case of Theorem 8-1).
-However, Theorem 8-2) still works and we may use Proposition 9 to check this condition.
-In the following we provide a weaker sufficient but operationally more efficient condition, for degenerate error candidate pruning.
+To maintain an overall complexity of $O(n^{3})$, the value of $m\choose c$} should be chosen accordingly.
+To achieve low computational complexity, OSD4-0 is often used in practice.
+Theorem 8 shows that when $n$-ADD is employed, only flips that induce nontrivial logical syndromes need to be considered.
+However, this argument does not directly apply to OSD4-0.
+In 0-ADD, candidates are compared solely by their weight.
+Multiplying an error by a stabilizer may change its weight while leaving its logical coset unchanged, and therefore a degenerate error can become a better candidate under 0-ADD.
+As a result, in OSD4-0 one cannot restrict bit flips solely based on logical syndromes as in Theorem 8-1).
+Nevertheless, the degeneracy condition in Theorem 8-2) still applies.
+We therefore introduce a weaker but computationally efficient sufficient condition for pruning degenerate candidates.
 Corollary 9.
-Let $\tilde{S}\in\{0,1\}^{m^{\prime}\times(2n-v)}$ be a shortened check matrix by the HRSR algorithm.
+Let $\tilde{S}\in\{0,1\}^{m^{\prime}\times(2n-v)}$ be a shortened check matrix by the RSR algorithm.
 Suppose that $\tilde{S}$ can be transformed into $\begin{bmatrix}{\mathbf{I}}&\tilde{{\mathbf{A}}}\end{bmatrix}$ after Gaussian elimination and a column permutation, where $\tilde{{\mathbf{A}}}$ is associated with the reliable bits.
-If all the columns of $\tilde{{\mathbf{A}}}$ are of weight less than $d-1$, then OSD4-$w$ on the corresponding reduced linear system is equivalent to OSD4-0 for any $w$.
-The operational degenerate error pruning condition in Corollary 10 provides a more operationally efficient method to determine whether order-$w$ OSD for $w>0$ is nontrivial by evaluating the weight of each column.
-This can be done in $O((n-k)\times(n+k))$.
-If there are many highly reliable bits, the number of stabilizers associated with these bits would be high.
-Consequently, the shortened check matrix $\tilde{\mathbf{S}}$ from the HRSR algorithm will have fewer rows and lower column weights.
-Therefore, Corollary 10 can be applied with high probability when the error rate is low.
-We combine all the aforementioned techniques into the approximate degenerate OSD (ADOSD) algorithm, as detailed in Algorithm 4, using 00-ADD.
-Note that if decoding performance is of higher priority, $\delta$-ADD can be utilized, and Proposition 9 can be integrated into Algorithm 4.
+If all the columns of $\tilde{{\mathbf{A}}}$ are of weight less than d1d-1, then OSD4-$w$ on the corresponding reduced linear system is equivalent to OSD4-0 for any $w$.
+V-A RSR + ADOSD
+If many bits are identified as highly reliable, the number of stabilizers associated with these bits would be high.
+Consequently, the shortened check matrix $\tilde{\mathbf{S}}$ from the RSR algorithm will have fewer rows and lower column weights.
+Consequently, the sufficient condition in Corollary 9 is satisfied with high probability at low error rates, making higher-order OSD unnecessary.
+We combine RSR, degeneracy-aware pruning, and order-adaptive OSD into a unified approximate degenerate OSD (ADOSD) framework, as detailed in Algorithm 4, using 0-ADD.
+Note that if decoding performance is of higher priority or if the problem size is small enough (e.g., after RSR reduction), $\delta$-ADD can be utilized and can be integrated into Algorithm 4.
 Algorithm 4 ADOSD4
-Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$ of rank $n-k$, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, BP output distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, soft reliability threshold $\theta$, hard-decision vector ${\mathbf{e}}\in\{0,1\}^{2n}$, maximum number of OSD flips $\Gamma$, and backup OSD order $w$.
-Output: A valid error estimate $\hat{{\mathbf{e}}}\in\{0,1\}^{2n}$.
+Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$ of rank nkn-k, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, BP output distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, soft reliability threshold $\theta$, hard-decision vector ${\mathbf{e}}\in\{0,1\}^{2n}$, maximum number of OSD flips $\Gamma$, and backup OSD order $w$.
+[…]
 Steps:
-Apply the HRSR algorithm.
-if  HRSR returns failure, then
-     use OSD4-$w$ to find an error estimate $\hat{{\mathbf{e}}}\in\{0,1\}^{2n}$ for the original problem, and return $\hat{{\mathbf{e}}}$;
-else HRSR returns ${\mathbf{s}}^{\prime}$, $\tilde{{\mathbf{H}}}$, ${\mathbf{e}}^{H}$, and $\sigma$.
-     if all the columns of $\tilde{{\mathbf{H}}}$ in reduced row echelon form have weight less than $d-1$then $w\leftarrow 0$;
+Apply the RSR algorithm.
+if RSR returns failure, then
+  apply OSD4-$w$ to the original problem, and return $\hat{{\mathbf{e}}}$;
+else RSR returns $\mathbf{s}$$^{\prime}$, $\tilde{{\mathbf{H}}}$, ${\mathbf{e}}^{R}$, and $\sigma$. Suppose $\tilde{{\mathbf{H}}}$ in reduced row echelon form is $I|\tilde{A}$].
+  if all the columns of $\tilde{A}$ have weight less than d1d-1, then $w\leftarrow 0$;
   else determine $w$ by (18).
   end if
-     Use OSD4-$w$ on ${\mathbf{s}}^{\prime}$, $\tilde{S}$, and the given data to find an error estimate $\hat{{\mathbf{e}}}^{\prime}\in\{0,1\}^{2n-v}$, and return $\sigma^{-1}\left(\begin{bmatrix}\hat{{\mathbf{e}}}^{\prime}&{\mathbf{e}}^{H}\end{bmatrix}\right)$.
+[…]
 end if
-VI Simulation results
-We simulate the performance of the MBP4+ADOSD4 scheme on LCS codes [60], GHP codes [43, 44], BB codes [56], and various 2D topological codes, including toric, rotated surface, color codes and twisted XZZX codes [23, 24, 25, 26, 27, 59, 29] (see [64, Table I]).
-We use the MBP4 decoder introduced in [32] as a predecoder for OSD.
-MBP4 features a parameter $\alpha$ that controls the step size of message updates during BP iterations: when $\alpha<1$, the step size is enlarged; when $\alpha>1$, the step size is reduced.
-For $\alpha=1$, MBP4 is equivalent to the refined BP4 in [13].
-We will specify the value of $\alpha$ only when $\alpha\neq 1$ is used.
-The optimization of $\alpha$ is discussed in Section VI-A.
-The maximum number of BP iterations, ${\mathrm{T}}$, is set to $100$ unless otherwise specified.
-Furthermore, our MBP4+ADOSD4 scheme demonstrates robustness to variations in the maximum number of iterations, as discussed in Section VI-B.
-The soft reliability threshold is set to $\theta=0.999995$ and its selection is discussed in Section VI-C.
-In ADOSD4 simulations, $\Gamma$ is chosen according to (16) to enable comparison with OSD-2.
-In Section VI-D, we compare the proposed BP+OSD schemes with various reliability measures using GHP codes.
-We then evaluate decoding performance on BB codes and topological codes in Sections VI-E, and VI-F.
-Finally, Section VI-G analyzes the execution time of these schemes, demonstrating the efficiency of ADOSD.
+VI Extension to Circuit-Level Noise via Detector Error Models
+Our framework is not confined to code-capacity noise; it admits a natural extension to realistic circuit-level noise through detector error models, as we describe below.
+At the circuit level, gates, measurements, and ancilla qubits are faulty, and quantum error correction typically involves multiple rounds of stabilizer syndrome extraction circuits.
+The decoding task is to infer the most likely error configuration given the circuit description and the observed measurement outcomes.
+This circuit-level decoding problem can be formulated as solving a linear system of equations [60, 70].
+Specifically, the error syndrome of a location error is the set of measurement outcomes produced when only that error occurs in the syndrome extraction circuits, while the logical syndrome is given by the commutation relations between the logical operators and the residual data-qubit error after circuit evolution.
+Collecting the error and logical syndromes of all basis location errors yields a circuit-level parity-check matrix.
+For an $n$-qubit stabilizer code, the resulting circuit-level decoding problem typically has size $O(n^{3})$ using $O(\sqrt{n})$ syndrome extraction rounds.
+STIM [71] provides a detector error model (DEM) for a given syndrome-extraction circuit, which yields a binary circuit-level parity-check matrix ${\mathbf{H}}_{\mathrm{DEM}}$ and a corresponding binary logical matrix ${\mathbf{L}}_{\mathrm{DEM}}$.
+The error syndrome is the binary vector of detector variables that represent the syndrome difference between two consecutive rounds.
+Moreover, location errors that produce identical error and logical syndromes are indistinguishable and are therefore merged by STIM into equivalence classes, resulting in a nonuniform probability vector ${\mathbf{p}}_{\mathrm{DEM}}$ that specifies the approximate probability of each error class.
+Since STIM is a Clifford circuit simulator, it records only commuting observables; in particular, either logical $X$ operators or logical $Z$ operators are monitored.
+In the following, we assume that logical $X$ operators are observed.
+[…]
+[…]
+[…]
+This abstraction maps a fault-tolerant stabilizer circuit to a binary parity-check matrix with independent error probabilities, which allows our decoder to be applied without modification.
+For our purposes, this binary decoding problem can be embedded into the quaternary stabilizer formalism by treating ${\mathbf{H}}_{\mathrm{DEM}}$ as $Z$-type stabilizers and ${\mathbf{L}}_{\mathrm{DEM}}$ as logical $Z$ operators, while neglecting commutation relations.
+Importantly, circuit-level DEMs are extremely sparse and highly biased: most columns correspond to single-fault events with very small support.
+As a result, when BP is applied to ${\mathbf{H}}_{\mathrm{DEM}}$, a large fraction of error variables become highly reliable, making RSR particularly effective.
+In the low-noise regime, RSR typically reduces the effective system size by one to two orders of magnitude, transforming a nominal $O(n^{3})$ circuit-level problem into a reduced system whose size scales roughly linearly with the code distance when error rate is low.
+VII Simulation results
+We simulate the performance of the MBP4+ADOSD4 scheme on LCS codes [57], BB codes [56], and various 2D topological codes, including toric, rotated surface, color, and twisted XZZX codes [23, 24, 25, 26, 27, 72, 29] (see [82, Table I]) under the code-capacity noise model, and we further simulate rotated surface codes under a circuit-level noise model.
 The decoding performance of a quantum code ${\cal C}({\cal S})$ and a decoder is assessed through the logical error rate (LER).
@@
 In the worst case, a flatter error floor may appear earlier than expected.
-Each data point in the plots is based on at least 100 logical error events in the low error rate regime.
+Each data point in the plots is based on at least 100–1000 logical error events in the low error rate regime.
 At LER $10^{-6}$, this represents an error bar less than $2\times 10^{-7}$, which is roughly the size of the markers for the data points.
 Thus the error bars are neglected for clarity.
-We observe that both parallel and serial schedules in MBP4+ADOSD4 yield similar results for toric codes and surface codes.
-However, the parallel schedule outperforms the serial schedule on the [[882, 48, 16]] GHP code.
-We believe that this is because that the parallel schedule typically results in unsettled oscillations in BP, which can be resolved by OSD, whereas the serial schedule may make premature decisions on some bits, leading to incorrect convergence in the decoding process.
-Therefore, we consider the parallel schedule for the following simulations of MBP4+ADOSD4.
-Similar observations are also made in [55, 65].
-decoder source note
-MBP4+ADOSD4 based on Algorithm 4
-MBP4+OSD4-$w$ this paper reliability metric based on Definition 4
-MBP4+mOSD4-$w$ reliability metric based on Definition 5
-AMBP4 [32] No postprocessing. Can act as
-a predecoder for ADOSD4 and OSD4
-BP+mOSD($w$,$\lambda$) [43] binary BP+ order-w OSD4
-with additional flips on $\lambda$ positions
-[…]
-a greedy search on (${\lambda\choose 2}$ positions
-BP+GDG [55] binary BP with guided decimation
-guessing based on a decision tree
-TABLE I: Various decoders compared in this paper.
-For comparison, we consider several decoders listed in Table I. Except for AMBP, we do not implement these decoders ourselves but instead compare our results with those reported in the literature.
-In [66], it was shown that a random order schedule improves BP decoding performance.
-Here, we implement the AMBP decoder with a random order serial schedule, achieving better performance than the results reported in [32].
-Note that AMBP incorporates various techniques for performance enhancement, and we do not focus on optimizing its performance in the following simulations.
-VI-A Optimization of the parameter $\alpha$ in MBP4 for ADOSD.
-In this subsection, we demonstrate how to optimize the parameter $\alpha$ of our MBP4+ADOSD4 scheme using LCS codes [60] with ${\mathrm{T}}=100$ and $\theta=0.999995$.
+We use only the parallel schedule for the following simulations of MBP4+ADOSD4.
+In contrast, [83] showed that a random-order schedule can improve BP decoding performance.
+Accordingly, we implement the AMBP decoder with a random-order serial schedule, achieving better performance than the results reported in [32].
+In ADOSD4 simulations, $\Gamma$ is chosen according to (17).
+VII-A Parameter Selection and Tuning
+We use the MBP4 decoder introduced in [32] as a predecoder for OSD.
+MBP4 features a parameter $\alpha$ that controls the step size of message updates during BP iterations: when $\alpha<1$, the step size is enlarged; when $\alpha>1$, the step size is reduced.
+For $\alpha=1$, MBP4 is equivalent to the refined BP4 in [13].
+We will specify the value of $\alpha$ only when $\alpha\neq 1$ is used.
+The optimization of $\alpha$ is discussed in Section VII-A1.
+The maximum number of BP iterations, $\mathrm{T}$}, is set to $100$ and the soft reliability threshold to $\theta=0.999995$ for simulations under the code-capacity noise model to ensure high accuracy, while for the circuit-level noise model, $\mathrm{T}$} is set to $10$ and $\theta=0.99$ to prioritize computational efficiency.
+VII-A1 Optimization of the parameter $\alpha$ in MBP4 for ADOSD
+In this subsection, we demonstrate how to optimize the parameter $\alpha$ of our MBP4+ADOSD4 scheme using LCS codes [57].
 For most nondegenerate codes, $\alpha=1$ is sufficient to achieve good decoding performance.
@@
 This optimization should be conducted for each code and potentially for each error rate, though a single value of $\alpha$ often applies across an entire code family.
-LCS codes are a specialized type of quantum code that combines the features of surface codes and lift-product constructions.
+LCS codes are a specialized type of quantum codes that combines the features of surface codes and lift-product constructions.
 Decoding LCS codes presents the most challenging problem for AMBP4 among all the quantum codes considered in this paper.
@@
 Figure 2: MBP4+ADOSD4 decoding with various values of $\alpha$ for the $[[175,7,7]]$ LCS code.
-The curve MLE is taken from [60].
-The notation $\alpha\in(1.6,-0.01,0.5)$ means that $\alpha$ is tested in the sequence $1.60,1.59,1.58,\dots,0.51,0.50$.
+The curve MLE is taken and rescaled from [57].
+The notation $\alpha\in(1.6,-0.01,0.5)$ means that $\alpha$ is tested in the sequence $1.60,1.59,1.58,\dots,0.51,0.50$ with ${\mathrm{T}}=100$ and $\theta=0.999995$.
 Figure 2 compares various decoders on the $[[175,7,7]]$ LCS code, including an MLE (independent X/Z) reference curve from [57].
+To compare with the MLE curve under $X$-only noise, we rescale the physical error rate by a factor of $3/2$ so that the effective $X$ error probability matches that under depolarizing noise.
 This MLE curve serves as a benchmark, as a decoder capable of correcting errors of weight up to $3$ should exhibit a performance curve parallel to it at low error rates.
@@
 With $\alpha=1.6$, its performance curve indicates that it can correct nearly all weight-3 errors at LER $10^{-7}$.
-In [60], the BP-OSD-CS60 decoder was shown to closely match the MLE curve on this code.
+In [57], the BP-OSD-CS60 decoder [51] was shown to closely match the MLE curve on this code.
 Since the decoding of $X$ and $Z$ errors is performed independently, the MLE curve represents optimal performance under the assumption that $X$ and $Z$ errors are uncorrelated.
@@
 We observe a similar performance of MBP4+ADOSD4 for $[[15,3,3]]$, $[[65,5,5]]$ and $[[369,9.9]]$ LCS codes as well, while AMBP has an even higher error floor around LER $10^{-2}$ for the $[[369.9,9]]$ LCS code and a lower error floor around $10^{-5}$ for the $[[65,5,5]]$ LCS code.
-For improved performance, $\alpha$ can be adaptively chosen for each $\epsilon$, as demonstrated in Section VI-D.
-VI-B Robustness of the maximum iterations in MBP4+ADOSD4
-Figure 3: MBP4+ADOSD4 decoding with various maximum number of iterations ${\mathrm{T}}$ for the $[[175,7,7]]$ LCS code.
+VII-A2 Robustness of the maximum iterations in MBP4+ADOSD4
+Figure 3: MBP4+ADOSD4 decoding with various maximum number of iterations $\mathrm{T}$} for the $[[175,7,7]]$ LCS code with $\theta=0.999995$.
 Next, we demonstrate that the performance of ADOSD4 remains robust even with variations in the maximum number of iterations in MBP4.
-We note that in ADOSD4, OSD-$w$ is invoked when HRSR fails.
+We note that in ADOSD4, OSD-$w$ is invoked when RSR fails.
 As a result, the overall algorithm maintains high accuracy.
-However, the failures of the HRSR algorithm can impact the simulation time, as OSD-$w$ is significantly more time-consuming.
 Figure 3 illustrates the performance of MBP4+ADOSD4 decoding for the [[175,7,7]] LCS code with varying maximum numbers of iterations, ${\mathrm{T}}=1,2,3,4,5,100,200$.
@@
 Avg. iter. 194.46 97.67 23.51 6.31 2.44 1.70 1.23
-TABLE II: The average number of iterations for MBP4 decoding on the $[[175,7,7]]$ LCS code with ${\mathrm{T}}=200$.
-Table II presents the average number of iterations for ${\mathrm{T}}=200$.
+TABLE I: The average number of iterations for MBP4 decoding on the $[[175,7,7]]$ LCS code with ${\mathrm{T}}=200$.
+Table I presents the average number of iterations for ${\mathrm{T}}=200$.
 It can be observed that the average number of iterations rapidly decreases to one or two when $\epsilon<0.01$.
 Since we are targeting lower error rates, to balance accuracy and efficiency, we typically choose ${\mathrm{T}}=100$.
-VI-C Selection of Reliability Threshold
-We now discuss the selection of the reliability threshold. Figure 4 presents an example for the $[[175,7,7]]$ LCS code, which is representative of our simulation results.
-Based on our experience, at high error rates, the reliability threshold has little impact on performance.
-However, at low error rates, where higher accuracy is required, the choice of $\theta$ depends on both ${\mathrm{T}}$ and the specific code.
-For instance, MBP4+ADOSD4 achieves excellent performance on the $[[175,7,7]]$ LCS code with ${\mathrm{T}}=2$ and $\theta=0.5$ but performs poorly on the (4.8.8) color code with $d=5$ under the same setting.
-With ${\mathrm{T}}=100$ and $\theta=0.999995$, we generally observe good and stable decoding performance across all codes considered, as shown in Figure 4.
-Therefore, we adopt these values for our analysis in the following subsections.
-Figure 4: MBP4+ADOSD4 decoding with various values of reliability threshold $\theta$ for the $[[175,7,7]]$ LCS code.
-VI-D Generalized hypergraph-product codes
-The $[[882,48,16]]$ GHP code [43] is highly degenerate, which causes conventional BP to perform poorly.
-Figure 5 presents the performance curves of different MBP and OSD combinations, where $\alpha$ is either fixed at 1.6 or adaptively chosen as $\alpha(\epsilon)$.
-For comparison, we also include the performance of AMBP4 [32] and BP+mOSD($w$, $\lambda$[43].
-BP+mOSD($w$, $\lambda$) exhibits an early error floor around $\epsilon=0.09$.
-Both MBP4+ADOSD4 and MBP4+OSD4-2 with $\alpha=1.6$ outperform BP+mOSD($15$, $15$[43] by more than an order of magnitude at $\epsilon=0.05$.
-[…]
-Figure 5: Comparison of various decoders on the $[[882,48,16]]$ GHP code. The curve BP+mOSD(15,15) is from [43].
-Figure 6: Comparison of various metrics for reliability order in MBP4-OSD4 on the [[882, 48, 16]] GHP code with $\alpha=1.6$.
-In ADOSD4, we propose using integrated hard-decision and soft reliabilities as a reliability metric (Definition 4).
-Herein we also simulate the performances of MBP4+mOSD based on the reliability metrics of Entropy or Max as shown in Figure 6, including mOSD4  as well.
-Two conclusions are as follows:
-  1. 1.
-    The reliability metric in Definition 4, which incorporates historical message information, outperforms other metrics that rely solely on the last belief distributions.
-  2. 2.
-    The performances of mOSD based on Definition 5, Entropy, or Max are mostly comparable; mOSD-$w$ based on Definition 5 is slightly better than the other two metrics at lower error rate for $w=2$.
-VI-E Bivariate bicycle codes
+VII-B Code capacity noise model decoding simulations
+VII-B1 Bivariate bicycle codes
 BB codes are a class of QLDPC codes proposed in [56], featuring a high error threshold and a significantly higher code rate than topological codes.
@@
 However, the BP+CB decoder in [52] does not perform as well as BP-OSD-0.
-The BP+GDG decoder in [55] shows improved performance, surpassing BP-OSD-0 and being comparable to BP2-OSD-CS10 over independent X errors, as shown in [55, Figure 4].
-Figure 7 presents the simulation results of MBP4+ADOSD4 and AMBP4 decoding for BB codes, both of which demonstrate significantly better performance than GDG decoding.
+The BP+GDG decoder in [53] shows improved performance, surpassing BP-OSD-0 and being comparable to BP2-OSD-CS10 over independent X errors, as shown in [53, Figure 4].
+Figure 4 presents the simulation results of MBP4+ADOSD4 and AMBP4 decoding for BB codes, both of which demonstrate significantly better performance than BP+GDG decoding [53].
 For instance, consider the [[144,12,12]] BB code.
@@
 The $[[72,12,6]]$ and $[[90,8,10]]$ codes closely follow these reference curves, while the other three codes exhibit lower error floors below LER $10^{-7}$.
-Figure 7: AMBP4 and MBP4+ADOSD4 decoding performance for various BB codes.
-VI-F 2D topological codes
-Figure 8: MBP4+OSD4-2, MBP4+ADOSD4, and AMBP4 decoding performance for various $[[d^{2},1,d]]$ rotated surface codes.
-Figure 9: (a) The threshold of MBP4+OSD4-2 on the $[[d^{2},1,d]]$ rotated surface codes with $\alpha=1.0$ is about $17.67\%$ from the simulations.
+Figure 4: AMBP4 and MBP4+ADOSD4 decoding performance for various BB codes with ${\mathrm{T}}=100$ and $\theta=0.999995$.
+VII-B2 2D topological codes
+Figure 5: MBP4+OSD4-2, MBP4+ADOSD4, and AMBP4 decoding performance for various $d^{2},1,d]]$ rotated surface codes with ${\mathrm{T}}=100$ and $\theta=0.999995$.
+Figure 6: (a) The threshold of MBP4+OSD4-2 on the $d^{2},1,d]]$ rotated surface codes with $\alpha=1.0$, $T=100$, and $\theta=0.999995$ is about $17.67\%$ from the simulations.
 For each data point, we collected at least 10,000 logical error events.
@@
 Examining LER around $10^{-6}$ is important, as it helps identify whether an iterative decoder exhibits error floor behavior.
-Figure 8 presents the performance curves of MBP4+ADOSD4 and MBP4+OSD4-2 on $[[d^{2},1,d]]$ rotated surface codes with distances ranging from 5 to 17.
+Figure 5 presents the performance curves of MBP4+ADOSD4 and MBP4+OSD4-2 on $d^{2},1,d]]$ rotated surface codes with distances ranging from 5 to 17.
 They exhibit excellent performance, closely following the dotted reference curves.
-The threshold analysis for $[[d^{2},1,d]]$ rotated surface codes, conducted using the scaling ansatz method [67, 68], is shown in Figure 9.
+The threshold analysis for $d^{2},1,d]]$ rotated surface codes, conducted using the scaling ansatz method [84, 85], is shown in Figure 6.
 The results show a threshold of approximately 17.67% for BP4+OSD4-2, with MBP4+ADOSD4 exhibiting nearly identical performance.
@@
 Similar analyses are performed for toric, twisted XZZX, (6.6.6), and (4.8.8) color codes using both BP4+OSD4-2 and BP4+ADOSD4.
-The summarized threshold results are presented in Table III.
+The summarized threshold results are presented in Table II.
 ADOSD4 and OSD4-2 exhibit nearly identical performances on toric, surface, and twisted XZZX codes, with OSD4-2 performing slightly better on the (6.6.6) and (4.8.8) color codes.
-Figure 10: AMBP4 and MBP4+ADOSD4 decoding performance for various (4.8.8) color codes.
-TABLE III: Thresholds for 2D topological codes subjecte to depolarizing errors. The parameter for MBP4 is set to $\alpha=1.0$.
+TABLE II: Thresholds for 2D topological codes under depolarizing errors. The parameters are $\alpha=1.0$, $T=100$, and $\theta=0.999995$
 […]
-The thresholds for the color codes are lower than those of toric and surface codes.
-This is evident from the performance curves of the (4.8.8) color codes as shown in Figure 10.
-While the slopes of these curves increase as the error rate decreases, they do not closely match the slopes of the dotted reference curves.
-This suggests that some low-weight errors, even those with weight less than $t+1$, may remain uncorrected by these decoders.
-Consequently, additional techniques or higher-order OSD may be necessary.
-TABLE IV: Comparison of the time consumption for surface codes at LER about $10^{-6}$.
+TABLE III: Comparison of the time consumption for surface codes at LER about $10^{-6}$ with $\alpha=1.0$, $T=100$, and $\theta=0.999995$.
 code $\epsilon$ MBP4 time per MBP4 time per time per average iterations ADOSD4 time
-distance failure rate iteration ($\mu$s) ADOSD4 ($\mu$s) OSD4-2 ($\mu$s) when MBP4 succeeds in MBP4 iterations
+[…]
 11 1.7% 21.46% 0.01464 0.0272 0.667 1.002 2
@@
 $^{\dagger}:$ the time per ADOSD4 operation divided by the time per MBP4 iteration.
-VI-G Time consumption
-In the previous subsections, we have demonstrated the effectiveness of our BP+OSD schemes. In this subsection we analyze the time efficiency of our OSD schemes. We fix $\alpha=1$ in MBP4.
-VI-G1 Time Consumption of ADOSD4 vs. OSD4-2 and BP
-Table IV lists the average time consumptions of MBP4, ADOSD4, and OSD4-2 for surface codes with distances of 11, 13, and 15 at LER arouond $10^{-6}$.
-The unit of time is microseconds ($\mu$s) measured using a single-thread desktop computer.
-To achieve LER $10^{-6}$ using MBP4+ADOSD4, the physical error rates are set to 0.017, 0.025, and 0.033 for codes with distances 11, 13, and 15, respectively.
-At least $10^{5}$ samples are collected for each case.
+VII-B3 Time Consumption of ADOSD4 vs. OSD4-2 and BP
+[…]
 It can be observed that the time consumption of ADOSD4 is approximately 2% to 4% of that of OSD4-2, with similar results observed for other quantum codes in this study.
 In summary, ADOSD4 is at least 25 times faster than OSD4-2 at an LER of around $10^{-6}$.
-Table IV also lists the time consumption of ADOSD4 in terms of MBP4 iterations.
+Table III also lists the time consumption of ADOSD4 in terms of MBP4 iterations.
 Specifically, the average time consumption per codeword for ADOSD4 is approximately equivalent to 1.76 to 2.21 MBP4 iterations.
 This shows that the time consumption for ADOSD4 is roughly equivalent to 1 to 2 successful MBP4 executions.
-VI-G2 HRSR and OSD-0 in ADOSD4
-We have shown that ADOSD4 is efficient, thanks to Algorithm 3 and Corollary 10.
-These methods reduce the effective length using reliable subset reduction and allow us to perform only OSD4-0 by identifying degenerate error cosets.
-Now, we examine the effectiveness of these techniques on the surface code of distance 11, with similar results observed for distances 13 and 15.
-Table V enumerates the percentage of ADOSD4 trials that only perform OSD4-0 using Corollary 10.
-At $\epsilon=4.1\%$, 93.61% of ADOSD4 trials already perform only OSD4-0, increasing to 99.84% at $\epsilon=1.7\%$.
-Thus, at lower error rates, almost all trials execute OSD4-0.
-Moreover, the HRSR algorithm reduces the dimensions of the check matrix in both rows and columns.
-Table V also provides the percentages of trials where the number of rows and columns in the check matrix is reduced to 20% or 30% of the original (denoted as $20\%$ and $30\%$ effective dimensions in the table).
-These percentages increase quickly as the error rate decreases.
-With fewer than 30% effective dimensions, the time consumption for Gaussian elimination reduces to roughly $0.3^{3}=2.7\%$, consistent with the previous results.
-TABLE V: Percentage of OSD-0 in ADOSD4 on the surface code of distance $11$
-depolarizing rate $\epsilon$ 1.7% 2.7% 4.1%
-LER $10^{-6}$ $10^{-5}$ $10^{-4}$
-OSD-0 99.84% 98.85% 93.61%
-20% effective 44.13% 26.30% 5.13%
-dimensions
-30% effective 94.04% 84.46% 57.91%
-dimensions
-VII Conclusion
-We proposed quaternary OSD with integrated hard- and soft-decision reliabilities to retain the $X/Z$ correlations.
-Our BP-OSD schemes have demonstrated highly effective decoding capabilities for various sparse quantum codes in the code capacity noise model, including both CSS and non-CSS codes.
-Additionally, when the logical error rate is around $10^{-6}$, ADOSD4 can further enhance the decoding process by offering accelerated decoding while maintaining comparable performance.
-As a result, ADOSD4 is applicable to small or medium-sized codes with $n\approx 1000$, significantly enhancing the performance of BP in the code capacity noise model.
-In our simulations, we restricted the number of error candidates to the size of order-2 OSD; better performance of ADOSD4 can be achieved by allowing for greater options.
-We note that while all the techniques developed in this paper are presented in the context of quaternary error variables, they can naturally be adapted to binary OSD for handling CSS codes.
-This includes hard-decision reliability, highly reliable subset reduction, approximate degenerate decoding, and degenerate candidate pruning conditions, by restricting the approach to either $X$- or $Z$-stabilizers.
-Furthermore, these techniques can be integrated with other methods available in the literature.
-In our investigations, we primarily focused on MBP4+OSD4-2 and ADOSD4.
-However, MBP4+OSD4-0 has also exhibited excellent performance for most codes, aligning with the findings in [43].
-In particular, we found that ADOSD4 mostly performs only MBP4+OSD4-0 at low error rates.
-On the other hand, at low error rates (e.g., below $10^{-3}$), there are small non-isolated errors, so the effective length for OSD is relatively small, allowing us to perform higher-order OSD.
-We have observed some instances of surface codes with order $w$ up to 13 in (16).
-An interesting observation is that we can identify bit-flips in error candidates corresponding to multiplying a stabilizer operator, as stated in Theorem 8 and Corollary 10.
-This information can be used to improve the algorithm by flipping effective bits in approximate degenerate decoding.
-Additionally, another potential improvement is extending the candidate pruning condition by Fossorier et al[47] to quantum codes.
-We chose the binary representation of quantum codes and syndrome correction for developing the OSD algorithm.
-A quantum code can be interpreted as a quaternary additive code with binary error syndromes [6].
-Thus, one might consider formulating a system consisting of $n$ quaternary variables and $n-k$ additive constraints and employ a Gaussian elimination process for quaternary additive groups.
-However, for practical implementation, we chose to binarize this process.
-There are various reliability metrics that consider the (weighted) accumulation of the log-likelihood ratio (LLR) signs for each bit during BP iterations [72, 73, 74, 75].
-Our hard-decision reliability can be viewed as an extension of this idea, focusing on the length of the last run in the hard-decision vectors.
-The reliability order in Definition 4 can be adapted to employ alternative metrics, such as Entropy or Max [43].
-From our experiments, the integrated hard- and soft-decision reliabilities yielded the best results.
-Finally, there is significant progress in circuit-level error decoding [76, 77, 78, 56, 79].
-Various decoders have been shown to work for circuit-level noise models, including BP-OSD-CS, GDG, ambiguity clustering[57], LSD [58], and ordered Tanner forest [80].
-Our BP-OSD scheme can be extended to handle both data and syndrome errors in the phenomenological noise model [81, 82], using techniques proposed in [83, 84, 85], as demonstrated in our recent work [86].
+VII-C Circuit-Level Surface Code Decoding
+We simulate circuit-level decoding of the $d^{2},1,d]]$ rotated surface codes using DEM matrices generated by STIM [71]. For a distance-$d$ code, we simulate $d$ rounds of syndrome extraction.
+All faulty locations are assumed to have the same error rate $\epsilon$.
+We use stim.Circuit with experiment surface_code:rotated_memory_z to generate the syndrome-extraction circuits, and convert the resulting DEM into the binary matrices ${\mathbf{H}}_{\mathrm{DEM}}$, ${\mathbf{L}}_{\mathrm{DEM}}$, and the probability vector ${\mathbf{p}}_{\mathrm{DEM}}(\epsilon)$, where ${\mathbf{p}}_{\mathrm{DEM}}(\epsilon)$ depends on the physical error rate $\epsilon$.
+In the default setting, the first round of detector variables contains detectors for only the $\frac{d^{2}-1}{2}$ $Z$ stabilizers.
+The subsequent d1d-1 rounds contain detector variables for all $d^{2}-1$ $X$ and $Z$ stabilizers.
+In addition, there is a final round of detector variables for the $Z$ stabilizers, which can be regarded as noiseless syndrome measurements.
+As a result, the default STIM circuit for a $d^{2},1,d]]$ rotated surface code yields a total of $d(d^{2}-1)$ parity checks.
+Note that decoders which treat $X$ and $Z$ errors separately, such as the standard MWPM decoder, do not make use of both types of detector variables.
+For comparison, we therefore also generate decoding problems in which the $X$-type detector variables are removed.
+The parameters of the resulting DEM decoding problems are summarized in Table IV.
+Since the default DEM contains more detector variables and hence more syndrome bits, it produces a larger number of error classes, i.e., a larger $N$, because some $Y$ errors can be distinguished from $Z$ errors by the presence of $X$-type detector variables.
+Distance 3 5 7 9
+NN 219 1677 5471 12705
+# Checks 24 120 336 720
+(a) DEM decoding parameters with both $X$ and $Z$ detector variables.
+Distance 3 5 7 9
+NN 55 301 883 1945
+# Checks 16 72 192 400
+(b) DEM decoding parameters with only $Z$ detector variables.
+TABLE IV: Parameters of the DEM decoding problems for rotated surface codes.
+Subtable (a) shows the default setting with both $X$ and $Z$ detector variables, and subtable (b) shows the modified setting with only $Z$ detector variables.
+In our simulations, an error vector $\mathbf{e}$} is sampled according to ${\mathbf{p}}_{\mathrm{DEM}}(\epsilon)$ and its syndrome ${\mathbf{H}}_{\mathrm{DEM}}{\mathbf{e}}^{\top}$ is computed.
+Given ${\mathbf{H}}_{\mathrm{DEM}}$, ${\mathbf{L}}_{\mathrm{DEM}}$, the measured syndrome ${\mathbf{H}}_{\mathrm{DEM}}{\mathbf{e}}^{\top}$, and ${\mathbf{p}}_{\mathrm{DEM}}(\epsilon)$, we apply MBP4+ADOSD4 to this binary decoding problem to obtain an estimate $\hat{{\mathbf{e}}}$.
+The logical $X$ error rate is then estimated using (21) and (22).
+We collected 1000 logical error events for each data point in the plots, except for the cases with $d=9$ and $N=12705$ at $\epsilon=0.002$ and $\epsilon=0.001$.
+VII-C1 Decoding performance
+For computational efficiency, the maximum number of BP iterations is set to $T=10$.
+In particular, we use only soft reliability with threshold $\theta=0.99$, without employing hard-decision history.
+By lowering the reliability threshold, RSR produces a smaller reduced decoding problem.
+The step size is fixed to $\alpha=1.5$, following the choice used in the code-capacity noise model without further optimization.
+Figure 7: MBP4+ADOSD4 and MWPM decoding performance for various $d^{2},1,d]]$ rotated surface codes with $\alpha=1.5$, $T=10$, and $\theta=0.99$ under the circuit-level noise model.
+Error bars represent 95% confidence intervals.
+The simulation results are shown in Fig.
+7, where the MWPM decoder is implemented using PyMatching2 [89] for comparison.
+We observe that MWPM performs better than MBP4+ADOSD4 for DEM instances without $X$-type detector variables.
+However, when $X$ detector variables are included, MBP4+ADOSD4 outperforms MWPM by exploiting the additional syndrome information.
+Moreover, for higher physical error rates ($p>0.003$), the $d=9$ MBP4+ADOSD4 curve outperforms the $d=11$ MWPM curve.
+The error threshold is also improved to approximately $0.76\%$, compared with the $0.7\%$ reported for PyMatching2.
+For reference, as shown in Ref. [90], RelayBP performs worse than MWPM for rotated surface codes. We also remark that MWPM can be adapted to exploit $X/Z$ correlations [91, 92, 93].
+Furthermore, we include a direct comparison with BP+LSD-0 from [55] in Fig.
+8, where our decoder consistently achieves a lower logical error rate.
+Although it is unclear whether [55] uses the default DEM generated by stim, both of our performance curves (with and without $X$ detector variables) outperform BP+LSD-0.
+Figure 8: MBP4+ADOSD4 and BP–LSD-0 decoding performance for $d^{2},1,d]]$ rotated surface codes with $\alpha=1.5$, $T=10$, and $\theta=0.99$ under the circuit-level noise model.
+The BP+LSD-0 curves are from [55] and the algorithm uses a maximum of 30 BP iterations.
+VII-C2 Decoding statistics
+We summarize the decoding statistics for circuit-level surface-code decoding of distance $9$ without $X$-type detector variables in Table V.
+The row “Samples” reports the total number of Monte Carlo shots, while “OSD calls” counts the number of samples for which BP does not converge to a valid solution and OSD is therefore invoked.
+The row “Average length after RSR” shows that the effective code length after reduction is below $5\%$ of the original for $\epsilon\leq 0.005$, and below $1\%$ for $\epsilon=0.001$, demonstrating the strong reduction capability of RSR.
+From the statistics, we observe no stage-1 failures and only a small number of stage-2 failures.
+Even if all stage-2 failures are conservatively counted as logical errors, the overall logical error rate is only marginally affected.
+The row “Avg iterations w/o OSD” corresponds to samples that are successfully decoded by BP alone without invoking OSD.
+Typically, only about 3366 BP iterations are required.
+This justifies setting the maximum number of BP iterations to $T=10$, which is sufficient for RSR to accurately identify reliable and unreliable bits.
+“Time per BP iteration” denotes the average time of a single BP iteration, averaged over all runs.
+“Time per OSD” is the average time of one OSD call, including both ADOSD and the fallback regular OSD when RSR fails.
+We observe that the time per BP iteration is essentially independent of the depolarizing rate $\epsilon$.
+In contrast, the OSD time decreases significantly as $\epsilon$ becomes smaller.
+At $\epsilon=0.01$, one OSD call is equivalent to roughly 250 BP iterations, while at $\epsilon=0.001$ it drops to about 18 BP iterations.
+For reference, in the reported simulations of Relay-BP-5, $R=601$ and $T_{r}=60$, resulting in more than $3.6\times 10^{4}$ BP iterations [90].
+Finally, we remark that the current OSD runtimes do not yet fully reflect the potential speedup enabled by RSR.
+Our implementation embeds the binary DEM decoding problem into a quaternary formalism, which is not optimized for this task and can be further accelerated by a dedicated binary implementation.
+TABLE V: Statistics of circuit-level surface decoding of distance $9$ without $X$ detector variables. $N=1945$.
+depolarizing rate $\epsilon$ 0.01 0.007 0.005 0.003 0.002 0.001
+Samples 100,000 100,000 115990 1311328 9943799 10,000,000
+OSD calls 99966 96895 91954 483438 1632156 389602
+Average length after RSR 229.76 113.37 62.64 32.69 22.67 14.99
+relative code length 0.12 0.06 0.03 0.02 0.012 0.008
+Stage-1 Failures 0 0 0 0 0 0
+Stage-2 Failures 15 43 109 1142 3927 916
+OSD-0 0 0 0 2 107 219
+Avg iterations w/o OSD 5.824 5.691 5.144 4.525 4.059 3.353
+Time per BP iteration (ms) 0.1565 0.1498 0.1289 0.1248 0.1219 0.1185
+Time per OSD (ms) 37.65 27.04 20.71 19.90 8.632 1.987
+VIII Related Work
+Our decoder essentially follows a BP+RSR+OSD pipeline: 1) run BP; 2) if BP does not converge to a valid codeword, use the reliability statistics generated by BP to perform RSR and obtain a reduced linear system; 3) apply OSD to solve the reduced system.
+In this way, the postprocessing stage is significantly accelerated.
+Related work such as Ambiguity Clustering (AC) [54] and Localized Statistics Decoding (LSD) [55] also aim to speed up the postprocessing stage.
+AC employs a divide-and-conquer strategy by decomposing the global decoding problem into multiple localized clusters.
+Each cluster is solved independently to determine its local logical effect, and the local solutions are then linearly combined to reconstruct a global logical correction.
+LSD follows a cluster-growth paradigm inspired by the Union-Find (UF) decoder [94] to localize syndromes.
+It further uses linear-algebraic verification to govern cluster merging and syndrome satisfaction, allowing LSD to be applied to codes without a geometric structure.
+Since AC and LSD can be viewed as high-efficiency alternatives to OSD for postprocessing, it is natural to integrate them into a unified pipeline such as BP+RSR+AC or BP+RSR+LSD.
+Note that RSR requires performing rank-revealing linear-algebraic operations on the reduced system.
+Importantly, essentially the same class of linear-algebraic operations (e.g., pivoting, rank tests, and column selections) is also required by OSD, AC, and LSD in order to initialize clusters, perform algebraic verification, or generate candidate solutions.
+Therefore, applying AC or LSD after RSR does not introduce a fundamentally new computational stage.
+We further note that AC applies degenerate decoding when solving the local subproblems.
+Because these subproblems are much smaller, it becomes feasible to perform fully degenerate decoding by optimizing over logical cosets.
+This observation motivates the use of $\delta$-ADD with $\delta>0$ as a hard-decision rule in OSD after RSR has reduced the problem size sufficiently.
+To summarize, degeneracy is exploited in our algorithm in two distinct ways.
+(1) Since OSD-$w$ is a list decoder, we first use a degeneracy condition (Theorem 8 and Corollary 9) to determine whether multiple logical cosets are possible, and hence whether generating a candidate list is necessary at all.
+(2) Once a reduced problem has been obtained by RSR and a candidate list is generated, we apply the approximate degenerate decoding rule $\delta$-ADD to select the most likely logical coset from the OSD list.
+IX Conclusion
+We proposed an RSR procedure that reduces the decoding problem size using BP-generated statistics with both hard- and soft-decision reliabilities.
+RSR serves as a general-purpose preprocessor that enables post-processing decoders, including OSD, LSD, and AC, to operate efficiently on the reduced problems.
+We further introduced a degeneracy-aware pruning condition to identify when higher-order OSD is beneficial.
+Combined with MBP, this yields the MBP–RSR–ADOSD family of decoders, which achieves strong and consistent performance across a broad range of sparse quantum codes under the code-capacity noise model, including both CSS and non-CSS codes, as well as under circuit-level DEM noise for rotated surface codes.
+A key practical result is that MBP–RSR–ADOSD4 scales to decoding problems with more than $10^{4}$ error variables.
+At low physical error rates (e.g., below $10^{-3}$), errors are typically sparse and isolated, so RSR produces a small effective problem size, making higher-order OSD both computationally feasible and highly effective in this regime.
+More general circuit-level decoding problems with strongly correlated syndrome structures [70] are left for future work.
+The source files of our simulations can be found in [95].

### v4

@@
 Through extensive simulations, we demonstrate improved performance over MWPM and Localized Statistics Decoding for a variety of CSS and non-CSS codes under the code-capacity noise model, and for rotated surface codes under realistic circuit-level noise.
-At low physical error rates, RSR reduces the effective problem size to less than 5%, enabling higher-order OSD with accelerated runtime.
+At low physical error rates, RSR reduces the effective problem size to as little as 1% (e.g., for $\epsilon=0.001$ in surface-code DEM), enabling higher-order OSD with drastically reduced computational complexity.
 These results highlight the practical efficiency and broad applicability of the BP+ADOSD framework for both theoretical and realistic quantum error correction scenarios.
@@
 When BP fails to converge to a valid global solution, it still provides rich statistical information about the error configuration.
-In particular, motivated by density evolution arguments for sparse LDPC codes [58, 59], most error variables are typically resolved with high confidence, except for a small number of unsettled variables associated with trapping sets [30].
+In particular, observing the convergence behavior of belief propagation on sparse graphs, most error variables are typically resolved with high confidence, except for a small number of unsettled variables associated with trapping sets [30].
 Consequently, a large fraction of variables can be regarded as already determined and effectively part of the solution.
@@
 When both $X$- and $Z$-type detector variables are used, MBP4+ADOSD4 also surpasses the minimum-weight perfect-matching (MWPM) decoder [73, 74] and achieves an improved threshold of approximate $0.76\%$.
-Beyond improved decoding performance, a central outcome of our simulations is the strong reduction effect produced by reliable subset reduction (RSR): at low physical error rates, RSR typically reduces the effective decoding problem size by one to two orders of magnitude (e.g., below $5\%$ of the original size for $\epsilon\leq 0.005$, and below $1\%$ for $\epsilon=0.001$ in surface-code DEM problems).
+Beyond improved decoding performance, a central outcome of our simulations is the strong reduction effect produced by reliable subset reduction (RSR): at low physical error rates, RSR typically reduces the effective decoding problem size by one to two orders of magnitude (e.g., below $4\%$ of the original size for $\epsilon\leq 0.005$, and below $1\%$ for $\epsilon=0.001$ in surface-code DEM problems).
 This dramatic dimensionality reduction enables higher-order OSD to be applied efficiently even for problems with more than $10^{4}$ error variables, a regime that is far beyond the reach of conventional OSD-based postprocessing.
 As a result, our MBP–RSR–ADOSD approach achieves strong threshold performance for topological codes and substantially improved logical error rates in the low-error regime.
+Crucially, these results are achieved under a strict budget of at most 10 BP iterations.
+In practice, when MBP4 converges, it typically does so within 3 to 6 iterations.
 This paper is organized as follows.
@@
 The remaining unsettled variables are usually associated with trapping sets [30], and additional post-processing techniques are required to resolve them.
-In this section, we estimate the expected size of these unsettled error variables using density evolution [58, 59].
-III-A Density-Evolution-Guided Reliability Heuristic
-We consider a family of regular LDPC codes over a quaternary symmetric channel and use the syndrome-based Gallager-A algorithm [15] as a heuristic approximation of quantum stabilizer decoding under depolarizing noise.
-Density evolution is then employed to characterize the behavior of this iterative decoder.
-Gallager-A is a hard-decision message-passing algorithm defined on a Tanner graph.
-In our setting, variable nodes $v_{j}$ represent Pauli errors on physical qubits, while check nodes $c_{i}$ represent stabilizer generators.
-Since a Pauli operator either commutes or anticommutes with a stabilizer, all messages are binary: $1$ indicates anti-commutation and 0 indicates commutation.
-We consider a syndrome-based quaternary–binary variant of Gallager-A suitable for CSS-type stabilizer codes, where each stabilizer contains only $X$ or $Z$ operators.
-Given a check matrix $\mathbf{H}$} and syndrome ${\mathbf{s}}=(s_{1},\dots,s_{m})$, the decoding proceeds as follows.
-  1. 1.
-    Initialization: Set estimated error $\hat{E}_{j}=I$ for all qubits and initialize all variable-to-check messages
-    $m_{v_{j}\to c_{i}}^{(0)}=0.$
-  2. 2.
-    Check update: At iteration $\ell$, each check node computes a recommendation message
-    $m_{c_{i}\to v_{j}}^{(\ell)}=s_{i}\oplus\bigoplus_{k\in{\cal I}(c_{i})\setminus\{j\}}m_{v_{k}\to c_{i}}^{(\ell)}.$
-  3. 3.
-    Variable update: Each variable node $v_{j}$ updates its estimate according to the unanimous recommendations of its neighboring checks:
-    • $X$-flip: if all incoming messages from $Z$-type checks are $1$, set $\hat{E}_{j}\leftarrow\hat{E}_{j}X$;
-    • $Z$-flip: if all incoming messages from $X$-type checks are $1$, set $\hat{E}_{j}\leftarrow\hat{E}_{j}Z$.
-    (If both updates occur, a $Y$ correction is effectively applied up to a global phase.)
-  4. 4.
-    Message update: Set $m_{v_{j}\to c_{i}}^{(\ell+1)}=1$ if $\hat{E}_{j}$ anticommutes with the stabilizer of $c_{i}$, and 0 otherwise.
-  5. 5.
-    Termination: The algorithm stops if the current estimate $\hat{\mathbf{E}}$ is valid.
-For a CSS-type stabilizer code, $X$-type and $Z$-type errors are processed independently by the modified Gallager-A algorithm.
-The correlations between $X$ and $Z$ errors induced by the depolarizing channel are therefore neglected at this stage.
-Accordingly, we focus on the analysis of $X$-error decoding, which reduces to a classical binary decoding problem: estimating an $n$-bit vector ${\mathbf{e}}=(e_{1},\dots,e_{n})\in\{0,1\}^{n}$ from its syndrome.
-Suppose that the $Z$-type stabilizers define a $(d_{c},d_{v})$-regular LDPC code, where each check node is connected to $d_{c}$ variable nodes and each variable node is connected to $d_{v}$ check nodes.
-Proposition 2.
-[Heuristic Density Evolution Analysis] Consider a quantum code of length $n$ under depolarizing noise with rate $\epsilon$.
-Assume that the Tanner graph induced by the $Z$ stabilizers is locally tree-like and that decoding is performed using the modified Gallager-A algorithm.
-Let $x^{(\ell)}$ denote the probability that a variable-to-check message is incorrect at iteration $\ell$, and let $y^{(\ell)}$ denote the probability that a check-to-variable message is incorrect.
-Then the message error probabilities evolve according to
-$y^{(\ell)}$ $=\frac{1-(1-2x^{(\ell)})^{d_{c}-1}}{2},$ (6)
-$x^{(\ell+1)}$ $=p_{a}\big[1-(1-y^{(\ell)})^{d_{v}-1}\big]+(1-p_{a})(y^{(\ell)})^{d_{v}-1},$ (7)
-where $p_{a}=\frac{2\epsilon}{3}$ is the probability that a Pauli error anticommutes with a $Z$-type stabilizer, and $x^{(0)}=p_{a}$.
-The recursion (6)–(7) follows from the standard density-evolution analysis of Gallager-A decoding for regular LDPC codes over the binary symmetric channel [15, 58, 59].
-At error rates below the heuristic threshold, density evolution predicts that most variable nodes converge rapidly, leaving a fraction proportional to the channel error probability $\epsilon$ unresolved.
-The expected number of erroneous variable nodes at iteration $\ell$ is $n\cdot x^{(\ell)}$, and after an average convergence time $\ell_{\mathrm{succ}}$ the residual error set has size roughly $n\cdot x^{(\ell_{\mathrm{succ}})}$.
-For finite-length codes, cycles in the Tanner graph cause deviations from the infinite-length density evolution prediction.
-As $\epsilon$ decreases, $x^{(\ell_{\mathrm{succ}})}$ rapidly vanishes, leaving extremely sparse residual errors.
-Unlike classical LDPC ensembles, stabilizer Tanner graphs inherently contain many short cycles due to commutation constraints.
-As a result, the above density-evolution analysis should be interpreted as a heuristic guideline rather than a rigorous asymptotic prediction.
-Although one could in principle extend DE to full belief propagation with real-valued LLRs, the Gallager-A approximation already captures the main qualitative behavior of the decoding process.
-Nonetheless, our heuristic results serve as a useful reference for BP algorithms that exploit memory effects to mitigate the impact of short cycles [32, 77].
-III-B Reliable subset reduction for the syndrome decoding problem
+III-A Reliable subset reduction for the syndrome decoding problem
 In the syndrome decoding problem, we seek a binary vector ${\mathbf{e}}\in\{0,1\}^{2n}$ satisfying
@@
 When BP fails to converge to a valid solution within $\mathrm{T}$} iterations, a postprocessing algorithm such as OSD is typically required.
-Motivated by density evolution, we expect that after a few BP iterations, most variables have already converged to highly reliable values, while only a small subset remains ambiguous.
+We expect that after a few BP iterations, most variables have already converged to highly reliable values, while only a small subset remains ambiguous.
 We denote by $R\subset\{e_{1},\dots,e_{2n}\}$ a subset of reliable variables, whose values are assumed to be correct and fixed.
@@
 $\hat{{\mathbf{e}}}=\sigma^{-1}\big(\hat{{\mathbf{e}}}^{\prime},{\mathbf{e}}^{R}\big).$
-III-C Reliability order based on BP4
+III-B Reliability order based on BP4
 To quantify the reliability of each error-variable estimate, we combine two sources of information: (i) the hard-decision trajectory of BP across all iterations, and (ii) the belief distribution produced at the final BP iteration.
@@
 These techniques operate on the residual errors left by BP and can include methods such as OSD, LSD, or other syndrome-based decoders.
-In this section, we first review the basic OSD procedure, which relies on the reliability order defined in Section III-C.
-When combined with the RSR method introduced in Section III-B, OSD can operate on a reduced system of variables, thereby lowering computational complexity.
+In this section, we first review the basic OSD procedure, which relies on the reliability order defined in Section III-B.
+When combined with the RSR method introduced in Section III-A, OSD can operate on a reduced system of variables, thereby lowering computational complexity.
 In the context of OSD, a critical step involves identifying and sorting the accurate and reliable coordinates to establish $n+k$ linearly independent bits.
 Then, we can use these coordinates to generate a valid error that matches the syndrome.
-To accomplish this, we use the reliability orders defined in Section III-C.
+To accomplish this, we use the reliability orders defined in Section III-B.
 Let $\tilde{{\mathbf{e}}}$ be the hard-decision output from BP4 after $\mathrm{T}$} iterations.
@@
 4:Let $\mu(\pi(\tilde{{\mathbf{e}}}))=\left[\tilde{{\mathbf{e}}}^{U}~~\tilde{{\mathbf{e}}}^{R}\right]$, where $\tilde{{\mathbf{e}}}^{U}$ represents the independent and unreliable subset of nkn-k bits and $\tilde{{\mathbf{e}}}^{R}$ is the reliable subset of $n+k$ bits.
-Use the reliable subset $\tilde{{\mathbf{e}}}^{R}$ and the modified parity-check condition 12 to select feasible unreliable bits
+Use the reliable subset $\tilde{{\mathbf{e}}}^{R}$ and the modified parity-check condition 10 to select feasible unreliable bits
 $\tilde{{\mathbf{e}}}^{U}={{\mathbf{s}}^{\prime}}^{\top}\oplus{\tilde{{\mathbf{e}}}^{R}}{\mathbf{A}}^{\top}.$ (11)
@@
 IV-C Reliable Subset Reduction combined with OSD
-As introduced in Section III-B, the RSR procedure removes highly reliable error variables from the parity-check constraints, producing a reduced linear system.
+As introduced in Section III-A, the RSR procedure removes highly reliable error variables from the parity-check constraints, producing a reduced linear system.
 This reduced system can then be solved using OSD or any other suitable combinatorial decoder to estimate the remaining unreliable variables.
@@
 […]
-TABLE III: Comparison of the time consumption for surface codes at LER about $10^{-6}$ with $\alpha=1.0$, $T=100$, and $\theta=0.999995$.
-code $\epsilon$ MBP4 time per MBP4 time per time per average iterations ADOSD4 time
-[…]
-11 1.7% 21.46% 0.01464 0.0272 0.667 1.002 2
-13 2.5% 35.97% 0.0212 0.0469 1.342 1.79 2.21
-15 3.3% 51.47% 0.0287 0.0506 2.155 2.41 1.76
-$^{\dagger}:$ the time per ADOSD4 operation divided by the time per MBP4 iteration.
-VII-B3 Time Consumption of ADOSD4 vs. OSD4-2 and BP
-[…]
-It can be observed that the time consumption of ADOSD4 is approximately 2% to 4% of that of OSD4-2, with similar results observed for other quantum codes in this study.
-In summary, ADOSD4 is at least 25 times faster than OSD4-2 at an LER of around $10^{-6}$.
-Table III also lists the time consumption of ADOSD4 in terms of MBP4 iterations.
-Specifically, the average time consumption per codeword for ADOSD4 is approximately equivalent to 1.76 to 2.21 MBP4 iterations.
-This shows that the time consumption for ADOSD4 is roughly equivalent to 1 to 2 successful MBP4 executions.
 VII-C Circuit-Level Surface Code Decoding
@@
 For comparison, we therefore also generate decoding problems in which the $X$-type detector variables are removed.
-The parameters of the resulting DEM decoding problems are summarized in Table IV.
+The parameters of the resulting DEM decoding problems are summarized in Table III.
 Since the default DEM contains more detector variables and hence more syndrome bits, it produces a larger number of error classes, i.e., a larger $N$, because some $Y$ errors can be distinguished from $Z$ errors by the presence of $X$-type detector variables.
@@
 (b) DEM decoding parameters with only $Z$ detector variables.
-TABLE IV: Parameters of the DEM decoding problems for rotated surface codes.
+TABLE III: Parameters of the DEM decoding problems for rotated surface codes.
 Subtable (a) shows the default setting with both $X$ and $Z$ detector variables, and subtable (b) shows the modified setting with only $Z$ detector variables.
@@
 We collected 1000 logical error events for each data point in the plots, except for the cases with $d=9$ and $N=12705$ at $\epsilon=0.002$ and $\epsilon=0.001$.
-VII-C1 Decoding performance
 For computational efficiency, the maximum number of BP iterations is set to $T=10$.
@@
 The error threshold is also improved to approximately $0.76\%$, compared with the $0.7\%$ reported for PyMatching2.
-For reference, as shown in Ref. [90], RelayBP performs worse than MWPM for rotated surface codes. We also remark that MWPM can be adapted to exploit $X/Z$ correlations [91, 92, 93].
+For reference, as shown in [88], RelayBP performs worse than MWPM for rotated surface codes. We also remark that MWPM can be adapted to exploit $X/Z$ correlations [89, 90, 91].
 Furthermore, we include a direct comparison with BP+LSD-0 from [55] in Fig.
@@
 The BP+LSD-0 curves are from [55] and the algorithm uses a maximum of 30 BP iterations.
-VII-C2 Decoding statistics
-We summarize the decoding statistics for circuit-level surface-code decoding of distance $9$ without $X$-type detector variables in Table V.
-The row “Samples” reports the total number of Monte Carlo shots, while “OSD calls” counts the number of samples for which BP does not converge to a valid solution and OSD is therefore invoked.
-The row “Average length after RSR” shows that the effective code length after reduction is below $5\%$ of the original for $\epsilon\leq 0.005$, and below $1\%$ for $\epsilon=0.001$, demonstrating the strong reduction capability of RSR.
-From the statistics, we observe no stage-1 failures and only a small number of stage-2 failures.
+VII-D Decoding statistics
+We summarize the decoding statistics for circuit-level surface-code decoding of distance $9$ without $X$-type detector variables in Table IV, using $\alpha=1.5$, $T=10$, and $\theta=0.99$.
+The row “Shots” reports the total number of Monte Carlo shots, while “OSD calls/Shots” is the percentage of shots for which BP does not converge to a valid solution and OSD is therefore invoked.
+The row “Average length after RSR” shows that the effective code length after reduction is below $4\%$ of the original for $\epsilon\leq 0.005$, and below $1\%$ for $\epsilon=0.001$, demonstrating the strong reduction capability of RSR.
+From the statistics, we observe no stage-1 failures and only a small percentage of stage-2 failures.
 Even if all stage-2 failures are conservatively counted as logical errors, the overall logical error rate is only marginally affected.
@@
 This justifies setting the maximum number of BP iterations to $T=10$, which is sufficient for RSR to accurately identify reliable and unreliable bits.
-“Time per BP iteration” denotes the average time of a single BP iteration, averaged over all runs.
-“Time per OSD” is the average time of one OSD call, including both ADOSD and the fallback regular OSD when RSR fails.
-We observe that the time per BP iteration is essentially independent of the depolarizing rate $\epsilon$.
-In contrast, the OSD time decreases significantly as $\epsilon$ becomes smaller.
-At $\epsilon=0.01$, one OSD call is equivalent to roughly 250 BP iterations, while at $\epsilon=0.001$ it drops to about 18 BP iterations.
 For reference, in the reported simulations of Relay-BP-5, $R=601$ and $T_{r}=60$, resulting in more than $3.6\times 10^{4}$ BP iterations [88].
-Finally, we remark that the current OSD runtimes do not yet fully reflect the potential speedup enabled by RSR.
-Our implementation embeds the binary DEM decoding problem into a quaternary formalism, which is not optimized for this task and can be further accelerated by a dedicated binary implementation.
-TABLE V: Statistics of circuit-level surface decoding of distance $9$ without $X$ detector variables. $N=1945$.
+Finally, we remark that the current software runtimes do not yet fully reflect the intrinsic speedup enabled by RSR.
+Our current research-oriented implementation embeds the binary DEM decoding problem into a quaternary formalism, which incurs unnecessary computational overhead.
+A dedicated binary implementation is expected to further reduce the wall-clock time.
+TABLE IV: Statistics of circuit-level surface decoding ($d=9,N=1945$) without $X$ detector variables. Parameters: $\alpha=1.5,T=10,\theta=0.99$. $n_{red}$ denotes the reduced variable length by RSR.
 Depolarizing rate $\epsilon$ 0.01 0.007 0.005 0.003 0.002 0.001
-Samples 100,000 100,000 115990 1311328 9943799 10,000,000
-OSD calls 99966 96895 91954 483438 1632156 389602
-Average length after RSR 229.76 113.37 62.64 32.69 22.67 14.99
-relative code length 0.12 0.06 0.03 0.02 0.012 0.008
+Shots $10^{6}$ $10^{6}$ $10^{6}$ $10^{6}$ $10^{7}$ $10^{7}$
+OSD calls / Shots $99.97\%$ $96.85\%$ $79.12\%$ $36.88\%$ $16.41\%$ $3.90\%$
+Average length after RSR ($n_{red}$) 229.53 113.27 62.34 32.85 22.67 14.99
+Relative code length ($n_{red}/N$) 0.118 0.058 0.032 0.017 0.012 0.008
 Stage-1 Failures 0 0 0 0 0 0
-Stage-2 Failures 15 43 109 1142 3927 916
-OSD-0 0 0 0 2 107 219
-Avg iterations w/o OSD 5.824 5.691 5.144 4.525 4.059 3.353
-Time per BP iteration (ms) 0.1565 0.1498 0.1289 0.1248 0.1219 0.1185
-Time per OSD (ms) 37.65 27.04 20.71 19.90 8.632 1.987
+Stage-2 Failures / Shots $1.19\times 10^{-4}$ $4.33\times 10^{-4}$ $61\times 10^{-4} 8.8$$88\times 10^{-4}$ $3.95\times 10^{-4}$ $9.16\times 10^{-5}$
+Avg. BP iterations (converged) 6.21 5.67 5.15 4.53 4.07 3.50
+The order-of-magnitude reduction in the residual problem size ensures that the computational cost of the ADOSD stage remains comparable to only a few BP iterations.
+This low-complexity feature is consistent across both code-capacity and circuit-level noise models, facilitating the implementation of high-order OSD in large-scale quantum error correction systems.
 VIII Related Work